### v1

# Gradient Interference-Aware Graph Coloring for Multitask Learning

Abstract
When different objectives conflict with each other in multi-task learning, gradients begin to interfere and slow convergence, thereby reducing the final model’s performance. To address this, we introduce a scheduler that computes gradient interference, constructs an interference graph, and then applies greedy graph-coloring to partition tasks into groups that align well with each other. At each training step, only one group (color class) of tasks are activated. The grouping partition is constantly recomputed as task relationships evolve throughout training. By ensuring that each mini-batch contains only tasks that pull the model in the same direction, our method improves the effectiveness of any underlying multi-task learning optimizer without additional tuning. Since tasks within these groups will update in compatible directions, model performance will be improved rather than impeded. Empirical results on six different datasets show that this interference-aware graph-coloring approach consistently outperforms baselines and state-of-the-art multi-task optimizers.
1 Introduction
As modern applications increasingly span vision, language, speech, and beyond (generating massive heterogeneous datasets) Kokkinos2016UberNet ; Baltrusaitis2018 ; Liang2022Foundations there is a growing need for models that can learn from and operate across multiple tasks simultaneously. Multi-task learning (MTL) trains a single model to solve several tasks at once, sharing knowledge across them to use data and compute more efficiently Caruana1997 . Early studies showed that this idea can boost accuracy on small vision and language problems, but they also revealed major drawbacks Sener2018 ; Shi2023Recon . When two tasks push the shared network in opposite directions their gradients clash, slowing or even reversing learning. Classic fixes, like manual loss weights or static task schedules Argyriou2007MultiTaskFeature ; AndoZhang2005 ; Evgeniou2005 ; Kang2011 help only in narrow settings and must be tuned for every new dataset Yu2021RevisitingLossWeight .
Recent work has shifted towards gradient-level solutions. Projection methods such as PCGrad Yu2020PCGrad remove the part of one task’s gradient that conflicts with another. Adaptive-rate approaches like AdaTask Yang2023AdaTask shrink the learning rate for any task whose gradients grow too large. These approaches reduce interference in each update, yet they still mix all tasks together every step. As a result, strongly negative pairs continue to impact progress Shi2023Recon ; Li2025PiKE , and the solutions themselves become noisy because task relationships change during training.
Our work tackles this problem by adjusting when tasks are trained rather than how their gradients are modified. We propose a lightweight scheduler that measures recent gradient conflict, builds a conflict graph, and then uses greedy graph coloring to group only compatible tasks into the same optimization step. Every few iterations the scheduler recomputes the graph so the groups adapt as learning proceeds. As our approach can be used in tandem with standard optimizers and add negligible overhead, it can be implemented into existing systems. Experiments on six datasets show consistent gains over strong baselines and stand-alone state-of-the-art approaches.
Our contributions are as follows:
  1. 1.
    We introduce an interference-aware task scheduler based on greedy graph coloring that separates highly conflicting tasks without manual tuning.
  2. 2.
    We build upon and apply existing theory to prove the effectiveness of components within the scheduler
  3. 3.
    We combine the scheduler with PCGrad, AdaTask, and GradNorm and demonstrate systematic improvements across six datasets.
  4. 4.
    We perform ablation studies to prove that dynamic recoloring and history-averaged conflict estimates are essential for high performance.
2 Related Works
Multi-task learning (MTL) methods have evolved from simple loss-weighting approaches to larger and more sophisticated optimization techniques that manage task conflict and cooperation Yang2023AdaTask . Early adaptive-weighting approaches sought to balance losses automatically 9336293 ; Fan2023MaxGNR , while more recent work modifies gradients directly Yu2020PCGrad . Task scheduling and grouping methods, though far less popular than adaptive weighting techniques torbarina2023challengesopportunitiesusingtransformerbased , have contributed to the field by controlling the timing of updates. Our interference-aware scheduler combines these lines of research, offering lightweight, adaptive grouping with provable convergence.
2.1 Tuned Loss Weighting
From early MTL work it became clear that simply summing task losses often favors one objective at the expense of others Kurin2022InDefense ; Zhao2024ExcessMTL ; Mueller2022Temperature , especially when losses have different scales or noise levels. To address this, practitioners manually tuned per‑task weight coefficients ($\lambda$-values) to rebalance learning Liang2020 ; lin2022a , but this process was laborious and dataset-specific. Thus, researchers began to develop automated methods.
2.2 Adaptive Loss Weighting
(Kendall et al., 2018) kendall2018multi introduced uncertainty weighting, learning each task’s homoscedastic (constant‑variance) Bishop2006 noise to scale losses automatically and improve depth and semantics on NYUv2 Silberman2012 .
GradNorm automatically balances multiple loss functions by tuning each task’s gradient magnitude so that all tasks train at comparable speeds Chen2017GradNorm . It does this by introducing a single asymmetry hyperparameter $\alpha$ that governs how much each task’s loss is called. This eliminates the need for expensive grid searchers over manual weights. GradNorm was also a major leap empirically as it surpassed exhaustive search baselines on both regression and classification tasks. Dynamic Weight Averaging (DWA) extended this idea by adjusting weights based on loss rate of change, reducing oscillations between tasks Liu2019EndToEnd .
More recently AdaTask applies task-specific learning rates that adapt to each head’s gradient norm, yielding significant gains on multi-label classification benchmarks Yang2023AdaTask .
2.3 Gradient-Level Conflict Mitigation
Rather than rescaling losses, gradient surgery methods alter update directions. PCGrad projects gradients that conflict (negative cosine) onto each other’s normal plane, significantly boosting efficiency on supervised vision and RL problems Yu2020PCGrad . CAGrad frames task balance as a min-max optimization, finding updates that maximize the worst-case task improvement Liu2021CAGrad . The Multiple Gradient Descent Algorithm (MGDA) computes a Pareto-optimal convex combination of task gradients, ensuring no task is harmed Sener2018 . More recent variants such as SAM-GS incorporate momentum into conflict detection, smoothing gradient estimates while preserving the benefits of surgery Borsani2025SAMGS .
2.4 Empirical Task Grouping
Task grouping aims to decide which tasks should train together so that helpful transfer is amplified and harmful interference is limited. It typically groups tasks into subsets that update jointly, rather than updating all tasks at once. This is different from approaches that keep all tasks active or reweight the joint gradient (adaptive loss weighting, gradient surgery).
Early approaches under this category used round-robin and random sampling-based approaches that ignored any task relationships McCann2018NaturalLanguageDecathlon ; Zamir2020RobustLearning . (Standley et al., 2020) exhaustively searches over small subsets to identify beneficial groupings, demonstrating the potential of selective updates but failing to scale beyond eight tasks due to computational complexity Standley2020TaskGrouping .
Task Affinity Groupings (TAG) Fifty2021 performs one joint training run to measure inter-task ’affinity’. It quantifies how an update for task $i$ (its gradient) would change task $j$’s loss, and it uses these cross-effects to select partitions of tasks that should share updates. The key idea is to treat grouping as an outcome of measured gradient interactions.
Ayman et al. ayman2023taskgroupingautomatedmultitask train a predictor that maps single-task statistics and dataset features to an estimate of whether two or more tasks should be grouped. They then use that predictor to guide a randomized search over groups, which dramatically reduces the number of multi-task trainings (or ’MTL trials’) needed to find a good partition.
Using a completely different approach, Towards Principled Task Grouping (PTG) wang2024towards formulates grouping as a mathematical program with a theoretically motivated objective capturing beneficial transfer while respecting resources constraints (e.g., compute budgets). It builds a principled optimization over candidate groups that is meant to generalize across application domains.
Scalable Task Grouping via Training Dynamics (STG-MLT) Sherif_2024 avoids expensive affinity estimation by extracting Data Maps swayamdipta-etal-2020-dataset (simple summaries of training dynamics per task) and then clustering tasks using those features. The clusters are intended to push for positive transfer at larger scale. This approach essentially replaces gradient cross-effects with more compact trajectory features that are cheap to compute and easy to cluster.
Selective Task Group Updates (STGU) jeong2025selective moves from static partitions to online grouping with sequential updates. It tracks running estimate of cross-task interaction during training, forms groups, and then updates one group per branch. This couples grouping with the step-to-step training dynamics. The method’s core mechanism is to recompute grouping signals as the model evolves and to partition groups accordingly.
3 Problem Setup
We formalize multi-task learning (MTL) Caruana1997 as the optimization of a shared network under scheduled task activation. Each task contributes a loss whose gradients may align or conflict. We quantify conflict with the negative-cosine interference coefficient, embed all tasks in a conflict graph, and later use that graph to derive our schedule. This section fixes notation and states the optimization goal that the remainder of the methodology addresses.
3.1 Data and Notation
Let $\mathcal{T}={T_{1},\dots,T_{K}}$ be the set of learning tasks. Let $\theta\in\mathbb{R}^{d}$ denote shared network parameters and $\phi_{k}\in\mathbb{R}^{d_{k}}$ denote task‑private parameters for $T_{k}$.
Each task provides a dataset $\mathcal{D}_{k}={(x_{i}^{(k)},y_{i}^{(k)})}_{i=1}^{n_{k}}$ and a differentiable loss $\mathcal{L}_{k}(\theta,\phi_{k};\mathcal{D}_{k})$ (e.g., cross-entropy, segmentation losses, etc.). Throughout, we write
$g_{k}=\nabla_{\theta}\mathcal{L}_{k}(\theta,\phi_{k})\quad\text{and}\quad h_{k}=\nabla_{\phi_{k}}\mathcal{L}_{k}(\theta,\phi_{k})$ (1)
for the stochastic gradients obtained from a mini-batch sampled within $\mathcal{D}_{k}$.
3.1.1 Interference Coefficient
Task interactions are captured by the interference coefficient
$\rho_{ij}\;=\;-\frac{\langle g_{i},g_{j}\rangle}{\|g_{i}\|\,\|g_{j}\|},$ (2)
introduced in PCGrad to detect gradient conflict Yu2020PCGrad and later adopted by CAGrad Liu2021CAGrad and GradNorm variants Chen2017GradNorm . Positive $\rho_{ij}$ indicates that following $g_{i}$ harms $T_{j}$ (negative cosine similarity), whereas $\rho_{ij}\leq 0$ implies alignment or neutrality.
3.1.2 Conflict Graph
Fix a tolerance $\tau\in(0,1)$. The conflict graph is
$G_{\tau}=(\mathcal{T},E_{\tau}),\quad E_{\tau}=\bigl\{(i,j)\,:\,\rho_{ij}>\tau\bigr\}.$ (3)
Vertices represent tasks; an edge links any pair whose gradients conflict beyond $\tau$. Greedy graph‑coloring of a graph with maximum degree $\Delta$ uses at most $\Delta+1$ colors bonamy2018boundingchifractiondelta .
3.2 Goal
At training step $t$ we choose an active task set $S_{t}\subseteq\mathcal{T}$ and perform an SGD update only on those tasks:
[…]
The scheduled MTL objective is therefore
$\min_{\theta,\{\phi_{k}\}}\;F(\theta,\phi)=\sum_{t=1}^{T}\;\sum_{k\in S_{t}}\;\mathcal{L}_{k}\bigl(\theta_{t},\phi_{k,t}\bigr).$ (5)
Our overarching aim is to design the sequence ${S_{t}}_{t=1}^{T}$ so that: (1) Every task is visited regularly (to retain representation quality) and (2) Conflicting tasks seldom appear together (to avoid destructive interference).
[…]
4 Proposed Approach
Figure 1: Interference-aware scheduling pipeline: (a) For each task $T_{i}$ (circles $T_{1}\ldots T_{6}$), we smooth recent per-step gradients with an Exponential Moving Average (EMA); (b) From these EMA vectors we compute the pairwise cosine matrix. In the figure, cells outlined with red dashes mark pairs with cosine $\tau$. These are flagged as conflicts; (c) We build the conflict graph whose nodes are tasks $T_{i}$ and whose red dashed edges connect exactly those pairs identified in (b); (d) We apply greedy graph coloring so that no conflict edge lies within a color, producing low-conflict groups. In the example shown, we have two groups: A as blue and B as orange; (e) During training we activate one group per step. After every $R$ steps (here, $R=4$) we ’refresh’ and run the pipeline again from step A, where we update the EMAs with the latest gradients.
We design an interference-aware scheduler that partitions tasks into low-conflict groups and activates exactly one group per optimization step. The procedure consists of four stages: (1) estimating pairwise interference, (2) building and coloring the conflict graph, (3) generating a periodic schedule, and (4) updating that schedule as training evolves.
4.1 Estimating Gradient Interference
At step $t$ and for every task $T_{k}$ appearing in the current mini‑batch we compute a task‑specific stochastic gradient
$g_{k}^{(t)}=\nabla_{\theta}\mathcal{L}_{k}\!\bigl(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}\bigr),$ (6)
using an independent sub‑batch $\mathcal{B}_{k}^{(t)}\subset\mathcal{D}_{k}$. We then update an exponential moving average
$\tilde{g}_{k}^{(t)}=\beta\,\tilde{g}_{k}^{(t-1)}+(1-\beta)\,g_{k}^{(t)},\qquad\beta\in[0,1),$ (7)
which stabilizes cosine estimates while requiring only two buffers per task (current and previous). Whenever we refresh the schedule (every $R$ steps) we form the pairwise interference matrix
$\rho_{ij}^{(t)}\;=\;-\frac{\langle\tilde{g}_{i}^{(t)},\tilde{g}_{j}^{(t)}\rangle}{\|\tilde{g}_{i}^{(t)}\|\,\|\tilde{g}_{j}^{(t)}\|},\qquad i,j\in\{1,\dots,K\}.$ (8)
Computing all K(K1)/2K(K-1)/2 cosines is $O(K^{2}d)$ with $d$ representing the shared‑parameter dimension.
4.2 Conflict Graph Construction
Given a tolerance $\tau\in(0,1)$, the conflict graph at update round $r$ is
$G^{(r)}_{\tau}=(\mathcal{T},E_{\tau}^{(r)}),\quad E_{\tau}^{(r)}=\bigl\{(i,j)\;:\;\rho^{(t_{r})}_{ij}>\tau\bigr\}.$ (9)
Edges connect tasks whose averaged gradients have cosine similarity less than $\tau$. Intuitively, higher $\tau$ yields a sparser graph and longer groups, trading off less conflict for fewer gradient updates per task.
4.3 Partitioning via Greedy graph coloring
We apply the Welsh-Powell largest-first greedy heuristic Welsh1967 to color $G^{(r)}\tau$ and obtain color classes $C^{(r)}{1},\dots,C^{(r)}_{m_{r}}$. Classical graph-theory results West2000 guarantee the heuristic uses no more than $\Delta+1$ colors, where $\Delta$ is the maximum vertex degree. In practice $\Delta$ is small because many task pairs do not interfere, yielding concise schedules.
4.4 Schedule Generation and Execution
We create a periodic schedule of length $m_{r}$:
$S_{t}\;=\;C^{(r)}_{\,\bigl(t\bmod m_{r}\bigr)+1},\qquad t_{r}\leq t<t_{r+1}=t_{r}+R.$ (10)
Each training step activates exactly one color class; over one period every task in that class receives a gradient update, while conflicting tasks (edges in $E_{\tau}^{(r)}$) are guaranteed not to co‑occur.
4.4.1 Minimum coverage constraint
If the greedy coloring yields a singleton class for a rarely updated task, we replicate that task into each subsequent slot until its update frequency reaches a defined minimum $f_{\min}$.
4.4.2 Warm-up and annealing
We start with $\tau=1$ (no edges, full simultaneous training) for the first $T_{\mathrm{warm}}$ steps, then logarithmically anneal $\tau$ to a target value $\tau^{*}$. This mitigates noisy gradient signals early in training.
4.5 Overall Algorithm
[…]
4.6 Relationship to Existing Techniques
4.6.1 Versus PCGrad and CAGrad
We avoid projecting or re-weighting gradients. Instead, we temporally separate high-conflict tasks via disjoint scheduling. This eliminates destructive interference without altering step directions.
4.6.2 Versus Selective-Group Updates
Instead of manually defining or tuning task subsets, our scheduler automatically partitions tasks via greedy graph coloring on the conflict graph. This guarantees at most $\Delta+1$ groups and delivers a $\tau$-dependent convergence bound (Section 5.2), all without extra heuristics or manual scheduling.
4.6.3 Versus Re-weighting Techniques
Methods like GradNorm and AdaTask compute losses and back‑propagate gradients for all $K$ tasks at every step, so both memory usage and computational cost grow with $K$. In contrast, our scheduler activates only one color class per step, so memory and compute requirements scale with the size of the active group
5 Theoretical Analysis
We provide a high-level overview of the theory supporting our scheduler-based approach. For each section, we include expanded proofs and theoretical analysis in the appendix.
5.1 Descent Preservation under $\tau$-Compatibility
Proposition 1.
For any $\tau$-compatible task set $S$
$\Bigl\|\sum_{k\in S}g_{k}\Bigr\|^{2}\;\geq\;(1-\tau)\sum_{k\in S}\|g_{k}\|^{2}.$ (12)
Proof sketch.
Write
$\|{\textstyle\sum_{k}}g_{k}\|^{2}=\sum_{k}\|g_{k}\|^{2}+2\sum_{i<j}\langle g_{i},g_{j}\rangle$ (13)
Because $S$ is $\tau$‑compatible, each inner product satisfies $\langle g_{i},g_{j}\rangle\geq-\tau\|g_{i}\|\|g_{j}\|$. Substitute and bound the cross‑terms with
$-\tau\sum_{i<j}\|g_{i}\|\|g_{j}\|\leq-\tfrac{\tau}{2}\!\sum_{i\neq j}\|g_{i}\|\|g_{j}\|$ (14)
A second application of Cauchy–Schwarz cauchy1821cours converts the mixed sum into
$\tfrac{\tau}{2}(\sum_{k}\|g_{k}\|)^{2}\leq\tau\sum_{k}\|g_{k}\|^{2}.$ (15)
Rearranging yields Equation 13 Yu2020PCGrad . ∎
The inequality introduced in Propostion 1 shows that any task group chosen by our $\tau$-aware scheduler still produces a genuine descent direction for the shared network. Essentially, separating highly conflicting tasks never turns a training step into an ascent step, so the scheduler can avoid interference without negatively impacting progress. The full proof and a clearer explanation is made available in Appendix A.
5.2 Convergence Rate with $\tau$-Dependent Constant
Theorem 1.
Assume each $\mathcal{L}_{k}$ is $L$-smooth, stochastic gradients are unbiased with variance $\leq\sigma^{2}$, and the step size is fixed to $\eta=c/\sqrt{T}$, $c\leq 1/L$. Then the scheduler satisfies
$\min_{1\leq t\leq T}\,\mathbb{E}\!\left[\|\nabla F(\theta_{t})\|^{2}\right]\;\leq\;\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\,(1+\tau)\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}.$ (16)
Proof sketch.
Start from the descent lemma for smooth non-convex objectives (Ghadimi & Lan, 2013) GhadimiLan2013 . For each step replace $\|g_{t}\|^{2}$ by the lower bound from Proposition 1 with $S=S_{t}$. Summing over $T$ steps and telescoping the function values gives
$\sum_{t=1}^{T}\eta(1-\tfrac{\tau}{2})\,\mathbb{E}\!\bigl[\|\nabla F(\theta_{t})\|^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (17)
Divide by $T\eta(1-\frac{\tau}{2})$ and optimize in $\eta$ Nesterov2004 to obtain Equation 16. The rate is $O((1+\tau)/\sqrt{T})$; when $\tau=0$ we match the classical bound.
Theorem 1 confirms that the scheduler reaches the classical $O(1/\sqrt{T})$ speed of vanilla SGD while only facing a small $(1+\tau)$ constant for conflict avoidance. Thus our scheduling strategy preserves the well‑known convergence rate even as it enforces low‑interference updates. The full proof is provided in Appendix B.
5.3 Bounded Task Update Staleness
Greedy coloring partitions the task into at most $m\leq\Delta+1$ color groups. We execute those colors cyclically, so a given color (and the task it contains) reappears at most m1m-1 intervening steps. Thus every task’s parameters are never more than m1m-1 iterations old, meeting the usual bounded‑staleness assumption Recht2011Hogwild . Thus the bounded‑staleness bound m1m-1 provides the formal guarantee that our interference‑aware scheduler never lets a task’s parameters drift for more than $\Delta$ iterations, while still ensuring Goal (that highly conflicting tasks are never updated together). Extended proof and practical explanation are available in Appendix C.
5.4 Coloring Period Bound ($\Delta$ + 1)
Because every vertex has at most $\Delta$ neighbors, when you process the vertices one by one the worst case is that all $\Delta$ of those neighbors are already colored and each is using a different color. With the palette $\{1,...,\Delta+1\}$ that still leaves at least one unused color to assign. Repeating this for every vertex ensures the algorithm never needs more than $\Delta+1$ colors in total West2000 ; Diestel2017 . This color bound guarantees that the cycle length of the scheduler is never longer than a property of the conflict graph itself, keeping both memory usage and task-refresh latency predictability small even when the task set grows large. A formal proof and extended explanation is given in Appendix D.
5.5 Baseline Non‑Convex SGD Rate
The smoothness condition lets each SGD step decrease the expected objective up to a controllable error, while the unbiased‑variance term grows only linearly with the number of steps. Balancing these two effects with a step size $\eta=\frac{c}{\sqrt{T}}$ makes the cumulative descent dominate the cumulative noise, yielding the $O\bigl(1/\sqrt{T}\bigr)$ RobbinsMonro1951 ; Nemirovski2009 bound on the smallest expected gradient norm over $T$ iterations. Thus, when $\tau=0$ (no conflicts) the scheduler reduces to ordinary multi-task SGD and recovers the same guarantee Standley2020TaskGrouping ; Ruder2017Overview ; zhang2018overview . This confirms that our methods will never slow training in the absence of interference. A full proof and in-depth explanation is provided in Appendix E.
5.6 Exact Recovery of the Population Conflict Graph and Task Partition
We show that, after observing gradients for only a modest number of steps, the scheduler can exactly reconstruct the true conflict relations among tasks by average recent gradients (EMA), computing pairwise cosines, thresholding at $\tau$, and coloring the resulting graph. Under natural assumptions about tasks being meaningfully different (separation), noise, and slowly changing gradients (drift), the conflict graph we estimate from finite data agrees, with high probability, with the ideal ’population’ conflict graph. Essentially, the graph we build from averaged gradients matches the true conflict pattern you’d get with unlimited data, meaning the grouping the scheduler uses can effectively recover the ground-truth task partition. The result explains why the scheduler’s group structure is trustworthy and ties the required probe budget (i.e., how many gradient samples must be averaged before building the graph) to interpretable quantities (noise level, margin, number of tasks). Theory supporting this can be found in Appendix F.
5.7 Descent Bounds for Scheduled versus Aggregated Updates
We compare two ways to use the same gradient information from a refresh: (1) a scheduled sequence of per-group steps (i.e., the approach we propose with the scheduler), and (2) a single aggregated step that combines all groups at once. We do this to understand when the scheduler actually helps and by how much. We find that, when different groups’ gradients pull in opposing directions (so adding them together would cancel progress) the scheduler has an advantage. In that case, taking the updates one group at a time is provably better. Our theory guarantees a larger drop in the objective during that refresh than the one-shot step, even though both use the same step size and the same gradients. Theory supporting this can be found in Appendix H.
6 Experimental Setup
6.1 Datasets
Table 1: Information on the datasets utilized in experimentation. (*Samples were removed during preprocessing)
Dataset Main Tasks (+) Aux. Tasks (-) Aux Tasks Modalities Samples
NYUv2 Semantic Segmentation [2pt] Depth Estimation [2pt] Surface Normal Prediction Color Temp. Estimation Image 250*
CIFAR-10 Image Classification Quadrant Localization [2pt] Texture Classification Corruption‐Type Prediction [2pt] Rotation Angle Prediction Image 2,500*
AV-MNIST Digit Classification Digit Parity Audio, Image 56.0k
$M$-IMDb Genre Classification Release Decade Title-Initial Classification Image, Text 25.9k
STOCKS-F&B $4\times$ Stock Return Prediction Five-Day Rolling Volatility [2pt] Sector-Average Next-Day Return Day of the Week Prediction [2pt] Lag-0 Reconstruction of Today’s Open-Price Timeseries $\times 18$ 75.5k
STOCKS-HEALTH $7\times$ Stock Return Prediction Five-Day Rolling Volatility [2pt] Sector-Average Next-Day Return Day of the Week Prediction [2pt] Lag-0 Reconstruction of Today’s Open-Price Timeseries $\times 63$ 75.5k
We evaluate the proposed scheduler alongside numerous baselines and state-of-the-art models across multiple datasets to reliably assess its general performance relative to other approaches. In total, it is evaluated across 6 datasets (see Table 1).
Across all datasets, we incorporate positive and/or negative auxiliary tasks into training. Positive auxiliary tasks share structure or predictive signals with the main tasks (e.g., common features or correlated outputs) and so can improve the learned representations by providing relevant supervision. In contrast, negative auxiliary tasks are uncorrelated or directly conflicting with the main objectives, inducing gradient interference that can slow or degrade primary performance. Including these tasks allows us to stress-test the scheduler’s ability (relative to other models) to suppress harmful tasks and prioritize helpful ones.
6.1.1 NYUv2
The NYU Depth Dataset v2 (NYUv2) Silberman2012 consists of RGB-D indoor scenes with 1,449 densely labeled pairs of RGB and depth images. To demonstrate auxiliary task value in data-scarce conditions, we employ a subset of 250 training samples randomly selected from the original training set.
We formulate a multi-main-task setup with three primary objectives: (1) semantic segmentation (14 classes), (2) depth estimation where the model predicts per-pixel depth values from RGB images, and (3) surface normal prediction where 3-channel surface normals are estimated from RGB input. The negative auxiliary task is color temperature estimation, a synthetically generated task that predicts global color temperature properties designed to interfere with the main tasks by emphasizing global color distribution rather than local semantic and geometric features.
All tasks utilize RGB images as the sole input modality, with depth maps and surface normals serving as prediction targets rather than input features. A ResNet-18 he2016residual backbone trained from scratch processes the RGB input, with task-specific decoder heads for segmentation (with 32 $\times$ upsampling), depth regression, surface normal regression, and color temperature estimation.
6.1.2 CIFAR-10
The CIFAR-10 krizhevsky2009learning dataset contains 60,000 $32\times 32$ color images across 10 generic classes. To evaluate our interference-aware scheduler in a data-scarce environment where auxiliary tasks provide maximum benefit, we employ a subset of 2,500 training samples (250 per class) from the original 50,000 training images.
For the multi-task learning setup, we set image classification as the main task and construct three auxiliary tasks synthetically from the RGB images. The positive auxiliary tasks include: (1) quadrant localization, where the model predicts which quadrant contains the primary object, and (2) texture classification using Gabor filter responses clustered into 8 texture categories via k-means clustering. The negative auxiliary tasks consist of: (3) corruption-type prediction, where images are artificially corrupted using 15 different corruption types from the ImageNet-C corruption suite hendrycks2019robustness , and (4) rotation angle prediction, where images are rotated by 0°, 90°, 180°, or 270° and the model predicts the rotation angle.
All tasks share a ResNet-18 he2016residual backbone trained from scratch without pretraining, with task-specific heads for each auxiliary task.
6.1.3 AV-MNIST
The AV‑MNIST benchmark vielzeuf2019centralnet pairs MNIST images lecun1998gradient with a log‑mel spectrogram of the corresponding spoken digit from TIDIGITS leonard1993tidigits . It is a synthetic benchmark that has significant noise applied to audio and feature reduction applied to images, making it far more difficult than the original MNIST.
We use all paired samples in our experiments. Our primary task is 10‑way digit classification. Following (Vielzeuf et al., 2018), we encode images with a small 4‑layer convolutional network and spectrograms with a 2‑layer CNN, both built and trained from scratch. These embeddings are projected and fused for processing by a simple MLP in intermediate fusion boulahia2021early ; guarrasi2025systematic , as are the models trained on MM-IMDb and STOCKS. We include only one positive auxiliary class, Digital Parity. This task aims to identify the digits as either even or odd, which has been shown to be a positive auxiliary task for improving representations on MNIST-like datasets Tacchetti2018 ; Mohammadi2020OddOneOut .
6.1.4 MM-IMDb
The MM‑IMDb dataset arevalo2017gatedmultimodalunitsinformation contains 25,959 movies with genre annotations over 23 categories. We extract poster images and plot summaries for every movie in the dataset.
The images and summaries are encoded by a frozen VGG16 simonyan2014very and Google word2vec mikolov2013efficientestimationwordrepresentations model, respectively. Our main task is movie genre prediction. We add one positive auxiliary task, Release Decade, and one negative auxiliary task, the classification of the title’s first word as either a vowel or consonant.
6.1.5 STOCKS
The STOCKS datasets we use, introduced in (Liang et al., 2021) liang2021multibench , contain stock market timeseries data across two categories. Specifically: (1) STOCKS-F&B, which has 14 input and 4 output stocks in the GICS Restaurants or Packaged Food & Meats category MSCI2024GICs , and (2) STOCKS-HEALTH, which contains 56 input and 7 output stocks in the Health Care category.
Every input stock consists of 500 trading days, with the goal of predicting returns over the next day. We discretize the continuous return variable $R$ into three non-overlapping categories: (1) $Low$, where $0\leq R<0.1$, (2) $Medium$, where $0.1\leq R<0.5$, and (3) $High$, where $R\geq 0.5$. Mean Absolute Error (MAE) is calculated by mapping the three classes to numbers ($Low\rightarrow 0$, $Medium\rightarrow 1$, $High\rightarrow 2$) and then deriving MAE as usual. Each input series is encoded by the same CNN-BiLSTM network. This consists of 3 CNNs and 1 BiLSTM Cui2018 .
We augment the main prediction task with two positive auxiliaries and two negative auxiliaries. The first positive task, Five-Day Rolling Volatility, is calculated as the standard deviation of daily logarithmic returns over a sliding five‑trading‑day window. This feature captures short-term fluctuations in a stock’s price. In Sector-Average Next-Day Return, for each date we compute the mean of the actual next‑day returns of all stocks within the same GICS sector, providing a simple measure of sector‑level momentum and drift
The negative tasks focus on useless information that is meant to distract the model. Namely, day of the week prediction (in the range of Monday to Friday) and Lag-0 Open-Price Reconstruction, which requires the model to reproduce the same day’s opening price verbatim. The first is information that contains little to no signals that would contribute to overall performance, and the second is a trivial identity mapping that contributes no real predictive challenge.
6.2 Baseline and State-of-the-Art Comparisons
We evaluate our interference-aware scheduler against 10 different multi-task learning methods across all datasets. These methods include 3 oler baseline models, more recent state-of-the-art gradient-based optimization techniques, and novel combinations of our scheduler with existing methods to demonstrate complementary benefits.
6.2.1 Baseline Models
  1. 1.
    Uniform. This baseline assigns equal weights to all tasks throughout training, representing the simplest approach where all task losses are weighted equally.
  2. 2.
    Gradnorm Chen2017GradNorm . Balances task learning rates by normalizing gradient magnitudes relative to target loss ratios. This maintains consistent training dynamics across tasks.
  3. 3.
    MGDA Sener2018 . Formulates multi-task learning as a multi-objective optimization problem, finding Pareto-optimal solutions lockwood2008pareto ; pareto2014manual through gradient descent in the convex hull of gradients FliegeSvaiter2000 ; Miettinen1999 .
6.2.2 State-of-the-Art Models
  1. 1.
    PCGrad Yu2020PCGrad . Projects conflicting gradients onto orthogonal subspaces when negative cosine similarity is detected, eliminating destructive interference between task gradients.
  2. 2.
    CAGrad Liu2021CAGrad . Extends PCGrad by adaptively adjusting gradient magnitudes based on conflict severity. This proves more nuanced modifications to gradients than binary projection.
  3. 3.
    Adatask Yang2023AdaTask . Dynamically reweighs task losses using relative loss changes, adapting to varying task learning rates during training.
  4. 4.
    FAMO Liu2023FAMO . Fast Adaptive Multitask Optimization dynamically adjusts task weights to equalize each task’s rate of loss improvement. It uses an online, per-step rule (no pairwise gradient ops), adding negligible overhead while remaining robust to loss-scale differences.
  5. 5.
    Fair Resource Allocation in MTL (FairGrad) Ban2024FairMTL . Views the shared update as a limited resource and chooses it to maximize an $\alpha$-fair utility of per-task improvements. The parameter $\alpha$ controls the trade-off between average performance and fairness.
  6. 6.
    Nash-MTL Navon2022NashMTL . Frames multitask training as a bargaining game and computes a scale-invariant weighted combination of task gradients given by the Nash bargaining solution. Weights are obtained by solving a small inner problem (e.g., via CCP) using the gradient Gram matrix. Updates are balanced across tasks.
6.2.3 Scheduler Extension Models
  1. 1.
    Scheduler AdaTask. Combines our interference-aware task selection with AdaTask’s dynamic loss weighting, applying adaptive weights only to scheduler-selected tasks.
  2. 2.
    Scheduler GradNorm Warm Start. Initializes training with GradNorm for stable gradient magnitudes, then transitions to our scheduler after 3 epochs.
  3. 3.
    Scheduler PCGrad. Applied PCGrad’s gradient projection specifically to tasks selected by our scheduler, providing fine-grained conflict resolution within $\tau$-compatible groups.
6.3 Ablation Study
6.3.1 Static One-Shot Coloring
We run the greedy graph coloring once at the start of training, freeze the resulting task groups, and never recompute the conflict graph. All other hyperparameters ($\tau$, history length $H$, and update interval $R$) match the full scheduler. As training progresses we expect the fixed coloring to grow stale, mixing tasks whose inference relationships have changed. This ablation isolates the benefit of dynamic recoloring, showing how much performance depends on adapting the schedule to evolving gradient conflicts.
6.3.2 Single-Step Conflict Estimation
Here, we set the history length to $H=1$, so every recoloring step relies on only the most recent mini-batch gradients to estimate interference. Without aggregation over many past steps, the conflict graph should become highly noisy, causing unstable task groupings from one update window to the next. This variant tests the importance of historical conflict statistics in the scheduler.
7 Results and Discussion
Table 2: Performance of Models Across Datasets
Model Accuracy (%) $\uparrow$ F&B HEALTH NYUv2
CIFAR-10 AV‑MNIST MM‑IMDb Acc. (%) $\uparrow$ MAE $\downarrow$ Acc. (%) $\uparrow$ MAE $\downarrow$ Angle Error $\downarrow$ Seg. MIOU $\uparrow$ Depth RMSE $\downarrow$
Uniform 55 63 56 45 0.57 52 0.54 21.6 0.059 0.73
GradNorm 61 65 58 47 0.57 53 0.52 21.4 0.054 0.65
MGDA 59 62 56 44 0.57 53 0.53 21.8 0.63 0.75
PCGrad 61 65 58 50 0.55 58 0.48 20.9 0.07 0.69
CAGrad 59 62 57 46 0.58 53 0.52 21.9 0.65 0.73
AdaTask 63 67 59 47 0.59 55 0.52 20.3 0.69 0.65
FAMO 64 70 61 52 0.53 60 0.49 19.9 0.074 0.63
FairGrad 62 66 59 52 0.54 60 0.47 20.7 0.072 0.67
Nash-MTL 63 66 60 52 0.54 60 0.47 20.6 0.073 0.67
Static One-Shot 61 66 58 48 0.56 54 0.51 20.5 0.071 0.65
Single-Step 40 59 20 42 0.60 47 0.55 26.4 0.042 0.81
Scheduler GradNorm 62 69 59 51 0.53 59 0.49 19.6 0.073 0.64
Scheduler AdaTask 67 71 63 52 0.53 59 0.48 20.1 0.68 0.67
Scheduler PCGrad 65 70 60 54 0.52 62 0.45 19.7 0.076 0.62
Scheduler 65 69 61 51 0.53 58 0.50 19.8 0.073 0.59
Results for all models across every experiment are depicted in Table 2. Across ten metrics on six datasets, our conflict-aware schedulers consistently match or exceed all baseline methods.
7.1 Overall Performance Improvements
Overall, the conflict-aware approaches improve over the uniform baseline by 10%-20% on CIFAR-10 and by 7% on MM-IMDb. This reinforces the idea that grouping tasks according to measured interference is more effective than treating all tasks equally at every update. On NYUv2, we see similar improvements across all the metrics. These results suggest that the scheduler’s graph coloring cleanly separates high-conflict tasks, preserving the projection or LR-balancing advantages (stemming from PCGrad’s gradient projection and AdaTask’s learning‑rate adaptation, respectively) while removing residual interference.
7.2 Ablation Study on Scheduler Design
Our ablation studies further highlight the importance of how the scheduler is designed, particularly the adaptive scheduling and history smoothing components. The Static One-Shot ablation sees a drop in performance on most metrics. This suggests that dynamic recoloring captures the evolving "conflict landscape" between different tasks. Single-step performance, on the other hand, faces dramatic performance loss across every dataset. This demonstrates that noise-free and history-averaged conflict estimates are extremely important in our scheduler.
7.3 Additional Analysis
7.3.1 Optimizer-Task Alignment
Interestingly, we observe that AdaTask-based approaches tend to be the best on classification tasks (CIFAR-10, AV-MNIST, MM-IMDb) while PCGrad-based approaches tend to be the best on tasks that model regression (NYUv2).
We believe that this stems from unique differences in the features of classification and regression-based models. For example, cross-entropy gradients near decision boundaries tend to be bursty and high in variance Shrivastava2016OHEM ; Lin2017FocalLoss ; Hoffer2017TrainLonger . By scaling each task’s step size according to its running gradient norm, AdaTask smooths out these spikes.
On the other hand, we believe that Scheduler PCGrad performs particularly well on regression and dense-prediction tasks as their tasks tend to generate smooth, large-magnitude gradients whose directions change gradually. PCGrad removes only the small component of the gradient that conflicts across tasks, preserving the main descent direction while reducing inference.
7.3.2 Synergy Between Scheduling and Baselines
We believe that the superior results found in the combinations of the scheduler and baseline models can be traced to the way scheduling and optimization reinforce one another.
First, greedy graph coloring partitions tasks into $\tau$-compatible groups, segregating tasks with highly divergent gradients. This yields a guaranteed lower bound on descent (Proposition 1), directly improving optimization efficiency.
Within each low-conflict group, the optimizer can do its job under more ideal conditions. PCGrad can remove only the remaining minor conflicting components, preserving the majority of the descent direction. AdaTask can adjust each task’s learning rate without being impacted by large, adversarial gradients.
This $\Delta+1$ color bound ensures that every task is scheduled at least once per period. This prevents tasks from being essentially starved of updates.
Finally, by computing interference over a window, the scheduler smooths out gradient fluctuations. This prevents the erratic schedule changes that projection-only grouping methods have been shown to face Yu2020PCGrad ; Shi2023Recon ; Zhang2024Proactive , thereby better stabilizing convergence.
7.4 Time Complexity and Tradeoffs
7.4.1 Time Complexity
The proposed scheduler has a time complexity of $\Theta(K^{2}d)$ per refresh. However, unlike many MTL approaches, our scheduler concentrates its extra work in occasional refreshes. This time complexity therefore becomes $\Theta\!\big(K^{2}d/R\big)$ amortized per training step where $R$ is the refresh period (the number of training steps between conflict-graph rebuilds). It adds non-trivial overhead which grows quadratically with $K$ (number of tasks) but shrinks as $R$ grows. We provide a full analysis of the time complexity in Appendix G and discuss approaches to reducing time complexity under certain conditions in Appendix G.5.
7.4.2 Speed and Tradeoffs
For smaller values of $K$, or with larger refresh period $R$, the proposed scheduler’s overhead will typically be modest relative to backpropagation and can be competitive or faster than methods that do heavy work every step. For example, methods like Nash-MTL and FairGrad typically compute all $K$ task gradients each iteration and then solve for weights, so their cost grows linearly with $K$. FAMO is almost always faster than every other approach, with a time complexity of $\Theta(1)$ per step, while still achieving impressive performance (Table 2).
These contrasts demonstrate the tradeoffs between speed and fidelity to task interference. Faster methods like FAMO minimize overhead, while methods that actually model conflicts (ours, Nash-MTL, FairGrad) can improve accuracy. These tradeoffs will have to be assessed on a case-by-case basis, based on the values that factor into each approach’s respective time complexity as well as the importance of training speed versus performance on a given application.
8 Conclusion
We have presented a conflict-aware scheduling strategy that, when combined with modern optimizers, yields provable $\tau$-dependent convergence guarantees and clear gains across diverse multi-task benchmarks. Dynamic graph coloring and history-based interference smoothing are shown to be essential through both theory and ablation. When coupled with the scheduler, PCGrad and AdaTask consistently outperform their standalone forms. This work establishes a practical, low-overhead path to more reliable and efficient multi-task training, and opens more areas for adaptive thresholding and heterogeneous task integration.
Appendix A Descent Preservation Under $\tau$-Compatibility
Proposition 2.
Let $S\subseteq\{1,\dots,K\}$ be a $\tau$-compatible task set; that is, every pair of gradients satisfies
$\bigl\langle g_{i},\,g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert,\qquad\forall\,i\neq j\in S,0\leq\tau<1$ (A.1)
Then
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}\;\;\geq\;\;(1-\tau)\sum_{k\in S}\lVert g_{k}\rVert^{2}.$ (A.2)
Proof.
We begin with the polarization identity for any finite set of vectors:
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}=\sum_{k\in S}\lVert g_{k}\rVert^{2}+2\!\!\sum_{\begin{subarray}{c}i,j\in S\\ i<j\end{subarray}}\!\bigl\langle g_{i},g_{j}\bigr\rangle.$ (A.3)
A.0.1 Lower-bounding the cross terms
Because $S$ is $\tau$-compatible, inequality (A.1) gives
$\bigl\langle g_{i},g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.4)
Insert this bound into (A.3) to obtain
$\Bigl\lVert\textstyle\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-2\tau\!\!\sum_{i<j}\!\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.5)
A.0.2 Symmetrizing the mixed sum
Observe that
$\sum_{i<j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;=\;\frac{1}{2}\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.5)
Substituting (A.5) into (A.4) yields
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-\;\tau\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (A.6)
A.0.3 Bounding the mixed sum via Cauchy-Schwarz
Apply the Cauchy–Schwarz inequality in $\mathbb{R}^{|S|}$ to the vectors $a=(\lVert g_{1}\rVert,\dots,\lVert g_{|S|}\rVert)$ and $\mathbf{1}=(1,\dots,1)$:
$\sum_{k}\lVert g_{k}\rVert=\langle a,\mathbf{1}\rangle\;\leq\;\lVert a\rVert\,\lVert\mathbf{1}\rVert=\Bigl(\sum_{k}\lVert g_{k}\rVert^{2}\Bigr)^{1/2}\,\sqrt{|S|}.$ (A.7)
Using $\left(\sum_{k}a_{k}\right)^{2}\leq|S|\sum_{k}a_{k}^{2}$ and recognizing that
$\sum_{i\neq j}\|g_{i}\|\,\|g_{j}\|\;=\;\Bigl(\sum_{k}\|g_{k}\|\Bigr)^{2}\;-\;\sum_{k}\|g_{k}\|^{2},$ (A.8)
However, because $0\leq\tau<1$ and $|S|-1\geq 1$, a looser but dimension‑free bound suffices for (A.6):
$\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\bigl(\sum_{k}\lVert g_{k}\rVert\bigr)^{2}\;\leq\;\sum_{k}\lVert g_{k}\rVert^{2}\;\cdot\;|S|\;\;\Longrightarrow\;\;\tau\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\tau\,|S|\sum_{k}\lVert g_{k}\rVert^{2}.$ (A.9)
Choosing the tight one‑step bound from (A.8) but dropping the factor $|S|-1\leq|S|$ gives
$\tau\!\!\sum_{i\neq j}\!\!\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\tau\sum_{k}\lVert g_{k}\rVert^{2}.$ (A.10)
A.0.4 Combining bounds
Insert (A.9) into (A.6):
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}-\tau\sum_{k}\lVert g_{k}\rVert^{2}\;=\;(1-\tau)\sum_{k}\lVert g_{k}\rVert^{2},$ (A.11)
establishing (A.2).
A.1 Interpretation and practical implications
Equation (A.2) guarantees that whenever we restrict an SGD step to a $\tau$-compatible group (i.e. a set of tasks whose gradients are not too antagonistic) the resulting joint update cannot annul more than a $\tau$-fraction of the potential descent. Put differently, no matter how many tasks are batched together, as long as every pair’s cosine conflict stays below $\tau$, the shared update still makes at least $(1-\tau)$ of the progress we would have obtained by training them separately and summing their squared step lengths.
This bound plays two roles in later analysis:
  1. (i)
    Descent direction safety. Because $1-\tau>0$, the aggregated gradient is never the zero vector and never flips into an ascent direction
  2. (ii)
    Convergence‑rate constant. When invoking the smooth‑SGD descent lemma, we can replace $\lVert g_{t}\rVert^{2}$ by the right‑hand side of (A.2), yielding the $(1+\tau)$ constant in our overall $O((1+\tau)/\sqrt{T})$ rate.
In practice, choosing a smaller $\tau$ makes the guarantee tighter, since each grouped mini‑batch moves almost as far as an interference‑free update. However, this leads to shorter task groups and thus fewer updates per task per epoch.
Appendix B Convergence rate with $\tau$ -dependent constant
Theorem 2 (Restatement of Theorem 1).
Let $F(\theta)\;=\;\sum_{k=1}^{K}\mathcal{L}_{k}(\theta,\phi_{k})$ be the aggregate loss, where every $\mathcal{L}_{k}$ is LL–smooth in the shared parameters $\theta$. Assume the stochastic gradient $g_{t}$ obtained at step $t$ satisfies $\mathbb{E}[g_{t}]=\nabla F(\theta_{t})$ and $\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}]\leq\sigma^{2}$. Let the step size be fixed to $\eta=\tfrac{c}{\sqrt{T}}$ with $0<c\leq\frac{1}{L}$, and suppose the scheduler selects a $\tau$-compatible task set $S_{t}$ at each step. Then
$\min_{1\leq t\leq T}\;\mathbb{E}\!\left[\lVert\nabla F(\theta_{t})\rVert^{2}\right]\;\leq\;\frac{2\bigl(F_{0}-F^{\star}\bigr)}{c\sqrt{T}}\,(1+\tau)\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}.$ (B.1)
Proof.
We proceed in four steps.
B.0.1 Smoothness descent lemma
Because each $\mathcal{L}_{k}$ is LL–smooth, $F$ is also LL–smooth. For any $\eta\leq\tfrac{1}{L}$ the standard non‑convex SGD inequality ([43], Lemma 3.2) gives
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\;\leq\;\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\frac{\eta}{2}\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (B.2)
The only term requiring modification is the squared gradient norm, because our update uses the scheduled gradient $g_{t}=\sum_{k\in S_{t}}g_{k,t}$.
B.0.2 Incorporating $\tau$‑compatibility
By Proposition 1 we have for every step $\lVert g_{t}\rVert^{2}\geq(1-\tau)\sum_{k\in S_{t}}\lVert g_{k,t}\rVert^{2}.$ Taking conditional expectation w.r.t. the mini‑batch and using $\mathbb{E}[g_{t}]=\nabla F(\theta_{t})$ yields
$\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr]\;\geq\;(1-\tau)\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr].$ (B.3)
Re‑express inequality (B.2) as
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\;\leq\;\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\frac{\eta}{2(1-\tau)}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (B.4)
B.0.3 Summation and telescoping
Summing Equation B.4 over $t=0,\dots,T-1$ (where Equation B.4 is the descent lemma after injecting the $(1-\tau)$ bound, i.e. your current B.4) and telescoping gives
$\sum_{t=0}^{T-1}\frac{\eta}{2(1-\tau)}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (B.5)
Applying Equation B.3 ($\mathbb{E}\|g_{t}\|^{2}\geq(1-\tau)\mathbb{E}\|\nabla F\|^{2}$) then yields
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (B.6)
B.0.4 Optimizing the constant step size
Substitute the schedule $\eta=\tfrac{c}{\sqrt{T}}$ and $1/(1-\tau)\leq 1+\tau$:
$\min_{t}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\,(1+\tau)+\frac{cL\sigma^{2}}{\sqrt{T}},$ (B.7)
which is exactly (B.1). ∎
B.1 Discussion and intuition
Equation (B.1) extends the classical $O(1/\sqrt{T})$ rate for non‑convex SGD to the scheduled multi‑task setting. The multiplicative factor $(1+\tau)$ quantifies the loss of descent efficiency incurred by allowing up to $\tau$-level conflict inside each scheduled group. When $\tau=0$ (perfect alignment) the bound matches the Ghadimi–Lan constant; as $\tau\to 1$ the constant doubles, mirroring the worst‑case scenario where half of the descent power can be cancelled by conflicting gradients.
Appendix C Bounded Staleness via Greedy Graph Coloring
Proposition 3 (Staleness Bound).
Let $G=(\mathcal{T},E)$ be the task–conflict graph whose vertices are tasks and whose edges connect pairs with interference coefficient exceeding the threshold $\tau$. Denote by $\Delta$ its maximum degree. Greedy graph coloring produces a proper coloring $C_{1},\dots,C_{m}$ with
$m\;\leq\;\Delta+1.$ (C.1)
If the scheduler activates the color classes in the cyclic order $C_{1}\!\to\!C_{2}\!\to\!\dots\!\to\!C_{m}\!\to\!C_{1}\!\to\!\dots$, then every task is updated at least once every
$s_{\max}\;=\;m-1\;\leq\;\Delta$ (C.2)
iterations. Hence the scheduler satisfies the bounded‑staleness condition of Recht et al. ([45]): the parameter vector used by any task is never older than $\Delta$ updates.
Proof.
We proceed in two parts.
Part A: Color count bound.
The greedy (Welsh–Powell) algorithm scans the vertices in non‑increasing order of degree and assigns to each vertex the smallest available color that is not used by its earlier colored neighbours. When the ii‑th vertex is reached, at most $\Delta$ of its neighbours have been colored, so at most $\Delta$ colors are unavailable. Therefore one of the first $\Delta+1$ colors is always free, implying $m\leq\Delta+1$ ([78].
Part B: Staleness of cyclic execution.
Fix any task $T\in\mathcal{T}$ and let it belong to color $C_{j}$ for some $1\leq j\leq m$. Under cyclic scheduling, $C_{j}$ is executed at steps $t=j,\;j+m,\;j+2m,\dots$. The number of intervening steps between two consecutive executions of $C_{j}$ is exactly m1m-1. Hence task $T$ never waits more than $s_{\max}=m-1$ iterations for an update. Combining with Equation C.1 yields $s_{\max}\leq\Delta$.
C.1 Interpretation
The staleness bound (equation C.2) guarantees that the shared parameters used by any task cannot lag behind the most recent update by more than $\Delta$ iterations, even in the worst case where the conflict graph is a clique of size $\Delta+1$. This criterion matches the bounded‑delay assumption commonly used to analyse asynchronous SGD and lock‑free training, ensuring that the convergence proofs derived under that assumption apply unchanged to our scheduled setting. In practice $\Delta$ is often much smaller than the total number of tasks, so the scheduler achieves low interference and low parameter staleness simultaneously.
Appendix D Greedy Graph-Coloring Uses at Most $\Delta\!+\!1$ Colors
Proposition 4 (Coloring Period Bound).
Let $G=(V,E)$ be a finite, simple, undirected graph with maximum degree $\Delta\,\coloneqq\!\max_{v\in V}\deg(v)$. The greedy (Welsh–Powell) coloring algorithm [^1] Order the vertices in non–increasing degree and assign to each the smallest positive integer (color) not used by its previously colored neighbors. produces a proper vertex coloring with no more than
$\chi_{\text{greedy}}(G)\;\leq\;\Delta+1$ (D.1)
distinct colors. Consequently, when the scheduler activates the color classes in a cyclic order, the cycle length is bounded by $\Delta+1$—a quantity depending only on the structure of the conflict graph.
Proof.
Let the vertices be processed in the Welsh–Powell order $v_{1},v_{2},\dots,v_{|V|}\,$. Assume inductively that after coloring the first k1k-1 vertices the algorithm has used at most $\Delta+1$ colors. Consider vertex $v_{k}$. By construction, every neighbor of $v_{k}$ has degree at most $\Delta$, and at most $\Delta$ of those neighbors appear before $v_{k}$ in the ordering. Hence, at the moment of coloring $v_{k}$, at most $\Delta$ colors are forbidden (one for each previously colored neighbor). Because the palette $\{1,2,\dots,\Delta+1\}$ contains $\Delta+1$ colors in total, there is always at least one color still available. Assigning the smallest such color to $v_{k}$ maintains a proper coloring and never introduces a new color beyond $\Delta+1$.
Proceeding vertex‑by‑vertex, no step ever requires more than $\Delta+1$ colors, establishing Equation D.1. ∎
D.1 Implications for the scheduler
A coloring with at most $\Delta+1$ classes means the scheduler’s cycle period (the number of batches needed before every task reappears) is bounded by a graph invariant independent of the number of tasks. Even if thousands of tasks exist, as long as each one conflicts with at most $\Delta$ others, the memory footprint (one shared backbone plus $\Delta+1$ sets of head activations) and the maximum waiting time between successive updates for any task (bounded by $\Delta$; see Proposition 3) remain predictable and small. This guarantee is essential for scaling the scheduler to large, heterogeneous tasks while retaining bounded‑delay assumptions needed in standard convergence proofs of asynchronous and scheduled optimization algorithms.
Appendix E Baseline Non‑Convex SGD Convergence Rate
Theorem 3 (Classical $O(1/\sqrt{T})$ bound).
Let $F:\mathbb{R}^{d}\!\to\!\mathbb{R}$ be an $L$-smooth, possibly non‑convex objective and suppose the stochastic gradient $g_{t}$ computed at iteration $t$ satisfies
$\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t}),\qquad\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}\mid\theta_{t}]\;\leq\;\sigma^{2}.$ (E.1)
Run SGD with the constant step size $\eta=\tfrac{c}{\sqrt{T}},\;0<c\leq\frac{1}{L},$ for $T$ iterations starting from $\theta_{0}$. Then
$\min_{0\leq t<T}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;\frac{2\bigl(F_{0}-F^{\ast}\bigr)}{c\sqrt{T}}+\frac{cL\sigma^{2}}{\sqrt{T}},$ (E.2)
where $F^{\ast}=\inf_{\theta}F(\theta)$.
Proof.
The proof is a streamlined restatement of ([43, 48]). By $L$-smoothness,
$F(\theta_{t+1})\;\leq\;F(\theta_{t})+\langle\nabla F(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\lVert\theta_{t+1}-\theta_{t}\rVert^{2}.$ (E.3)
With $\theta_{t+1}=\theta_{t}-\eta\,g_{t}$ and taking conditional expectation,
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\leq\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\eta\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L}{2}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr].$ (E.4)
Decompose the squared stochastic gradient: $\mathbb{E}[\lVert g_{t}\rVert^{2}]=\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}]\leq\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\sigma^{2}.$ Thus
$\mathbb{E}[F(\theta_{t+1})]\leq\mathbb{E}[F(\theta_{t})]-\frac{\eta}{2}\,\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (E.5)
Summing from $t=0$ to T1T-1 and telescoping gives
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]\leq F_{0}-F^{\ast}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (E.6)
Dividing by $\eta T$ and inserting $\eta=c/\sqrt{T}$ yields (E.2). ∎
E.1 Connection to the scheduler
When the interference threshold is set to $\tau=0$, every pair of tasks is deemed incompatible and the scheduler activates exactly one color‑class per step. However, under a deterministic (cyclic) activation order, $g_{t}=\sum_{k\in S_{t}}g_{k,t}$ is not an unbiased estimator of the full gradient $\sum_{k=1}^{K}\nabla\mathcal{L}_{k}$, because
$\mathbb{E}\bigl[g_{t}\mid\theta_{t}\bigr]=\sum_{k\in S_{t}}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t})\;\neq\;\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t}).$ (E.7)
To recover the classical $O(1/\sqrt{T})$ rate in this case (which is applicable to the schedular), we need to do one of the following:
  • Randomized scheduling: Sample one of the $m$ color‑classes uniformly at random each step. Then
    $\mathbb{E}\bigl[g_{t}\mid\theta_{t}\bigr]=\frac{1}{m}\sum_{j=1}^{m}\sum_{k\in C_{j}}\nabla\mathcal{L}_{k}=\sum_{k=1}^{K}\nabla\mathcal{L}_{k},$ (E.8)
    so the usual unbiased‑SGD analysis (Theorem 3) applies directly.
  • Deterministic cyclic analysis: Keep the fixed periodic schedule but invoke convergence results for cyclic block‑coordinate descent in non‑convex smooth optimization (e.g. Saha & Tewari 2013; Cai et al. 2022). These guarantee an $O(1/\sqrt{T})$ rate up to constants depending on the number of blocks, by tracking the bias within each cycle and aggregating descent over a full sweep.
In either case, when $\tau=0$ the method recovers an $O(1/\sqrt{T})$ convergence guarantee, matching classical non‑convex SGD under appropriate scheduling.
Appendix F Exact Recovery of the Population Conflict Graph and Task Partition
F.1 Setting, definitions, and population objects
Let $K\geq 2$ be the number of tasks and $d\geq 1$ the parameter dimension. At designated refresh iterations, the scheduler:
  1. (i)
    computes a per-task exponential moving average (EMA) of stochastic gradients over a probe window of $R$ iterations,
  2. (ii)
    forms a cosine-similarity matrix from the $K$ EMA vectors,
  3. (iii)
    builds a conflict graph by thresholding negative cosines at a fixed level $\tau$ with $\tau\in(0,1)$,
  4. (iv)
    computes a proper coloring of the conflict graph, and
  5. (v)
    schedules one color class per iteration until the next refresh
Definition F.1.
At the beginning of a refresh window (i.e., at a fixed iterate $\theta$), let
$\mu_{i}\;\in\;\mathbb{R}^{d}\qquad(i=1,\dots,K)$ (F.1)
denote the population task gradients (or the window-stationary means). Define the population cosine matrix $C^{\star}\in[-1,1]^{K\times K}$ by
$C^{\star}_{ij}\;=\;\frac{\langle\mu_{i},\mu_{j}\rangle}{\|\mu_{i}\|\,\|\mu_{j}\|},\qquad i\neq j,\quad C^{\star}_{ii}=1.$ (F.2)
Definition F.2.
Fix $\tau\in(0,1)$. The population conflict graph $G^{\star}=(V,E^{\star})$ on vertex set $V=\{1,\dots,K\}$ has an edge $\{i,j\}$ iff $C^{\star}_{ij}<-\tau$. The true grouping $\mathcal{P}^{\star}$ is one of:
  1. (A)
    Component Model: the vertex partition given by the connected components of $G^{\star}$.
  2. (B)
    Multipartite model: a partition $V=\bigsqcup_{r=1}^{m}P_{r}$ (with $m\geq!$) such that $G^{\star}$ is the complete $m$-partite graph induced by $\{P_{r}\}_{r=1}^{m}$ (no edges within any $P_{r}$, all cross-part edges present)
When we later speak of group recovery, we mean equality of the empirical partition (defined from data) with $\mathcal{P}^{\star}$, up to label permutation in case (B).
F.2 Assumptions
We adopt the following assumptions, which are standard in analyses of stochastic-gradient methods and verifiable in practice (see, e.g., [79, 80, 81, 82, 83], for concentration of geometrically weighted and mixing sequences, see [84, 85]).
Assumption 1 (Separation margin around the threshold).
There exists $\gamma\in(0,1-\tau)$ such that for all $i\neq j$:
[…]
Assumption 2 (Probe noise model and EMA).
In the refresh window of length $R$, the per-iteration stochastic task gradients admit the decomposition
$g_{i,t}\;=\;\mu_{i}+\xi_{i,t},\qquad t=1,\dots,R,$ (F.4)
where $\{\xi_{i,t}\}_{t=1}^{R}$ are mean-zero, sub-Gaussian with parameter $\sigma^{2}$, and satisfy a $\phi$-mixing or martingale-difference condition ensuring concentration with geometric weights. The EMA for task $i$ is
$\tilde{g}_{i}\;=\;\sum_{t=1}^{R}w_{t}\,g_{i,t},\qquad w_{t}\;=\;\frac{(1-\beta)\,\beta^{R-t}}{1-\beta^{R}},\qquad\beta\in[0,1).$ (F.5)
Define the effective sample size $n_{\rm eff}$ by
$n_{\rm eff}^{-1}\;:=\;\sum_{t=1}^{R}w_{t}^{2}\quad\text{so that}\quad c_{1}\,\frac{1-\beta^{2}}{1-\beta^{2R}}\;\leq\;\frac{1}{n_{\rm eff}}\;\leq\;c_{2}\,\frac{1-\beta^{2}}{1-\beta^{2R}}$ (F.6)
for absolute constants $0<c_{1}\leq c_{2}<\infty$ (in particular $n_{\rm eff}=\Theta\!\big(\tfrac{1-\beta^{2R}}{1-\beta^{2}}\big)$).
Assumption 3 (Slow drift within a refresh).
Over the refresh window, the changes in $\mu_{i}$ are small enough to be absorbed in the concentration bounds below (equivalently, one can regard $\mu_{i}$ as constant within the window by working at the start-of-window iterate and moving any drift into the noise process).
Assumption 4 (Minimum norm and task inclusion).
There exists $m_{0}>0$ such that $\|\mu_{i}\|\geq m_{0}$ for all tasks included in the graph. In our implementation, we make it so that tasks with $\|\tilde{g}_{i}\|<\nu$ (for a small $\nu\ll m_{0}$) are temporarily excluded from graph construction until stabilized.
Assumption 5 (Threshold selection).
The threshold $\tau$ is fixed across refreshes or selected using data independent of the probe window used to form $\{\tilde{g}_{i}\}$ (e.g., via a separate pilot set). The analysis below treats $\tau$ as deterministic with respect to the probe sample.
F.3 Deterministic group recovery from the conflict graph
We begin with basic graph-thoeretic facts that we will use once we have established that the empirical conflict graph coincides with its population counterpart.
Proposition 5 (Chromatic number of a complete multipartite graph).
If $G^{\star}$ is complete $m$-partite with parts $\{P_{r}\}_{r=1}^{m}$, then $\chi(G^{\star})=m$.
Proof.
Picking one vertex from each part yields a clique of size $m$, hence $\chi(G^{\star})\geq m$. Coloring each part with a distinct color is proper, hence $\chi(G^{\star})\leq m$. Therefore $\chi(G^{\star})=m$. ∎
Theorem 4 (Identifiability via optimal coloring under model (B)).
Assume model (B), i.e., $G^{\star}$ is complete $m$-partite with parts $\{P_{r}\}_{r=1}^{m}$. Let $c:V\to\{1,\dots,m\}$ be a proper coloring of $G^{\star}$ that uses exactly $\chi(G^{\star})$ colors. Then each color class equals some part $P_{r}$ (up to relabeling).
Proof.
In a complete multipartite graph, any two vertices from different parts are adjacent. Thus, no color class can contain vertices from two different parts, so each color class is contained in some $P_{r}$. By Proposition 5, $\chi(G^{\star})=m$, so any optimal coloring uses exactly $m$ colors. Since there are $m$ nonempty parts, none can be split across two colors. Hence, the color classes coincide with $\{P_{r}\}_{r=1}^{m}$ up to permutation. ∎
Proposition 6 (Identifiability via components under model (A)).
Under model (A), the grouping $\mathcal{P}^{\star}$ equals the connected components of $G^{\star}$. Consequently, any procedure that returns the connected components of the empirical graph recovers $\mathcal{P}^{\star}$ whenever the empirical graph equals $G^{\star}$.
F.4 Uniform control of empirical cosines from EMA gradients
We now quantify the deviation of the empirical cosine matrix $\widehat{C}$ formed from $\{\tilde{g}_{i}\}$ relative to $C^{\star}$.
Lemma 1 (EMA vector concentration in directions of interest).
Assume Assumption 2 and Assumption 3. There exists a constant $c>0$ depending only on the mixing parameters such that for any fixed unit vector $u\in\mathbb{S}^{d-1}$ and any $\varepsilon>0$.
$\Pr\!\Big(\,\big|\langle\tilde{g}_{i}-\mu_{i},\ u\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (F.7)
In particular, for any finite set of unit vectors $\{u_{j}\}_{j=1}^{M}$, a union bound yields
$\Pr\!\Big(\,\max_{1\leq j\leq M}\big|\langle\tilde{g}_{i}-\mu_{i},\ u_{j}\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\,M\,\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (F.8)
Proof.
The scalar process $\{\langle\xi_{i,t},u\rangle\}_{t=1}^{R}$ is sub-Gaussian with variance proxy $\sigma^{2}$ and satisfies the same mixing condition. Exponential-weighted averages of such sequences obey Hoeffding-Azuma/Berstein-type tail bounds with variance proxy $\sigma^{2}\sum_{t}w_{t}^{2}=\sigma^{2}/n_{\rm eff}$. The stated inequality follows. ∎
Lemma 2 (Cosine stability under perturbations).
Assume Assumption 4 and let $\epsilon>0$. If for a pair $(i,j)$ we have
$\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,$ (F.9)
then
$\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \frac{6\,\epsilon}{m_{0}}\,+\,\frac{4\,\epsilon^{2}}{m_{0}^{2}}.$ (F.10)
Proof.
Write $\tilde{g}_{i}=\mu_{i}+\delta_{i}$, $\tilde{g}_{j}=\mu_{j}+\delta_{j}$. Decompose the numerator and denominator in the cosine:
$\langle\tilde{g}_{i},\tilde{g}_{j}\rangle-\langle\mu_{i},\mu_{j}\rangle=\langle\delta_{i},\mu_{j}\rangle+\langle\mu_{i},\delta_{j}\rangle+\langle\delta_{i},\delta_{j}\rangle,$ (F.11)
and
$\|\tilde{g}_{i}\|=\|\mu_{i}\|\sqrt{1+2\langle\delta_{i},\mu_{i}\rangle/\|\mu_{i}\|^{2}+\|\delta_{i}\|^{2}/\|\mu_{i}\|^{2}}$ (F.12)
Using Assumption 4,
$|\langle\delta_{i},\mu_{j}/\|\mu_{j}\|\rangle|\leq\epsilon$ (F.13)
and
$|\langle\delta_{i},\mu_{i}/\|\mu_{i}\|\rangle|\leq\epsilon$ (F.14)
imply
$|\langle\delta_{i},\mu_{j}\rangle|\leq\epsilon\|\mu_{j}\|$ (F.15)
and
$|\langle\delta_{i},\mu_{i}\rangle|\leq\epsilon\|\mu_{i}\|$ (F.16)
A second-order expansion of the cosine in $(\delta_{i},\delta_{j})$ with the above controls yields the bound. The constants $6$ and $4$ arise from collecting the linear and quadratic contributions in $\epsilon/m_{0}$. ∎
Combining Lemma 1 and Lemma 2 with a union bound over all unordered pairs $(i,j)$ shows that the empirical cosines are uniformly close to their population counterparts.
Proposition 7 (Uniform cosine accuracy with high probability).
Assume Assumption 2, Assumption 3, and Assumption 4. For any $\epsilon>0$ there exist absolute constants $c,C>0$ such that if
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\epsilon^{2}}\,\log\!\Big(\frac{K}{\delta}\Big)$ (F.17)
then, with probability $1-\delta$,
$\max_{i<j}\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \epsilon$ (F.18)
Proof.
For each unordered pair $(i,j)$, apply Lemma 1 with the four unit vectors $\mu_{j}/\|\mu_{j}\|$, $\mu_{i}/\|\mu_{i}\|$, and use Lemma 2 to convert these directional deviations into a cosine deviation bound. A union bound over the $O(K^{2})$ pairs yields the claimed logarithmic factor. The constants absorb the quadratic term in $\epsilon$ by requiring $\epsilon\leq m_{0}$. ∎
F.5 Exact edge recovery and group recovery
We first show that a uniform cosine error smaller than the margin $\gamma$ implies exact equality of empirical and population conflict graphs.
Theorem 5 (Exact conflict-graph recovery under the margin).
Assume Assumptions 15. If
$\max_{i<j}|\widehat{C}_{ij}-C^{\star}_{ij}|\leq\epsilon$ (F.19)
with $\epsilon<\gamma$
$\max_{i<j}|\widehat{C}_{ij}-C^{\star}_{ij}|\leq\epsilon\qquad\text{with}\qquad\epsilon<\gamma$ (F.20)
then
$\widehat{C}_{ij}\geq-(\tau-\gamma)-\epsilon>-\tau$ (F.21)
Proof.
For any pair $(i,j)$, if $C^{\star}_{ij}\leq-(\tau+\gamma)$, then $\widehat{C}_{ij}\leq-(\tau+\gamma)+\epsilon<-\tau$, hence $\{i,j\}\in\widehat{E}$. If $C^{\star}_{ij}\geq-(\tau-\gamma)$, then $\widehat{C}_{ij}\geq-(\tau-\gamma)-\epsilon>-\tau$, hence $\{i,j\}\notin\widehat{E}$. ∎
Combining Proposition 7 and Theorem 5 yields a high-probability statement.
Corollary 1 (High-probability exact recovery of $G^{\star}$).
Under Assumptions 15, if
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Big(\frac{K}{\delta}\Big)$ (F.22)
then $\Pr(\widehat{G}=G^{\star})\geq 1-\delta$.
We now translate exact edge recovery into group recovery.
Theorem 6 (Group recovery under the component model).
Under model (A) and the conditions of Corollary 1, with probability at least $1-\delta$, the connected components of $\widehat{G}$ equal $\mathcal{P}^{\star}$.
Proof.
Immediate from $\widehat{G}=G^{\star}$ and the definition of $\mathcal{P}^{\star}$. ∎
Theorem 7 (Group recovery under the multipartite model).
Under model (B) and the conditions of Corollary 1, with probability at least $1-\delta$, $\chi(\widehat{G})=m$ and any optimal coloring of $\widehat{G}$ yields color classes equal to $\{P_{r}\}_{r=1}^{m}$ up to label permutation.
Proof.
If $\widehat{G}=G^{\star}$, then $\widehat{G}$ is complete $m$-partite. Proposition 5 gives $\chi(\widehat{G})=m$. Theorem 4 implies identifiability up to permutation by any optimal coloring. ∎
F.6 Quantitative probe-budget requirement
Combining the bounds above yields the following sample-complexity statement.
Corollary 2.
Under assumptions 15, there exist absolute constants $c,C>0$ such that the following holds. If the EMA parameters $(R,\beta)$ are chosen to ensure
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Big(\frac{K}{\delta}\Big)\qquad\Big(\text{equivalently,}\quad\sum_{t=1}^{R}w_{t}^{2}\ \leq\ c\,\frac{m_{0}^{2}\,\gamma^{2}}{\sigma^{2}}\,\frac{1}{\log(K/\delta)}\Big)$ (F.23)
then $\Pr(\widehat{G}=G^{\star})\geq 1-\delta$, and consequently Theorems 67 apply. In particular, at fixed $\beta$ and for large $R$, $n_{\rm eff}=\Theta((1-\beta^{2})^{-1})$, recovering the usual $O(\log K)$ dependence on the number of tasks.
F.7 Summary of the recovery argument
We summarize the logical flow leading to consistency of the scheduler.
  1. (i)
    Assumptions: Assumptions 15 define the conditions in which in which across-group population cosines lie below $(\tau+\gamma)$, within-group cosines lie above $(\tau-\gamma)$, EMA gradients concentrate with effective sample size $n_{\mathrm{eff}}$, and all included tasks have non-negligible gradient norm.
  2. (ii)
    Uniform cosine accuracy: Lemmas 12 together with Proposition 7 yield a high-probability uniform cosine approximation:
    $\max_{i<j}\bigl|\widehat{C}_{ij}-C^{\star}_{ij}\bigr|\ \leq\ \epsilon,$ (F.24)
    with probability at least $1-\delta$, where $\epsilon$ decreases as $n_{\mathrm{eff}}$ increases.
  3. (iii)
    Exact recovery of edges: If the approximation tolerance satisfies $\epsilon<\gamma$, Theorem 5 converts the uniform bound into exact edge recovery of the conflict graph:
    $\widehat{G}\;=\;G^{\star}.$
  4. (iv)
    Recovery of the grouping: Given $\widehat{G}=G^{\star}$, Theorem 6 implies group recovery under the component model (groups are the connected components). Under the multipartite model, Proposition 5 and Theorem 4 yield $\chi(\widehat{G})=m$ and Theorem 7 shows that any optimal coloring returns the true parts (up to label permutation).
Quantitative consequence.
Assume Assumptions 15 and fix $\delta\in(0,1)$. Let $m_{0}=\min_{i}\|\mu_{i}\|$ and let $\sigma^{2}$ be the variance proxy from Assumption 2. If the EMA probe budget satisfies
$n_{\mathrm{eff}}\;\geq\;C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Bigl(\frac{K^{2}}{\delta}\Bigr)$ (F.25)
for a universal constant $C>0$, then with probability at least $1-\delta$ the empirical conflict graph equals the population graph: $\widehat{G}=G^{\star}$. Consequently:
  1. (i)
    under the component model (A), the connected components of $\widehat{G}$ coincide with $\mathcal{P}^{\star}$.
  2. (ii)
    under the multipartite model (B), $\chi(\widehat{G})=m$ and any optimal coloring of $\widehat{G}$ recovers $\mathcal{P}^{\star}$ up to permutation of labels.
Appendix G Computational Complexity of One Refresh (and Amortized Over Training)
We analyze the computational and memory complexity of the proposed interference-aware scheduler per refresh and its amortized cost over training. The former accounts for the cost of a single refresh operation while the latter represents the average cost distributed across all training steps. We distinguish the work required by the underlying mutli-task training objective (.e.g, backpropagatoin to obtain gradients) from the scheduler overhead (EMA maintenance, cosine computation, conflict graph construction, and color).
G.1 Notation
  • $K\in\mathbb{N}$number of tasks
  • $d\in\mathbb{N}$dimension of the gradient EMA vector per task
  • $R\in\mathbb{N}$refresh period (number of training steps between graph rebuilds)
  • $\beta\in[0,1)$exponential moving average (EMA) parameter
  • $T\in\mathbb{N}$total number of training steps
  • $\mathsf{G}>0$time to compute one backward pass to obtain a task gradient at a refresh
  • $\tau\in(0,1)$ – conflict threshold; an undirected edge $\{i,j\}$ is present iff $\widehat{C}_{ij}<-\tau$
  • $T_{\text{refresh}}>0$time cost of a single scheduler refresh
  • $S_{\text{refresh}}>0$peak additional memory used during a refresh
  • $N_{\text{refresh}}\in\mathbb{N}$number of refreshes over $T$ steps with period $R$ (satisfies $N_{\text{refresh}}\in\{\lfloor T/R\rfloor,\lceil T/R\rceil\}$ and $N_{\text{refresh}}\leq T/R+1$)
G.2 Per-refresh complexity (time and space)
At a refresh, the scheduler performs a finite sequence of deterministic operations on the current collection of task-wise exponential moving averages (EMAs) of gradients. Let
$M\in\mathbb{R}^{K\times d}$ (G.1)
denote the matrix whose $i$-th row $m_{i}^{\top}$ is the EMA for task $i$. A refresh first updates these rows through a scalar EMA rule
$m_{i}\leftarrow\beta m_{i}+(1-\beta)g_{i}$ (G.2)
using the most recent probe (or reused) gradient $g_{i}$. It then constructs the cosine-similarity matrix
$\widehat{C}=\widetilde{M}\widetilde{M}^{\top}$ (G.3)
where $\widetilde{M}$ is the row-normalized version of $M$. It thresholds $\widehat{C}$ at $\tau$ to obtain the conflict adjacency. Finally, it applies a graph-coloring routine to the resulting simple graph ([40]).
EMA maintenance uses a constant number of vector operations per task: one multiply-add on each of the $d$ coordinates of $m_{i}$. Aggregating over all $K$ tasks gives a time proportional to $Kd$. The storage required to hold all EMAs is the $K\times d$ array $M$, so the working set devoted to EMAs is $\Theta(Kd)$ numbers.
The construction of $\widehat{C}$ proceeds by normalizing each row of $M$ and then multiplying $\widetilde{M}$ by its transpose. Row normalization touches each entry exactly once and therefore costs $\Theta(Kd)$ time. The Gram product $\widetilde{M}\widetilde{M}^{\top}$ consists of $K^{2}$ dot products of length $d$, which is $\Theta(K^{2}d)$ time ([86]). The cosine matrix itself occupies $K^{2}$ entries. If it is retained after thresholding, it uses $\Theta(K^{2})$ space. If dropped right after graph construction, that $\Theta(K^{2})$ storage is only temporary.
Thresholding linearly scans the off-diagonal of $\widehat{C}$, adding an undirected edge when $\widehat{C}_{ij}<-\tau$; this costs $\Theta(K^{2})$ time. The result is either a dense $K\times K$ boolean array requiring $\Theta(K^{2})$ space, or a sparse adjacency whose size depends on the number of conflicts (e.g., $\Theta(kK)$ when retaining the $k$ most negative entries per row).
Putting these pieces together yields the following statement.
Proposition 8 (Per-refresh scheduler overhead).
Under the standard RAM model with dense matrix multiplication costed as $\Theta(K^{2}d)$, the time required by a single scheduler refresh is
$T_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}d)\;+\;\Theta(K^{2})\;+\;O(K^{2})\;=\;\Theta(K^{2}d),$ (G.4)
and the additional space required by the scheduler during the refresh is
$S_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}),$ (G.5)
where the $\Theta(K^{2})$ term is transient if $\widehat{C}$ is not retained after coloring.
Proof.
The EMA update costs $\Theta(Kd)$ by a direct count of coordinate-wise multiply-adds. Row normalization also costs $\Theta(Kd)$. The Gram matrix requires $K^{2}$ inner products of length $d$, which is $\Theta(K^{2}d)$. This term dominates $\Theta(Kd)$. Thresholding scans $O(K^{2})$ entries and is therefore $\Theta(K^{2})$. The greedy coloring performs a sort of $K$ keys and then assigns at most one color per edge incident on the current vertex, which is $O(K^{2})$ in the worst case. This is dominated by $\Theta(K^{2}d)$ whenever $d\geq 1$. Summing these contributions and absorbing lower-order terms yields $T_{\text{refresh}}=\Theta(K^{2}d)$. The EMA matrix occupies $\Theta(Kd)$ memory, and storing $\widehat{C}$ uses $\Theta(K^{2})$. But if $\widehat{C}$ is discarded immediately after thresholding, only $\Theta(Kd)$ remains. ∎
G.3 Amortized cost over training
Let $R\in\mathbb{N}$ denote the refresh period as the scheduler executes a refresh once every $R$ training steps. Consider a training run of length $T$ steps. The number of refreshes executed is $\lfloor T/R\rfloor$ or $\lceil T/R\rceil$ depending on whether one refresh occurs at step 0. In either case it is bounded by $T/R+1$. Multiplying the per-refresh time $T_{\text{refresh}}$ by the number of refreshes and dividing by $T$ shows that the amortized scheduler time per training step satisfies
$\frac{1}{T}\,N_{\text{refresh}}\,T_{\text{refresh}}\;\leq\;\frac{1}{T}\,\Big(\frac{T}{R}+1\Big)\,T_{\text{refresh}}\;=\;\frac{1}{R}\,T_{\text{refresh}}+\frac{1}{T}\,T_{\text{refresh}}$ (G.6)
Letting $T\to\infty$ (or simply taking $T$ large compared to one refresh) eliminates the $T^{-1}T_{\text{refresh}}$ boundary term, yielding the asymptotic amortized bound
$\frac{1}{R}\,T_{\text{refresh}}\;=\;\ \frac{1}{R}\,\Theta(K^{2}d)\$(G.7)
If probe gradients are computed only at refreshes, their contribution $K\mathsf{G}$ per refresh adds $\frac{1}{R}\Theta(K\mathsf{G})$ to the amortized time per step. If, instead, the training loop already computes task-wise gradients each step and these are reused to update the EMAs, then the probe term is absent and the amortized scheduler overhead remains $\frac{1}{R}\Theta(K^{2}d)$.
The amortized space usage is simpler. The EMA matrix $M$ must be retained throughout training and therefore contributes $\Theta(Kd)$ at all times. The cosine matrix $\widehat{C}$ and the adjacency are constructed only during the refresh. They’re released after coloring, so the $\Theta(K^{2})$ space does not persist. Consequently, the persistent memory overhead attributable to the scheduler is $\Theta(Kd)$, while the peak overhead during a refresh is $\Theta(Kd)+\Theta(K^{2})$.
G.4 Conditions for negligible overhead
Let the amortized per-step costs be
$C_{\text{sched}}=\frac{a}{R}\,K^{2}d\qquad\text{and}\qquad C_{\text{probe}}=\frac{b}{R}\,K\,\mathsf{G},$
where $a,b>0$ are platform-dependent constants and $\mathsf{G}$ denotes the per-task backpropagation cost of the optional probe at a refresh. For fixed $R$,
$\frac{C_{\text{sched}}}{C_{\text{probe}}}=\frac{a}{b}\,\frac{K^{2}d}{K\,\mathsf{G}}=\frac{a}{b}\,\frac{Kd}{\mathsf{G}}.$
Hence $C_{\text{sched}}$ is negligible relative to $C_{\text{probe}}$ whenever
$\frac{C_{\text{sched}}}{C_{\text{probe}}}\leq\varepsilon\quad\text{for some }0<\varepsilon\ll 1\quad\Longleftrightarrow\quad Kd\;\leq\;\frac{b}{a}\,\varepsilon\,\mathsf{G}.$
G.5 Reducing time complexity
In this section, we detail approaches that can be taken under certain circumstances to optimize time complexity.
G.5.1 Random projections
We replace the EMA matrix $M\in\mathbb{R}^{K\times d}$ by a lower-dimensional sketch $\widetilde{M}=MR$ with $R\in\mathbb{R}^{d\times r}$ and $r\ll d$ ([87]). The sketching multiply costs $O(Kdr)$ and the cosine Gram becomes $O(K^{2}r)$ instead of $\Theta(K^{2}d)$. Storage for the sketched EMAs is $O(Kr)$. By the Johnson-Lendnstrauss (JL) random projection guarantee, if we map the $K$ task-EMA vectors from $\mathbb{R}^{d}$ to $\mathbb{R}^{r}$ using a suitable random matrix with $r=\Theta(\epsilon^{-2}\log K)$, then after row normalization all pairwise inner products (hence cosines) are preserved within $\pm\epsilon$ with high probability. We assume a uniform row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ (which can be enforced in practice by skipping tasks with $\|m_{i}\|<\nu\ll m_{0}$) so cosine errors remain controlled. Choosing $\epsilon<\gamma$, where $\gamma$ is the cosine margin from the recovery analysis, ensures that every pair remains on the same side of the threshold $\tau$. Therefore the set $\{(i,j):\widehat{C}_{ij}<-\tau\}$ and the resulting coloring are unchanged with high probability.
In short, dimensionality drops from $d$ to $r$, the refresh cost drops from $\Theta(K^{2}d)$ to $O(Kdr+K^{2}r)$, and decisions are preserved as long as the chosen $r$ makes the embedding error smaller than the margin.
G.5.2 Deterministic covariance sketching via frequent directions
We maintain a deterministic sketch $B\in\mathbb{R}^{\ell\times d}$ of the row space of $M$ using Frequent Directions and either project rows onto $\mathrm{span}(B)$ or form an approximate Gram from the sketch ([88, 89]). Maintaining the sketch costs $O(Kd\ell)$, the cosine Gram in the sketch space costs $O(K^{2}\ell)$, and storage for the sketch is $O(\ell d)$. Frequent Directions gives a spectral-norm bound
$|MM^{\top}-\widehat{MM^{\top}}|_{2}\leq\epsilon|M|_{F}^{2}$ (G.8)
when $\ell=\Theta(\epsilon^{-2})$, which yields a uniform bound on inner-product and squared-norm errors. Assuming a row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ and applying a standard cosine perturbation bound after row normalization, one obtains
$\big|\cos(m_{i},m_{j})-\widehat{\cos}(m_{i},m_{j})\big|\;\leq\;\frac{2\,\epsilon\,\|M\|_{F}^{2}}{m_{0}^{2}}\;+\;O\!\left(\frac{\epsilon^{2}\,\|M\|_{F}^{4}}{m_{0}^{4}}\right)\$(G.9)
Taking $\epsilon$ small enough so that the right-hand side is $\gamma$ ensures that all threshold decisions and the resulting coloring are preserved deterministically. Thus the effective dimension drops from $d$ to $\ell$ in the worst case, and the refresh cost becomes $O(Kd\ell+K^{2}\ell)$.
G.5.3 Edge sampling for conflict graphs with adaptive refinement
We reduce the number of cosine evaluations by computing $\widehat{C}_{ij}$ for only $\tilde{O}(K\log K)$ randomly chosen task pairs to build a provisional conflict graph and then refining by evaluating additional pairs that are near the threshold or needed to certify connectivity and chromatic structure. We still compute all $K$ row norms once in $O(Kd)$ time for normalization, and the first pass costs $O(Kd\log K)$ for the sampled dot products. The total cost adds only the refinement work, which remains small when only few pairs are ambiguous. Under a planted separation model with margin $\gamma$ and reasonably dense cross-group conflicts, one can show with high probability that the sampled graph already captures the correct inter-group connectivity, so the coloring or component structure is recovered after the first pass and only boundary pairs need refinement. This reduces the pairwise work from $K^{2}$ to near $K\log K$ while preserving the final decisions under stated assumptions ([90]).
G.5.4 Incremental gram updates
We avoid rebuilding the full cosine matrix when only a small subset of tasks has meaningfully changed since the last refresh. If $s$ rows of $M$ cross a chosen change threshold, we first renormalize these rows and then recompute both the corresponding $s$ rows and $s$ columns of the Gram by taking dot products against all $K$ rows, which costs $O(sKd)$, with an additional $O(sd)$ to update norms, instead of $\Theta(K^{2}d)$, and we leave all unchanged entries as they are. This update is exact for the affected entries, so conflict edges and coloring decisions are preserved by construction, and the reduction is deterministic whenever $s\ll K$. To prevent slow drift in the unchanged entries, we can periodically force a full rebuild and reset the change counters.
Appendix H Descent Bounds for Scheduled versus Aggregated Updates
We compare two update procedures over a single refresh: a scheduled sequence of per-group steps (i.e., the approach we propose in our paper) and a single aggregated step that combines all groups at once. Both use the same step size $\eta$ and the same gradient information measured at the start of the refresh, and our analysis operates at the level of $L$-smooth (descent) upper bounds. We identify when the scheduled bound is strictly tighter and summarize implications under PL / strong convexity.
Throughout, $F:\mathbb{R}^{d}\to\mathbb{R}$ is differentiable and $L$-smooth, i.e.
$F(y)\ \leq\ F(x)+\langle\nabla F(x),y-x\rangle+\tfrac{L}{2}|y-x|^{2},\quad\forall x,y.$ (H.1)
We write $\nabla F(x)=\sum_{r=1}^{m}G_{r}(x)$, where each $G_{r}(x)$ is the group gradient for color $r$ (any fixed linear aggregator of task gradients assigned to color $r$ for the current refresh). We use a refresh step size $\eta\in(0,\,1/L]$.
H.1 Single refresh baselines and notation
H.1.1 Single aggregated step
Definition H.1 (Aggregated step).
Starting from the same point $x$, with step size $\eta\in(0,1/L]$ and group gradients $G_{r}^{0}:=G_{r}(x)$ (with $\nabla F(x)=\sum_{r=1}^{m}G_{r}^{0}$), define
$x^{\mathrm{agg}}\ :=\ x-\eta\,\sum_{r=1}^{m}G_{r}^{0}.$ (H.2)
One-shot $L$-smoothness bound.
Applying $L$-smoothness with $y=x^{\mathrm{agg}}$ yields
$F(x^{\mathrm{agg}})\ \leq\ F(x)\;-\;\eta\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\tfrac{L\eta^{2}}{2}\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}.$ (H.3)
H.1.2 Scheduled group sequence over one refresh
Definition H.2 (Scheduled refresh).
Starting from the same point $x$, define
$x_{0}:=x,\qquad x_{r}:=x_{r-1}-\eta\,G_{r}(x_{r-1})\quad(r=1,\ldots,m),\qquad x^{\mathrm{sch}}:=x_{m}.$ (H.4)
Order and notation.
The within refresh order $(1,\ldots,m)$ may be fixed or randomly permuted each refresh. We write $H(\cdot)$ for the Hessian of $F$ and take $\eta\in(0,1/L]$.
Our goal is to compare upper bounds derived from $L$-smoothness for $F(x^{\text{sch}})$ and $F(x^{\text{agg}})$.
H.2 Telescoping bound for scheduled updates
Lemma 3 (Smoothness Expansion for Two Scheduled Groups).
Let $m=2$ and $G_{r}^{0}:=G_{r}(x)$. For any $\eta\in(0,1/L]$,
$F(x^{\text{sch}})$ $\leq F(x)\;-\;\eta\,\langle\nabla F(x),\,G_{1}^{0}\rangle\;+\;\tfrac{L\eta^{2}}{2}\,\|G_{1}^{0}\|^{2}$ (H.5)
$\quad-\;\eta\,\langle\nabla F(x),\,G_{2}(x_{1})\rangle\;+\;\tfrac{L\eta^{2}}{2}\,\|G_{2}(x_{1})\|^{2}\;+\;\eta^{2}\!\int_{0}^{1}\!\!\big\langle H(x{-}t\eta G_{1}^{0})\,G_{1}^{0},\ G_{2}(x_{1})\big\rangle dt.$
Proof sketch.
Apply the $L$-smoothness inequality at the first step to bound $F(x_{1})$. For the second step, use $L$-smoothness at $x_{1}$ and expand
$\nabla F(x_{1})=\nabla F(x)-\int_{0}^{1}H(x{-}t\eta G_{1}^{0})\,\eta G_{1}^{0}\,dt$ (H.6)
by the fundamental theorem of calculus along the segment $x\to x_{1}$. ∎
H.2.1 Start-of-refresh reduction under per-group lipschitzness
We adopt the following assumption whenever we compare bounds solely in terms of start-of-refresh measurements. It will be used throughout Sections H.3H.6
Assumption 6 (Per-group lipschitzness).
Each group map $G_{r}(\cdot)$ is $L_{r}$-lipschitz:
$\|G_{r}(u)-G_{r}(v)\|\ \leq\ L_{r}\,\|u-v\|\quad\text{for all }u,v.$ (H.7)
Under this assumption, for $m=2$ we have $G_{2}(x_{1})=G_{2}^{0}+\delta_{2}$ with $\|\delta_{2}\|\leq L_{2}\eta\|G_{1}^{0}\|$, hence
$\|G_{2}(x_{1})\|\leq\|G_{2}^{0}\|+L_{2}\eta\|G_{1}^{0}\|$ (H.8)
For general $m$
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\,\eta\sum_{p<r}\|G_{p}^{0}\|\qquad(r=2,\dots,m)$ (H.9)
When these substitutions are made in scheduled bounds, the induced drift contributions are collected into a nonnegative penalty $R_{m}(x;\eta)$
H.3 Upper bounds for scheduled and aggregated updates (general $m$)
Applying $L$-smoothness $m$ times yields the scheduled upper bound
$\mathrm{UB}_{\text{sch}}(x;\eta)$ $=F(x)\;-\;\eta\,\sum_{r=1}^{m}\big\langle\nabla F(x),\,G_{r}(x_{r-1})\big\rangle\;+\;\frac{L\eta^{2}}{2}\,\sum_{r=1}^{m}\|G_{r}(x_{r-1})\|^{2}$ (H.10)
$\quad+\;\eta^{2}\sum_{1\leq p<q\leq m}\ \int_{0}^{1}\!\big\langle H\!\big(x-t\eta G_{p}(x_{p-1})\big)\,G_{p}(x_{p-1}),\ G_{q}(x_{q-1})\big\rangle\,dt.$
The aggregated upper bound is the one-shot bound from Equation H.3, restated as
$\mathrm{UB}_{\text{agg}}(x;\eta):=F(x)\;-\;\eta\,\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\frac{L\eta^{2}}{2}\,\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}$ (H.11)
The integrals in Equation H.10 are over ordered pairs $p<q$ along the specific sequence $x_{0}\to x_{1}\to\cdots\to x_{m}$; the bound therefore depends on the within-refresh order. Randomizing the order yields an expected version.
In Sections H.4H.6 we express the scheduled bound in terms of $\{G_{r}^{0}\}$ under the per-group lipschitzness assumption. The associated drift terms are aggregated into $R_{m}(x;\eta)$.
H.4 Scheduled and aggregated gap at a common linearization
Define the shorthand
$I_{pq}(x;\eta)\ :=\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle\,dt$ (H.12)
By expanding $\mathrm{UB}_{\text{sch}}$ around $\{G_{r}^{0}\}$ and collecting the lipschitz drift penalties into $R_{m}(x;\eta)\geq 0$, we obtain:
Theorem 8 (Upper-bound gap under per-group lipschitzness).
Assuming per-group lipschitzness, for any partition $\{G_{r}\}$ and $\eta\in(0,1/L]$,
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\;\leq\;\eta^{2}\!\sum_{1\leq p<q\leq m}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;+\;I_{pq}(x;\eta)\Big)\;+\;R_{m}(x;\eta).$ (H.13)
Using $\|H(\cdot)\|_{\mathrm{op}}\leq L$ and Cauchy-Schwarz
$I_{pq}(x;\eta)\ \leq\ L\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|$ (H.14)
which gives the envelope
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ L\eta^{2}\!\sum_{p<q}\!\big(\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|-\langle G_{p}^{0},G_{q}^{0}\rangle\big)\;+\;R_{m}(x;\eta)\ \ \geq\ 0$ (H.15)
Interpretation
This shows that without additional structure, the scheduled smoothness bound can be looser than the aggregated bound. The gap is governed by Hessian-weighted cross terms $I_{pq}$
Proposition 9 (Drift penalty bound under per-group lipschitzness).
Assume each group map $G_{r}$ is $L_{r}$-lipschitz. Then for $r\geq 2$,
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\eta\sum_{p<r}\|G_{p}^{0}\|\ :=\ \|G_{r}^{0}\|+L_{r}\eta\,S_{r-1},$
and the scheduled start substitution error satisfies
$R_{m}(x;\eta)\ \leq$ $\ \eta^{2}\Big(\sum_{p=1}^{m}\|G_{p}^{0}\|\Big)\,\sum_{r=2}^{m}L_{r}\,S_{r-1}$
$\ +\ \frac{L\eta^{2}}{2}\sum_{r=2}^{m}\Big(2\,\|G_{r}^{0}\|\,L_{r}\eta S_{r-1}\;+\;(L_{r}\eta S_{r-1})^{2}\Big),$
so $R_{m}(x;\eta)=O(\eta^{2})$ with constants controlled by $\{L_{r}\}$ and $\{\|G_{r}^{0}\|\}$.
H.5 Sufficient conditions for a tighter scheduled bound
The terms $I_{pq}(x;\eta)$ encode Hessian-weighted interactions between groups and determine when scheduling is advantageous at the bound level.
Assumption 7 (Hessian-weighted negative cross-terms).
There exist nonnegative margins $\{\Gamma_{pq}\}_{p<q}$ such that
$I_{pq}(x;\eta)\ =\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle dt\ \leq\ -\,\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\quad\text{for all }p<q$ (H.16)
Theorem 9 (Strict upper-bound improvement under per-group lipschitzness and negative Hessian-weighted cross-terms).
Assuming per-group lipschitzness and H.16, for any $\eta\in(0,1/L]$,
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ \eta^{2}\!\sum_{p<q}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;-\;\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\Big)\ +\ R_{m}(x;\eta)$ (H.17)
In particular, if
$\sum_{p<q}\!\Big(\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|+L\,\langle G_{p}^{0},G_{q}^{0}\rangle\Big)\ >\ \frac{R_{m}(x;\eta)}{\eta^{2}}$ (H.18)
then $\mathrm{UB}_{\text{sch}}(x;\eta)\ <\ \mathrm{UB}_{\text{agg}}(x;\eta)$
H.6 PL or strong convexity: standard rate and upper-bound gains for scheduling
Assume $F$ satisfies the Polyak–Łojasiewicz (PL) inequality with parameter $\mu>0$:
$\tfrac{1}{2}|\nabla F(x)|^{2}\ \geq\ \mu\,(F(x)-F^{\star}),\qquad\forall x$ (H.19)
For any $\eta\in(0,1/L]$, the single aggregated update satisfies the standard GD bound
$F(x^{\text{agg}})\ \leq\ F(x)\;-\;\eta\Big(1-\tfrac{L\eta}{2}\Big)\,|\nabla F(x)|^{2}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)$ (H.20)
Define the upper-bound gain (under per-group lipschitzness, so both bounds are expressed at start-of-refresh):
$\Delta{\text{UB}}(x;\eta)\ :=\ \mathrm{UB}_{\text{agg}}(x;\eta)-\mathrm{UB}_{\text{sch}}(x;\eta)\ \geq\ 0$ (H.21)
whenever H.18 holds. Since $F(x^{\text{sch}})\leq\mathrm{UB}_{\text{sch}}(x;\eta)$ and $\mathrm{UB}_{\text{agg}}(x;\eta)$ upper-bounds the one-shot decrease term in H.20, we obtain the bound-level contraction
$F(x^{\text{sch}})-F^{\star}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)\;-\;\Delta_{\text{UB}}(x;\eta).$ (H.22)
Consequently, under per-group lipschitzness and H.18, the scheduled refresh satisfies the standard gradient-descent contraction and, in addition, achieves an extra nonnegative decrement $\Delta_{\mathrm{UB}}(x;\eta)$ in the upper bound.
H.7 Why the assumptions are mild
The assumptions we use are mild. They are standard and naturally align with our training pipeline.
H.7.1 $L$-smoothness
This is the same regularity used throughout the main paper and in our baselines. Each task loss we optimize is $L_{i}$-smooth, so the overall objective is $L$-smooth. We only use this to apply the standard smoothness (descent) inequality [44, 91].
H.7.2 Per-group lipschitzness of $G_{r}$
Each $G_{r}$ is a fixed linear combination of the task gradients assigned to group $r$. If each task gradient is $L_{i}$-lipschitz, then $G_{r}$ is lipschitz with constant $Lr\leq\sum{i\in r}L_{i}$. In other words, this property falls out of task-level smoothness. The same smoothness estimates we already use for step-size selection upper-bound the $L_{r}$.
H.7.3 Negative Hessian-weighted cross-terms
The condition we use asks that, over the short moves we actually take ($\eta\leq 1/L$), groups that are separated by the scheduler continue to exhibit negative interaction under the local Hessian (i.e., the Hessian-weighted cross-terms remain negative). This aligns with how the scheduler is built. It separates tasks that exhibit sustained negative interactions and it periodically refreshes assignments so the local geometry does not drift far. Thus the assumption matches the mechanism we deploy.
H.7.4 PL and strong convexity
We invoke PL only to convert a per-refresh decrease into a standard contraction factor. We do not require global strong convexity. A local PL inequality around the iterates is enough, which is commonly observed after warm-up and annealing we already use ([92, 93, 94]).
H.8 Concluding remarks
This section formalized a bound-level comparison between scheduled and aggregated updates. Without additional structure the scheduled bound need not be tighter, but under per-group lipschitzness and negative Hessian-weighted cross-terms it becomes strictly tighter, and under PL the scheduled refresh inherits the standard GD contraction with an additional nonnegative decrement. In practice, these conditions arise naturally once the task-group assignments stabilize, so the scheduler will typically achieve tighter descent bounds without changing step sizes or gradient information.