### v2

# Graph Coloring for Multi-Task Learning

Abstract
When different objectives conflict with each other in multi-task learning, gradients begin to interfere and slow convergence, thereby potentially reducing the final model’s performance. To address this, we introduce SON-GOKU, a scheduler that computes gradient interference, constructs an interference graph, and then applies greedy graph-coloring to partition tasks into groups that align well with each other. At each training step, only one group (color class) of tasks are activated, and the grouping partition is constantly recomputed as task relationships evolve throughout training. By ensuring that each mini-batch contains only tasks that pull the model in the same direction, our method improves the effectiveness of any underlying multi-task learning optimizer without additional tuning. Since tasks within these groups will update in compatible directions, multi-task learning will improve model performance rather than impede it. Empirical results on six different datasets show that this interference-aware graph-coloring approach consistently outperforms baselines and state-of-the-art multi-task optimizers. We provide extensive theory showing why grouping and sequential updates improve multi-task learning, with guarantees on descent, convergence, and the ability to accurately identify what tasks conflict or align.
1 Introduction
Multi-task learning (MTL) trains a single model to solve several tasks simultaneously, sharing knowledge across them to learn more effectively (Caruana, 1997). This allows models to generalize better and converge faster. However, a key issue known as negative transfer arises when tasks don’t align very well with each other (Sener & Koltun, 2018; Shi et al., 2023). When two tasks push the shared network in different directions their gradients clash, slowing or even reversing learning. Prior work addresses this issue primarily via (1) gradient manipulation, which reshapes task gradients to reduce conflicts, and (2) loss reweighting, which rescales task objectives to balance their influence. While effective in some specific settings, these strategies typically treat conflict locally at the level of shared-parameter updates and often overlook the evolving global structure of interactions among tasks throughout training.
Some recent works focus on partitioning tasks into subsets (groups) and updating those groups separately. These approaches have been found to improve accuracy and training stability by forming groups with high measured affinity and then updating one group at a time (Fifty et al., 2021; Jeong & Yoon, 2025). Grouping can outperform gradient manipulation and loss reweighting when tasks form clusters with aligned gradients, because each update then reduces direct clashes in the shared layers, lowers gradient variance within the step, and lets compatible tasks reinforce one another while conflicting tasks wait for their turn.
However, grouping methods often face a few key limitations: (1) many rely on dense pairwise affinities that grow noisy and costly as the number of tasks rises (Fifty et al., 2021; Standley et al., 2020; Sherif et al., 2024), and (2) others predetermine or rarely update groups, so they drift as task relations change (Wang et al., 2024; Ruder, 2017), and (3) several use local heuristics that fail to enforce global compatibility or to specify how groups should rotate over time (Zhang & Yang, 2018; Malhotra et al., 2022).
We present SON-GOKU (Scheduling via Optimal INterference-aware Graph-COloring for TasK Grouping in MUltitask Learning). We measure gradient interference, build a graph of tasks from those measurements, greedily color the graph to form non-conflicting compatible task groups, and update one color group per step during training. This design addresses the earlier issues. We estimate the interference graph from lightweight minibatch statistics and keep it sparse, which avoids noisy dense matrices and scales to many tasks. We recolor the graph at regular intervals so the groups track changing relations during training. Greedy graph coloring ensures we update only compatible tasks in each step, and the color order gives a simple way to cycle through the groups. Our proposed scheduler does not have to work in isolation, it can function on top of existing loss-reweighting and gradient-manipulation MTL approaches.
In our theoretical analysis (Section 5) we show that, under standard conditions, SON-GOKU tends to group tasks whose gradients are, on average, aligned within each group, with high probability. We further show that, over a refresh window, sequentially updating these low-conflict groups yields at least as much expected descent as a single mixed update, and strictly more when between-group interference is sufficiently negative. We also prove that SON-GOKU preserves descent and reaches the usual non-convex SGD rate under mild assumptions, with only a small factor that depends on the within-group conflict level. In Appendix D we discuss the scheduler’s amortized time complexity and the tradeoffs it offers between speed and performance. We discuss ways in which practitioners can reduce its time complexity under certain conditions.
Empirical results from experiments demonstrate that SON-GOKU consistently improves outcomes compared to other MTL approaches, especially when SON-GOKU is coupled with existing approaches. Our contributions are as follows:
  • We propose SON-GOKU, an interference-aware scheduler that measures cross-task gradient conflict, builds a conflict graph, colors it to form compatible groups, and activates one group per step. It can be used on top of standard MTL optimizers.
  • We provide theoretical analysis that offers guarantees on SON-GOKU’s grouping, convergence, scheduling behavior, and more.
  • Across six datasets, SON-GOKU improves over strong baselines and pairs well with methods like PCGrad, AdaTask, and GradNorm, delivering consistent gains.
  • We perform an ablation study showing that dynamic recoloring and history-averaged conflict estimates are key contributors to performance.
2 Related Work
Many MTL methods (especially earlier ones) adjust task influence by learning or adapting loss weights. Examples include uncertainty-based scaling (Kendall et al., 2018), rate-based schemes such as DWA (Liu et al., 2019), and fast bilevel formulations like FAMO (Liu et al., 2023). FAMO in particular is notable for its $\mathcal{O}(1)$ per-step time complexity. These approaches keep all tasks active each step while modulating relative magnitudes. A completely different approach, which emerged in 2018 with MGDA (Sener & Koltun, 2018), focuses on updating shared-parameter update directions to mitigate interference. Methods like PCGrad (Yu et al., 2020), CAGrad (Liu et al., 2021), and MGDA (Sener & Koltun, 2018) modify the geometry of the shared update to reduce cross-task conflicts while still updating all tasks each step. A smaller body of work forms subsets of tasks to update together, using offline affinity estimation or training-dynamics signals (Fifty et al., 2021; Standley et al., 2020; Wang et al., 2024; Sherif et al., 2024). Most recently, Selective Task Group Updates proposes online grouping with sequential updates, reporting that update order can influence task-specific learning (Jeong & Yoon, 2025). SON-GOKU differs in mechanism from existing approaches (Section 4). It complements loss reweighting and gradient surgery, and we provide explicit guarantees on descent, convergence, and graph/partition recovery. An expanded discussion of more related work is provided in Appendix M.
3 Problem Setup
We formalize multi-task learning (MTL) (Caruana, 1997) as optimizing a shared network while activating only a subset of tasks at each step. Each task contributes a loss whose gradients may align or conflict. We quantify conflict using (the negative of) cosine similarity, embed tasks in a conflict graph, and later use that graph to derive a schedule. This section fixes notation and states the optimization goal that the proposed approach addresses.
3.1 Data and Notation
Let $\mathcal{T}=\{T_{1},\dots,T_{K}\}$ be the set of tasks. The model has shared parameters $\theta\in\mathbb{R}^{d}$ and task-specific parameters $\phi_{k}\in\mathbb{R}^{d_{k}}$ for $T_{k}$. Each task draws examples $(x,y_{k})$ from a distribution $\mathcal{D}_{k}$ and defines a per-example loss $\ell_{k}(\theta,\phi_{k};x,y_{k})$. Its population loss is
$L_{k}(\theta,\phi_{k}):=\mathbb{E}_{(x,y_{k})\sim\mathcal{D}_{k}}\big[\ell_{k}(\theta,\phi_{k};x,y_{k})\big].$ (1)
We minimize the standard weighted MTL objective
$F(\theta,\phi_{1},\dots,\phi_{K})\;=\;\sum_{k=1}^{K}w_{k}\,L_{k}(\theta,\phi_{k}),$ (2)
with nonnegative task weights $w_{k}$ (default $w_{k}=1$). Note that, for simplicity in later sections, we absorb $w_{k}$ into the per-task gradient estimates. This is permissible since positive scalings do not change cosine signs or the induced conflict graph.
At step $t$, for any task $k$ that is active we compute stochastic gradients on a mini-batch $\mathcal{B}_{k}^{(t)}\subset\mathcal{D}_{k}$:
$g_{k}^{(t)}:=\nabla_{\theta}L_{k}(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}),\qquad h_{k}^{(t)}:=\nabla_{\phi_{k}}L_{k}(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}).$ (3)
In our proposed method, we form exponential moving averages (EMA) of per-task gradients within a refresh window to stabilize cosine estimates so that they do not become stale (Sec. 4).
3.1.1 Interference Coefficient
We quantify pairwise interaction with the interference coefficient
$\rho_{ij}\;=\;-\frac{\langle\tilde{g}_{i},\tilde{g}_{j}\rangle}{\|\tilde{g}_{i}\|\,\|\tilde{g}_{j}\|},$ (4)
where $\tilde{g}_{i}$ and $\tilde{g}_{j}$ are the EMA-smoothed gradients at refresh. Positive $\rho_{ij}$ indicates conflict (negative cosine). $\rho_{ij}\leq 0$ indicates alignment or neutrality.
3.1.2 Conflict Graph
Fix a tolerance $\tau\in(0,1)$. The conflict graph is
$G_{\tau}=(\mathcal{T},E_{\tau}),\qquad E_{\tau}=\bigl\{(i,j):\rho_{ij}>\tau\bigr\}.$ (5)
Vertices are tasks. An edge between a pair means to not update that pair together. We will utilize $G_{\tau}$ for coloring and scheduling in Section 4
3.2 Goal
At training step $t$ we choose an active set $S_{t}\subseteq\mathcal{T}$ and update only those tasks:
$\theta_{t+1}=\theta_{t}-\eta_{t}\sum_{k\in S_{t}}g_{k}^{(t)},\qquad\phi_{k,t+1}=\begin{cases}\phi_{k,t}-\eta_{t}h_{k}^{(t)},&k\in S_{t},\\ \phi_{k,t},&k\notin S_{t}.\end{cases}$ (6)
The problem the scheduler addresses is to design the sequence $\{S_{t}\}_{t=1}^{T}$ so that: (1) every task is visited regularly; and (2) conflicting tasks seldom appear together. We instantiate this via greedy graph coloring in Section 4 and analyze the guarantees in Section 5.
4 Proposed Approach
Figure 1: Interference-aware scheduling pipeline: (a) For each task $T_{i}$ (circles $T_{1}\ldots T_{6}$), we smooth recent per-step gradients with an Exponential Moving Average (EMA); (b) From these EMA vectors we compute the pairwise cosine matrix. In the figure, cells outlined with red dashes mark pairs with cosine $\tau$. These are flagged as conflicts; (c) We build the conflict graph whose nodes are tasks $T_{i}$ and whose red dashed edges connect exactly those pairs identified in (b); (d) We apply greedy graph coloring so that no conflict edge lies within a color, producing low-conflict groups. In the example shown, we have two groups: A as blue and B as orange; (e) During training we activate one group per step. After every $R$ steps (here, $R=4$) we ’refresh’ and run the pipeline again from step A, where we update the EMAs with the latest gradients.
We design an interference-aware scheduler that partitions tasks into low-conflict groups and activates exactly one group per optimization step. The procedure consists of four stages: (1) estimating pairwise interference, (2) building and coloring the conflict graph, (3) generating a periodic schedule, and (4) updating that schedule as training evolves. An overview of the scheduler is provided as Algorithm 1 in Appendix A. A visualization of SON-GOKU is provided in Figure 1 alongside a simple summary in the Figure caption.
4.1 Estimating Gradient Interference
We absorb task weights into per-task losses, so $g_{k}^{(t)}$ is the gradient of the weighted loss $w_{k}L_{k}$. Cosine calculations and graph construction are not impacted by applying positive scaling.
At step $t$ and for every task $T_{k}$ appearing in the current mini-batch we compute a task‑specific stochastic gradient
$g_{k}^{(t)}=\nabla_{\theta}\mathcal{L}_{k}\!\bigl(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}\bigr),$ (7)
using an independent sub-batch $\mathcal{B}_{k}^{(t)}\subset\mathcal{D}_{k}$. We then update an exponential moving average
$\tilde{g}_{k}^{(t)}=\beta\,\tilde{g}_{k}^{(t-1)}+(1-\beta)\,g_{k}^{(t)},\qquad\beta\in[0,1),$ (8)
which stabilizes cosine estimates while requiring only two buffers per task (current and previous). Whenever we refresh the schedule (every $R$ steps) we form the pairwise interference matrix
$\rho_{ij}^{(t)}\;=\;-\frac{\langle\tilde{g}_{i}^{(t)},\tilde{g}_{j}^{(t)}\rangle}{\|\tilde{g}_{i}^{(t)}\|\,\|\tilde{g}_{j}^{(t)}\|},\qquad i,j\in\{1,\dots,K\}.$ (9)
Computing all K(K1)/2K(K-1)/2 cosines is $O(K^{2}d)$ with $d$ representing the shared‑parameter dimension. We also write $h_{k}^{(t)}=\nabla_{\phi_{k}}\mathcal{L}_{k}\!\bigl(\theta_{t},\phi_{k,t};\mathcal{B}_{k}^{(t)}\bigr)$ for the gradient with respect to the task-specific parameters $\phi_{k}$.
4.2 Conflict Graph Construction
Given a tolerance $\tau\in(0,1)$, the conflict graph at update round $r$ is
$G^{(r)}_{\tau}=(V,E_{\tau}^{(r)}),\quad V=\{1,\dots,K\}E_{\tau}^{(r)}=\bigl\{(i,j)\;:\;\rho^{(t_{r})}_{ij}>\tau\bigr\}.$ (10)
To clarify, tasks are indexed by integers $1\ldots K$ in Equation 10. Edges connect tasks whose averaged gradients have cosine similarity less than $\tau$. Intuitively, larger $\tau$ yields a sparser conflict graph, typically fewer colors (larger per-step groups), and more frequent updates per task. Smaller $\tau$ results in a denser graph, more colors (smaller per-step groups), and less frequent updates per task.
4.3 Partitioning via Greedy Graph Coloring
We apply the Welsh-Powell largest-first greedy heuristic (Welsh & Powell, 1967) to color $G^{(r)}_{\tau}$ and obtain color classes $C^{(r)}_{1},\dots,C^{(r)}_{m_{r}}$. Classical graph-theory results (West, 2000; Diestel, 2017) guarantee the heuristic uses no more than $\Delta+1$ colors, where $\Delta$ is the maximum vertex degree. In practice $\Delta$ is small because many task pairs do not interfere, yielding concise schedules.
4.4 Schedule Generation and Execution
We create a periodic schedule of length $m_{r}$:
$S_{t}\;=\;C^{(r)}_{\,\bigl(t\bmod m_{r}\bigr)+1},\qquad t_{r}\leq t<t_{r+1}=t_{r}+R.$ (11)
Each training step activates exactly one color class; over one period every task in that class receives a gradient update, while conflicting tasks (edges in $E_{\tau}^{(r)}$) are guaranteed not to co-occur.
4.4.1 Minimum update frequency
If the greedy coloring yields a singleton class for a rarely updated task, we increase its update frequency by duplicating it only into steps whose active color has no conflict edge to that task.
4.4.2 Warm-up and Annealing
We start with $\tau=1$ (no edges, full simultaneous training) for the first $T_{\mathrm{warm}}$ steps, then logarithmically anneal $\tau$ to a target value $\tau^{*}$. This mitigates noisy gradient signals early in training.
4.5 Time Complexity
The proposed scheduler has a time complexity of $\Theta(K^{2}d)$ per refresh. However, unlike many MTL approaches, our scheduler concentrates its extra work in occasional refreshes. This time complexity therefore becomes $\Theta\!\big(K^{2}d/R\big)$ amortized per training step where $R$ s the refresh period (the number of training steps between conflict-graph rebuilds). It adds non-trivial overhead which grows quadratically with $K$ (number of tasks) but shrinks as $R$ grows. We provide a full analysis of the time complexity in Appendix D and discuss approaches to reducing time complexity under certain conditions in Appendix D.5.
5 Theoretical Analysis
We discuss some of the main guarantees behind SON-GOKU. For a very brief overview: (1) Updating groups of tasks whose gradients are mostly low-conflict (no internal edges) reduces the objective on average and still achieves the usual $1/\sqrt{T}$ convergence rate; (2) Over a refresh window, scheduling several group updates can beat one mixed update that uses all tasks at once; and (3) With a small number of recent gradient measurements per task (via EMA) and a margin separating conflicts, the estimated conflict graph matches the ideal one, giving a short schedule where every task is updated at least once every $\Delta+1$ steps ($\Delta$ is the maximum number of conflicts for any task). We provide expanded assumptions, definitions, proofs, reasoning, analysis, etc. in Appendix 5.4I.
5.1 Descent Preservation Within a Low-conflict Group
If the active set $S_{t}$ at step $t$ is $\tau$-compatible, then the combined update is a descent direction with a quantitative lower bound:
$\Bigl\|\sum_{k\in S_{t}}g_{k,t}\Bigr\|^{2}\;\geq\;\Bigl(1-\tau\,(|S_{t}|-1)\Bigr)\sum_{k\in S_{t}}\|g_{k,t}\|^{2}$ (12)
Thus the step cannot flip to ascent whenever $\tau(|S_{t}|-1)<1$. This is proved by expanding the polarization identity and controlling cross terms under the $\tau$-compatibility condition (see Appendix E). Essentially, this means that SON-GOKU’s per-step updates are safe when groups are low conflict. The aggregate direction keeps pointing downhill and the cancellation is quantitatively limited by $\tau$ and group size.
5.2 Nonconvex Convergence at the Standard Rate up to a Small Factor
Under standard smoothness and noise conditions (see Appendix I) and with steps $\eta=c/\sqrt{T}$, SON-GOKU achieves the usual nonconvex SGD rate, with a mild $(1+\tau)$ factor that reflects within-group conflict:
$\min_{t<T}\mathbb{E}\,\|\nabla F(\theta_{t})\|^{2}\;\leq\;\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\,(1+\tau)\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}$ (13)
When $\tau=0$, the constant matches the classical bound (Bottou et al., 2018; Ghadimi & Lan, 2013); as $\tau\to 1$, it at most doubles, matching the intuition that conflict can cancel up to half of the progress. This demonstrates that scheduling does not degrade asymptotic progress. SON-GOKU preserves the $1/\sqrt{T}$ decay of the gradient norm while controlling the constant through the compatibility threshold $\tau$. In other words, we keep the standard rate of SGD and trade a small constant for reduced interference.
5.3 When Scheduled Groups Outperform a Single Mixed Update
We compare two ways to use the same gradients gathered at a refresh: a scheduled sequence of per-group steps (i.e., the scheduler used in SON-GOKU) versus a single aggregated step. Using a telescoping $L$-smooth bound and evaluating both trajectories at a common linearization—i.e., expanding $F$ at the refresh start $\theta_{t_{r}}$ and applying the same first-order model with the same step size—the scheduled bound is never worse and is strictly better when cross-group interaction terms are sufficiently negative (so mixed updates would cancel progress).
Essentially, when different groups’ gradients pull in opposing directions (so adding them together would cancel progress) the scheduler has an advantage. In that case, taking the updates one group at a time is provably better. Our theory guarantees a larger drop in the objective during that refresh than the one-shot step, even though both use the same step size and the same gradients. Under the PL condition, the scheduled path maintains the usual contraction factor and gains a nonnegative extra decrease term over the window.
5.4 Exact Recovery of the Population Conflict Graph and Task Partition
We show that, after observing gradients for only a modest number of steps, the scheduler can exactly reconstruct the true conflict relations among tasks by averaging recent gradients (EMA), computing pairwise cosines, thresholding at $\tau$, and coloring the resulting graph. Under a separation margin $\gamma$ around the threshold (tasks are meaningfully different), bounded noise, and bounded drift within each refresh window, the conflict graph estimated from finite data agrees, with high probability, with the ideal population conflict graph $G^{\star}\tau$ (defined from the pairwise cosines of the true mean gradients $\{\mu_{i}\}_{i=1}^{K}$ at the start of the refresh window). Equivalently, when the uniform cosine estimation error is below $\gamma$, we have $\widehat{G}\tau=G^{\star}\tau$ and the resulting grouping recovers the ground-truth task partition. This explains why the scheduler’s group structure is trustworthy and ties the required number of recent gradient measurements per task to interpretable quantities such as noise level, margin, and the number of tasks. For example, an effective sample size of $n_{\mathrm{eff}}\;\gtrsim\;\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\bigl(K/\delta\bigr)$ suffices in our analysis.
5.5 Scheduling Properties with Few Groups and Bounded Staleness
Welsh-Powell greedy coloring uses at most $\Delta+1$ colors on a graph whose maximum degree is $\Delta$ (Bonamy et al., 2018). Running the colors in a fixed cycle means each task is updated at least once every $m\leq\Delta+1$ steps. Equivalently, no task waits more than $\Delta$ steps between updates (bounded staleness).
This means that the schedule length is controlled by the worst conflict degree $\Delta$ rather than by the total number of tasks $K$. This results in two important benefits: (1) a minimum update-frequency guarantee, since every task receives an update at least once per cycle of length $\leq\Delta+1$; and (2) compatibility with standard bounded-delay conditions used in analyses of asynchronous SGD (e.g., Niu et al. 2011; Lian et al. 2015), with delay parameter at most $\Delta$. When $\Delta\ll K$, we achieve both low interference (few conflicts per step) and low staleness (short update gaps).
6 Experimental Setup
6.1 Datasets
We evaluate across six benchmarks spanning vision, multimodal, and time-series. For each dataset we specify a small set of primary tasks and add positive and negative auxiliaries to stress interference. Architectures are standard backbones (e.g., ResNet-18 for image tasks, CNN/BiLSTM for time-series) with task-specific heads. Full dataset and task definitions, auxiliary construction, and architecture details (including preprocessing and head designs) appear in Appendix J and Table 3.
6.2 Baseline and State-of-the-Art Comparisons
We compare against loss-weighting (Uniform, GradNorm, AdaTask), multi-objective (MGDA, Nash-MTL, FairGrad), projection/surgery (PCGrad, CAGrad), and fast adaptive weighting (FAMO). We provide short method notes in Appendix K and discuss these approaches in Section 2.
6.3 Scheduler Extension Models
In addition to standalone models, we also evaluate combinations of the scheduler with existing approaches.
  1. 1.
    SON-GOKU ++ AdaTask. Combines our interference-aware task selection with AdaTask’s dynamic loss weighting, applying adaptive weights only to scheduler-selected tasks.
  2. 2.
    SON-GOKU ++ GradNorm Warm Start. Initializes training with GradNorm for stable gradient magnitudes, then transitions to our scheduler after 3 epochs.
  3. 3.
    SON-GOKU ++ PCGrad. Applied PCGrad’s gradient projection specifically to tasks selected by our scheduler, providing fine-grained conflict resolution within $\tau$-compatible groups.
6.4 Ablation Study
6.4.1 Static One-Shot Coloring
We run the greedy graph coloring once at the start of training, freeze the resulting task groups, and never recompute the conflict graph. All other hyperparameters ($\tau$, history length $H$, and update interval $R$) match the full scheduler. As training progresses we expect the fixed coloring to grow stale, mixing tasks whose interference relationships have changed. This ablation isolates the benefit of dynamic recoloring, showing how much performance depends on adapting the schedule to evolving gradient conflicts.
6.4.2 Single-Step Conflict Estimation
Here, we set the history length to $H=1$, so every recoloring step relies on only the most recent mini-batch gradients to estimate interference. Without aggregation over many past steps, the conflict graph should become highly noisy, causing unstable task groupings from one update window to the next. This variant tests the importance of historical conflict statistics in the scheduler.
7 Results and Discussion
Table 1: Performance of Evaluated Approaches Across Datasets
Model Accuracy (%) $\uparrow$ F&B HEALTH NYUv2
CIFAR-10 AV‑MNIST MM‑IMDb Acc. (%) $\uparrow$ MAE $\downarrow$ Acc. (%) $\uparrow$ MAE $\downarrow$ Angle Error $\downarrow$ Seg. MIOU $\uparrow$ Depth RMSE $\downarrow$
Uniform 55 63 56 45 0.57 52 0.54 21.6 0.059 0.73
GradNorm 61 65 58 47 0.57 53 0.52 21.4 0.054 0.65
MGDA 59 62 56 44 0.57 53 0.53 21.8 0.63 0.75
PCGrad 61 65 58 50 0.55 58 0.48 20.9 0.07 0.69
CAGrad 59 62 57 46 0.58 53 0.52 21.9 0.65 0.73
AdaTask 63 67 59 47 0.59 55 0.52 20.3 0.69 0.65
FAMO 64 70 61 52 0.53 60 0.49 19.9 0.074 0.63
FairGrad 62 66 59 52 0.54 60 0.47 20.7 0.072 0.67
Nash-MTL 63 66 60 52 0.54 60 0.47 20.6 0.073 0.67
Static One-Shot 61 66 58 48 0.56 54 0.51 20.5 0.071 0.65
Single-Step 40 59 20 42 0.60 47 0.55 26.4 0.042 0.81
SON-GOKU ++ GradNorm 62 69 59 51 0.53 59 0.49 19.6 0.073 0.64
SON-GOKU ++ AdaTask 67 71 63 52 0.53 59 0.48 20.1 0.68 0.67
SON-GOKU ++ PCGrad 65 70 60 54 0.52 62 0.45 19.7 0.076 0.62
SON-GOKU 65 69 61 51 0.53 58 0.50 19.8 0.073 0.59
Results for all models across every experiment are depicted in Table 1. Across ten metrics on six datasets, our conflict-aware schedulers consistently match or exceed all baseline methods.
7.1 Overall Performance Improvements
Overall, the conflict-aware approaches improve over the uniform baseline by 10%-20% on CIFAR-10 and by 7% on MM-IMDb. This reinforces the idea that grouping tasks according to measured interference is more effective than treating all tasks equally at every update. On NYUv2, we see similar improvements across all the metrics. These results suggest that the scheduler’s graph coloring cleanly separates high-conflict tasks, preserving the projection or LR-balancing advantages (stemming from PCGrad’s gradient projection and AdaTask’s learning‑rate adaptation, respectively) while removing residual interference.
7.2 Ablation Study on Scheduler Design
Our ablation study (Section 6.4) further highlights the importance of how SON-GOKU is designed. The results show that: (1) Dynamic recoloring matters. Static One-Shot underperforms the full scheduler on most metrics, indicating that task relations change enough during training that frozen groups become stale. This supports the need for periodic refresh; and (2) History smoothing is very important. Single-Step is markedly worse across datasets, consistent with our claim that per-batch cosines are too noisy to define stable groups. This aligns with our recovery analysis, which requires concentration of EMA gradients within a refresh window.
7.2.1 Interpretation of Ablation Study Results
Dynamic recoloring lets the schedule track how task interactions change over time, so partitions do not go stale. Averaging gradients over a short history makes the conflict signal less noisy, which results in more stable groups and reliable progress (Section 5.1 and 5.4). Together, these choices satisfy the conditions we use in our theory. They create low-conflict groups that ensure safe per-step descent (Section 5.15.2) and provide enough concentration to recover the population conflict graph within each refresh (Section 5.4).
7.3 Additional Analysis
7.3.1 Optimizer-Task Alignment
Interestingly, we observe that AdaTask-based approaches tend to be the best on classification tasks (CIFAR-10, AV-MNIST, MM-IMDb) while PCGrad-based approaches tend to be the best on tasks that model regression (NYUv2).
We believe that this stems from unique differences in the features of classification and regression-based models. For example, cross-entropy gradients near decision boundaries tend to be bursty and high in variance (Shrivastava et al., 2016; Lin et al., 2017; Hoffer et al., 2017). By scaling each task’s step size according to its running gradient norm, AdaTask smooths out these spikes.
On the other hand, we believe that PCGrad under the scheduler performs particularly well on regression and dense-prediction tasks as their tasks tend to generate smooth, large-magnitude gradients whose directions change gradually. PCGrad removes only the small component of the gradient that conflicts across tasks, preserving the main descent direction while reducing interference.
7.3.2 Synergy Between Scheduling and Baselines
We believe that the superior results found in the combinations of the scheduler and baseline models can be traced to the way scheduling and optimization reinforce one another.
First, greedy graph coloring partitions tasks into $\tau$-compatible groups, segregating tasks with highly divergent gradients. This yields a guaranteed lower bound on descent (Proposition 6), directly improving optimization efficiency.
Within each low-conflict group, the optimizer can do its job under more ideal conditions. PCGrad can remove the remaining minor conflicting components, preserving the majority of the descent direction. AdaTask can adjust each task’s learning rate without being impacted by large adversarial gradients.
This $\Delta+1$ color bound ensures that every task is scheduled at least once per period. This prevents tasks from being essentially starved of updates.
Finally, by computing interference over a window, the scheduler smooths out gradient fluctuations. This prevents the erratic schedule changes that projection-only grouping methods have been shown to face (Yu et al., 2020; Shi et al., 2023; Zhang et al., 2024), thereby better stabilizing convergence.
7.4 Speed and Tradeoffs
Table 2: Wall-clock time (seconds $\pm$ standard deviation) vs. number of tasks $K$.
Method (R if applicable) K=3 K=6 K=16 K=40
Uniform 0.2656 $\pm$ 0.1201 0.3240 $\pm$ 0.0629 0.3798 $\pm$ 0.1050 0.4054 $\pm$ 0.1190
GradNorm 5.4714 $\pm$ 0.7137 5.1201 $\pm$ 0.6112 4.9042 $\pm$ 0.5869 4.7372 $\pm$ 0.9286
AdaTask 2.1816 $\pm$ 0.0934 2.1032 $\pm$ 0.1012 2.2853 $\pm$ 0.0718 2.2278 $\pm$ 0.1370
PCGrad 3.6212 $\pm$ 0.3517 23.1266 $\pm$ 0.8773 176.7566 $\pm$ 2.8171 1127.1337 $\pm$ 34.2603
MGDA 97.1081 $\pm$ 5.4645 121.4371 $\pm$ 9.0923 132.4913 $\pm$ 3.1752 134.0878 $\pm$ 2.2621
FAMO 2.0725 $\pm$ 0.2073 1.9980 $\pm$ 0.1998 2.1710 $\pm$ 0.2171 2.1164 $\pm$ 0.2116
FairGrad 3.8020 $\pm$ 0.5703 15.2079 $\pm$ 2.2812 108.1450 $\pm$ 16.2218 675.9065 $\pm$ 101.3860
Nash-MTL 5.7030 $\pm$ 1.1406 22.8118 $\pm$ 4.5624 162.2176 $\pm$ 32.4435 1013.8598 $\pm$ 202.7720
SON-GOKU ($R=32$) 1.9896 $\pm$ 0.3651 3.3202 $\pm$ 0.5745 6.0897 $\pm$ 0.9425 12.1432 $\pm$ 1.2044
SON-GOKU ++ AdaTask ($R=32$) 3.7718 $\pm$ 0.9654 5.0511 $\pm$ 0.6531 7.5903 $\pm$ 1.1920 14.5182 $\pm$ 2.0660
SON-GOKU ++ GradNorm ($R=32$) 7.0202 $\pm$ 1.0711 8.1661 $\pm$ 0.9355 10.7227 $\pm$ 2.2088 16.5760 $\pm$ 1.8418
SON-GOKU ++ PCGrad ($R=32$) 1.9834 $\pm$ 0.3586 3.4971 $\pm$ 0.3840 6.1395 $\pm$ 0.9425 10.9097 $\pm$ 1.5263
The proposed scheduler has a time complexity of $\Theta\!\big(K^{2}d/R\big)$ amortized per training step (Section 4.5). Table 2 shows near-linear growth over this range of $K$ at R=32R{=}32, reflecting sparsity in the graphs and batched cosine computation. SON-GOKU’s time rises from around 2 seconds ($K=3$) to 12 seconds ($K=40$), remaining far below methods that perform heavy per-step conflict handling. For example, PCGrad, FairGrad, and Nash-MTL increase steeply with $K$. In contrast, FAMO and AdaTask are among the fastest and largely flat with $K$, as expected from their constant overhead.
These contrasts demonstrate the tradeoffs between speed and fidelity to task interference. Faster methods like FAMO minimize overhead, while methods that model conflicts can improve accuracy. These tradeoffs have to be assessed on a case-by-case basis, based on values that factor into each approach’s time complexity and the importance of training speed versus performance on an application.
8 Conclusion
We introduced SON-GOKU, an interference-aware scheduler that estimates cross-task alignment, builds a sparse conflict graph, and greedily colors it to activate one low-conflict group per step. Formally, we provide rigorous theoretical guarantees that justify the design and effectiveness of the scheduler. Empirically, across six benchmarks, SON-GOKU improves over strong baselines and recent approaches. It complements optimizers like PCGrad and AdaTask, indicating that scheduling and gradient shaping are synergistic. By modeling task interactions with a conflict graph and schedule, SON-GOKU offers a simple, scalable, and theory-backed mechanism for robust multitask training.
Appendix A Full Algorithm Block for Proposed Approach
Algorithm 1 SON-GOKU Scheduler
1:Initial shared params $\theta_{0}$, heads $\{\phi_{k}\}_{k=1}^{K}$, EMA buffers $\tilde{g}_{k}^{(0)}\!=\!0$, total steps $T$, learning-rate schedule $\{\eta_{t}\}$, refresh length $R$, warm-up $T_{\mathrm{warm}}$, target threshold $\tau^{\ast}$, minimum coverage $f_{\min}$, EMA parameter $\beta$
2:Gradients follow the weighted-loss convention (Sec. 4).
[…]
5:for $t=0,\dots,T-1$ do
6:  Warm-up/Anneal: $\tau\leftarrow\textsc{Anneal}(t)$ $\triangleright$ approach in Sec. 4.4
7:  Scheduling: $S_{t}\leftarrow C^{(r)}_{(t\bmod m_{r})+1}$
8:  Forward/Backward:
9:  for all $k\in S_{t}$ do
10:   compute per-task gradients $g_{k}^{(t)}$ and $h_{k}^{(t)}$ (defs: Sec. 4.1)
11:  end for
12:  Parameter update (shared): $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\sum_{k\in S_{t}}g_{k}^{(t)}$
13:  Parameter update (task-specific):
14:  for all $k\in S_{t}$ do
15:   $\phi_{k,t+1}\leftarrow\phi_{k,t}-\eta_{t}h_{k}^{(t)}$
16:  end for
17:  EMA:
18:  for all $k\in S_{t}$ do
19:   update $\tilde{g}_{k}^{(t+1)}$ (Eq. 8)
20:  end for
21:  if $(t+1)\bmod R=0$ then $\triangleright$ refresh
22:   EMA refresh: update all $\tilde{g}_{i}$ using small mini-batches (Sec. 4.1)
23:   Interference matrix: compute $\rho_{ij}^{(t+1)}$ via Eq. 9
24:   Conflict graph: build $G_{\tau}^{(r+1)}$ via Eq. 10
25:   Greedy coloring: Welsh–Powell $\rightarrow\{C^{(r+1)}_{1},\dots,C^{(r+1)}_{m_{r+1}}\}$
26:   Minimum coverage: enforce $f_{i}\!\geq\!f_{\min}$ using compatible-slot duplication (Sec. 4.4.1 )
27:   $r\leftarrow r+1$; $t_{r}\leftarrow t+1$
28:  end if
29:end for
Algorithm block 1 provides an overview of the SON-GOKU scheduler. At a high level, the procedure consists of four stages: (1) estimating pairwise interference, (2) building and coloring the conflict graph, (3) generating a periodic schedule, and (4) updating that schedule as training evolves.
Appendix B Exact Recovery of Population Conflict Graph & Task Partition
B.1 Setting, definitions, and population objects
Let $K\geq 2$ be the number of tasks and $d\geq 1$ the parameter dimension. At designated refresh iterations, the scheduler:
  1. (i)
    computes a per-task exponential moving average (EMA) of stochastic gradients over a probe window of $R$ iterations,
  2. (ii)
    forms a cosine-similarity matrix from the $K$ EMA vectors,
  3. (iii)
    builds a conflict graph by thresholding negative cosines at a fixed level $\tau$ with $\tau\in(0,1)$,
  4. (iv)
    computes a proper coloring of the conflict graph, and
  5. (v)
    schedules one color class per iteration until the next refresh
Definition B.1.
At the beginning of a refresh window (i.e., at a fixed iterate $\theta$), let
$\mu_{i}\;\in\;\mathbb{R}^{d}\qquad(i=1,\dots,K)$ (14)
denote the population task gradients (or the window-stationary means). Define the population cosine matrix $C^{\star}\in[-1,1]^{K\times K}$ by
$C^{\star}_{ij}\;=\;\frac{\langle\mu_{i},\mu_{j}\rangle}{\|\mu_{i}\|\,\|\mu_{j}\|},\qquad i\neq j,\quad C^{\star}_{ii}=1.$ (15)
Definition B.2.
Fix $\tau\in(0,1)$. The population conflict graph $G^{\star}=(V,E^{\star})$ on vertex set $V=\{1,\dots,K\}$ has an edge $\{i,j\}$ iff $C^{\star}_{ij}<-\tau$. The true grouping $\mathcal{P}^{\star}$ is one of:
  1. (A)
    Component Model: the vertex partition given by the connected components of $G^{\star}$.
  2. (B)
    Multipartite model: a partition $V=\bigsqcup_{r=1}^{m}P_{r}$ (with $m\geq!$) such that $G^{\star}$ is the complete $m$-partite graph induced by $\{P_{r}\}_{r=1}^{m}$ (no edges within any $P_{r}$, all cross-part edges present)
When we later speak of group recovery, we mean equality of the empirical partition (defined from data) with $\mathcal{P}^{\star}$, up to label permutation in case (B).
B.2 Assumptions
We adopt the following assumptions, which are standard in analyses of stochastic-gradient methods and verifiable in practice (see, e.g., Robbins & Monro 1951; Kushner & Yin 2003; Nemirovski et al. 2009; Bottou et al. 2018; Wainwright 2019; for concentration of geometrically weighted and mixing sequences, see Merlevède et al. 2011; De la Pena et al. 2009).
Assumption 1 (Separation margin around the threshold).
There exists $\gamma\in(0,1-\tau)$ such that for all $i\neq j$:
[…]
Assumption 2 (Probe noise model and EMA).
In the refresh window of length $R$, the per-iteration stochastic task gradients admit the decomposition
$g_{i,t}\;=\;\mu_{i}+\xi_{i,t},\qquad t=1,\dots,R,$ (17)
where $\{\xi_{i,t}\}_{t=1}^{R}$ are mean-zero, sub-Gaussian with parameter $\sigma^{2}$, and satisfy a $\phi$-mixing or martingale-difference condition ensuring concentration with geometric weights. The EMA for task $i$ is
$\tilde{g}_{i}\;=\;\sum_{t=1}^{R}w_{t}\,g_{i,t},\qquad w_{t}\;=\;\frac{(1-\beta)\,\beta^{R-t}}{1-\beta^{R}},\qquad\beta\in[0,1).$ (18)
Define the effective sample size $n_{\rm eff}$ by
$n_{\rm eff}^{-1}\;:=\;\sum_{t=1}^{R}w_{t}^{2}\;=\;\frac{(1-\beta)^{2}(1-\beta^{2R})}{(1-\beta^{R})^{2}(1-\beta^{2})}.$ (19)
In particular, as $R\to\infty$ (with fixed $\beta\in[0,1)$), we have $n_{\rm eff}\to\frac{1+\beta}{1-\beta}$.
Assumption 3 (Slow drift within a refresh).
Over the refresh window, the changes in $\mu_{i}$ are small enough to be absorbed in the concentration bounds below (equivalently, one can regard $\mu_{i}$ as constant within the window by working at the start-of-window iterate and moving any drift into the noise process).
Assumption 4 (Minimum norm and task inclusion).
There exists $m_{0}>0$ such that $\|\mu_{i}\|\geq m_{0}$ for all tasks included in the graph. In our implementation, we make it so that tasks with $\|\tilde{g}_{i}\|<\nu$ (for a small $\nu\ll m_{0}$) are temporarily excluded from graph construction until stabilized.
Assumption 5 (Threshold selection).
The threshold $\tau$ is fixed across refreshes or selected using data independent of the probe window used to form $\{\tilde{g}_{i}\}$ (e.g., via a separate pilot set). The analysis below treats $\tau$ as deterministic with respect to the probe sample.
B.3 Deterministic group recovery from the conflict graph
We begin with basic graph-theoretic facts that we will use once we have established that the empirical conflict graph coincides with its population counterpart.
Proposition 1 (Chromatic number of a complete multipartite graph).
If $G^{\star}$ is complete $m$-partite with parts $\{P_{r}\}_{r=1}^{m}$, then $\chi(G^{\star})=m$.
Proof.
Picking one vertex from each part yields a clique of size $m$, hence $\chi(G^{\star})\geq m$. Coloring each part with a distinct color is proper, hence $\chi(G^{\star})\leq m$. Therefore $\chi(G^{\star})=m$. ∎
Theorem 1 (Identifiability via optimal coloring under model (B)).
Assume model (B), i.e., $G^{\star}$ is complete $m$-partite with parts $\{P_{r}\}_{r=1}^{m}$. Let $c:V\to\{1,\dots,m\}$ be a proper coloring of $G^{\star}$ that uses exactly $\chi(G^{\star})$ colors. Then each color class equals some part $P_{r}$ (up to relabeling).
Proof.
In a complete multipartite graph, any two vertices from different parts are adjacent. Thus, no color class can contain vertices from two different parts, so each color class is contained in some $P_{r}$. By Proposition 1, $\chi(G^{\star})=m$, so any optimal coloring uses exactly $m$ colors. Since there are $m$ nonempty parts, none can be split across two colors. Hence, the color classes coincide with $\{P_{r}\}_{r=1}^{m}$ up to permutation. ∎
Proposition 2 (Identifiability via components under model (A)).
Under model (A), the grouping $\mathcal{P}^{\star}$ equals the connected components of $G^{\star}$. Consequently, any procedure that returns the connected components of the empirical graph recovers $\mathcal{P}^{\star}$ whenever the empirical graph equals $G^{\star}$.
B.4 Uniform control of empirical cosines from EMA gradients
We now quantify the deviation of the empirical cosine matrix $\widehat{C}$ formed from $\{\tilde{g}_{i}\}$ relative to $C^{\star}$.
Lemma 1 (EMA vector concentration in directions of interest).
Assume Assumption 2 and Assumption 3. There exists a constant $c>0$ depending only on the mixing parameters such that for any fixed unit vector $u\in\mathbb{S}^{d-1}$ and any $\varepsilon>0$.
$\Pr\!\Big(\,\big|\langle\tilde{g}_{i}-\mu_{i},\ u\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (20)
In particular, for any finite set of unit vectors $\{u_{j}\}_{j=1}^{M}$, a union bound yields
$\Pr\!\Big(\,\max_{1\leq j\leq M}\big|\langle\tilde{g}_{i}-\mu_{i},\ u_{j}\rangle\big|>\varepsilon\,\Big)\ \leq\ 2\,M\,\exp\!\Big(-c\,n_{\rm eff}\,\varepsilon^{2}/\sigma^{2}\Big).$ (21)
Proof.
The scalar process $\{\langle\xi_{i,t},u\rangle\}_{t=1}^{R}$ is sub-Gaussian with variance proxy $\sigma^{2}$ and satisfies the same mixing condition. Exponential-weighted averages of such sequences obey Hoeffding-Azuma/Berstein-type tail bounds with variance proxy $\sigma^{2}\sum_{t}w_{t}^{2}=\sigma^{2}/n_{\rm eff}$. The stated inequality follows. ∎
Lemma 2 (Cosine stability under perturbations).
Assume Assumption 4 and let $\epsilon>0$. If for a pair $(i,j)$ we have
$\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{i}-\mu_{i},\ \tfrac{\mu_{i}}{\|\mu_{i}\|}\rangle\big|\leq\epsilon,\qquad\big|\langle\tilde{g}_{j}-\mu_{j},\ \tfrac{\mu_{j}}{\|\mu_{j}\|}\rangle\big|\leq\epsilon,$ (22)
then
$\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \frac{6\,\epsilon}{m_{0}}\,+\,\frac{4\,\epsilon^{2}}{m_{0}^{2}}.$ (23)
Proof.
Write $\tilde{g}_{i}=\mu_{i}+\delta_{i}$, $\tilde{g}_{j}=\mu_{j}+\delta_{j}$. Decompose the numerator and denominator in the cosine:
$\langle\tilde{g}_{i},\tilde{g}_{j}\rangle-\langle\mu_{i},\mu_{j}\rangle=\langle\delta_{i},\mu_{j}\rangle+\langle\mu_{i},\delta_{j}\rangle+\langle\delta_{i},\delta_{j}\rangle,$ (24)
and
$\|\tilde{g}_{i}\|=\|\mu_{i}\|\sqrt{1+2\langle\delta_{i},\mu_{i}\rangle/\|\mu_{i}\|^{2}+\|\delta_{i}\|^{2}/\|\mu_{i}\|^{2}}$ (25)
Using Assumption 4,
$|\langle\delta_{i},\mu_{j}/\|\mu_{j}\|\rangle|\leq\epsilon$ (26)
and
$|\langle\delta_{i},\mu_{i}/\|\mu_{i}\|\rangle|\leq\epsilon$ (27)
imply
$|\langle\delta_{i},\mu_{j}\rangle|\leq\epsilon\|\mu_{j}\|$ (28)
and
$|\langle\delta_{i},\mu_{i}\rangle|\leq\epsilon\|\mu_{i}\|$ (29)
A second-order expansion of the cosine in $(\delta_{i},\delta_{j})$ with the above controls yields the bound. The constants $6$ and $4$ arise from collecting the linear and quadratic contributions in $\epsilon/m_{0}$. ∎
Combining Lemma 1 and Lemma 2 with a union bound over all unordered pairs $(i,j)$ shows that the empirical cosines are uniformly close to their population counterparts.
Proposition 3 (Uniform cosine accuracy with high probability).
Assume Assumption 2, Assumption 3, and Assumption 4. For any $\epsilon>0$ there exist absolute constants $c,C>0$ such that if
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\epsilon^{2}}\,$ (30)
then, with probability $1-\delta$,
$\max_{i<j}\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \epsilon$ (31)
Proof.
For each unordered pair $(i,j)$, apply Lemma 1 with the four unit vectors $\mu_{j}/\|\mu_{j}\|$, $\mu_{i}/\|\mu_{i}\|$, and use Lemma 2 to convert these directional deviations into a cosine deviation bound. A union bound over the $O(K^{2})$ pairs yields the claimed logarithmic factor. The constants absorb the quadratic term in $\epsilon$ by requiring $\epsilon\leq m_{0}$. ∎
B.5 Exact edge recovery and group recovery
We first show that a uniform cosine error smaller than the margin $\gamma$ implies exact equality of empirical and population conflict graphs.
Theorem 2 (Exact conflict-graph recovery under the margin).
Assume Assumptions 15. If
$\max_{i<j}\big|\widehat{C}_{ij}-C^{\star}_{ij}\big|\ \leq\ \epsilon\qquad\text{with}\qquad\epsilon<\gamma,$ (32)
then the empirical conflict graph equals the population graph:
$\widehat{G}\;=\;G^{\star}.$ (33)
Equivalently, for every $i\neq j$,
$C^{\star}_{ij}\leq-(\tau+\gamma)\ \Rightarrow\ \widehat{C}_{ij}<-\tau\quad\text{and}\quad C^{\star}_{ij}\geq-(\tau-\gamma)\ \Rightarrow\ \widehat{C}_{ij}>-\tau.$ (34)
Proof.
For any pair $(i,j)$, if $C^{\star}_{ij}\leq-(\tau+\gamma)$, then $\widehat{C}_{ij}\leq-(\tau+\gamma)+\epsilon<-\tau$, hence $\{i,j\}\in\widehat{E}$. If $C^{\star}_{ij}\geq-(\tau-\gamma)$, then $\widehat{C}_{ij}\geq-(\tau-\gamma)-\epsilon>-\tau$, hence $\{i,j\}\notin\widehat{E}$. ∎
Combining Proposition 3 and Theorem 2 yields a high-probability statement.
Corollary 1 (High-probability exact recovery of $G^{\star}$).
Under Assumptions 15, there exists a universal constant $C>0$ such that if
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Big(\frac{K^{2}}{\delta}\Big),$ (35)
then $\Pr(\widehat{G}=G^{\star})\geq 1-\delta$.
Theorem 3 (Group recovery under the component model).
Under model (A) and the conditions of Corollary 1, with probability at least $1-\delta$, the connected components of $\widehat{G}$ equal $\mathcal{P}^{\star}$.
Proof.
Immediate from $\widehat{G}=G^{\star}$ and the definition of $\mathcal{P}^{\star}$. ∎
Theorem 4 (Group recovery under the multipartite model).
Under model (B) and the conditions of Corollary 1, with probability at least $1-\delta$, $\chi(\widehat{G})=m$ and any optimal coloring of $\widehat{G}$ yields color classes equal to $\{P_{r}\}_{r=1}^{m}$ up to label permutation.
Proof.
If $\widehat{G}=G^{\star}$, then $\widehat{G}$ is complete $m$-partite. Proposition 1 gives $\chi(\widehat{G})=m$. Theorem 1 implies identifiability up to permutation by any optimal coloring. ∎
B.6 Quantitative probe-budget requirement
Combining the bounds above yields the following sample-complexity statement.
Corollary 2.
Under assumptions 15, there exist absolute constants $c,C>0$ such that the following holds. If the EMA parameters $(R,\beta)$ are chosen to ensure
$n_{\rm eff}\ \geq\ C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\qquad\Big(\text{equivalently,}\quad\sum_{t=1}^{R}w_{t}^{2}\ \leq\ c\,\frac{m_{0}^{2}\,\gamma^{2}}{\sigma^{2}}\,\frac{1}{\log(K/\delta)}\Big)$ (36)
then $\Pr(\widehat{G}=G^{\star})\geq 1-\delta$, and consequently Theorems 34 apply. In particular, for fixed $\beta$ and large $R$, $n_{\rm eff}\to\frac{1+\beta}{1-\beta}$ (i.e., it saturates). Thus, to meet the required budget as $K$ grows, one increases $n_{\rm eff}$ by choosing $\beta$ closer to 11 (e.g., $1-\beta\asymp 1/\log(K^{2}/\delta)$), or by switching to a unnormalized averaging approach.
B.7 Summary of the recovery argument
We summarize the logical flow leading to consistency of the scheduler.
  1. (i)
    Assumptions: Assumptions 15 define the conditions in which in which across-group population cosines lie below $(\tau+\gamma)$, within-group cosines lie above $(\tau-\gamma)$, EMA gradients concentrate with effective sample size $n_{\mathrm{eff}}$, and all included tasks have non-negligible gradient norm.
  2. (ii)
    Uniform cosine accuracy: Lemmas 12 together with Proposition 3 yield a high-probability uniform cosine approximation:
    $\max_{i<j}\bigl|\widehat{C}_{ij}-C^{\star}_{ij}\bigr|\ \leq\ \epsilon,$ (37)
    with probability at least $1-\delta$, where $\epsilon$ decreases as $n_{\mathrm{eff}}$ increases.
  3. (iii)
    Exact recovery of edges: If the approximation tolerance satisfies $\epsilon<\gamma$, Theorem 2 converts the uniform bound into exact edge recovery of the conflict graph:
    $\widehat{G}\;=\;G^{\star}.$ (38)
  4. (iv)
    Recovery of the grouping: Given $\widehat{G}=G^{\star}$, Theorem 3 implies group recovery under the component model (groups are the connected components). Under the multipartite model, Proposition 1 and Theorem 1 yield $\chi(\widehat{G})=m$ and Theorem 4 shows that any optimal coloring returns the true parts (up to label permutation).
Quantitative consequence.
Assume Assumptions 15 and fix $\delta\in(0,1)$. Let $m_{0}=\min_{i}\|\mu_{i}\|$ and let $\sigma^{2}$ be the variance proxy from Assumption 2. If the EMA probe budget satisfies
$n_{\mathrm{eff}}\;\geq\;C\,\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\Bigl(\frac{K^{2}}{\delta}\Bigr)$ (39)
for a universal constant $C>0$, then with probability at least $1-\delta$ the empirical conflict graph equals the population graph: $\widehat{G}=G^{\star}$. Consequently:
  1. (i)
    under the component model (A), the connected components of $\widehat{G}$ coincide with $\mathcal{P}^{\star}$.
  2. (ii)
    under the multipartite model (B), $\chi(\widehat{G})=m$ and any optimal coloring of $\widehat{G}$ recovers $\mathcal{P}^{\star}$ up to permutation of labels.
Appendix C Descent Bounds for Scheduled versus Aggregated Updates
We compare two update procedures over a single refresh: a scheduled sequence of per-group steps (i.e., the approach we propose in our paper) and a single aggregated step that combines all groups at once. Both use the same step size $\eta$ and the same gradient information measured at the start of the refresh, and our analysis operates at the level of $L$-smooth (descent) upper bounds. We identify when the scheduled bound is strictly tighter and summarize implications under PL / strong convexity.
Throughout, $F:\mathbb{R}^{d}\to\mathbb{R}$ is differentiable and $L$-smooth, i.e.
$F(y)\ \leq\ F(x)+\langle\nabla F(x),y-x\rangle+\tfrac{L}{2}|y-x|^{2},\quad\forall x,y.$ (40)
We write $\nabla F(x)=\sum_{r=1}^{m}G_{r}(x)$, where each $G_{r}(x)$ is the group gradient for color $r$ (any fixed linear aggregator of task gradients assigned to color $r$ for the current refresh). We use a refresh step size $\eta\in(0,\,1/L]$.
C.1 Single refresh baselines and notation
C.1.1 Single aggregated step
Definition C.1 (Aggregated step).
Starting from the same point $x$, with step size $\eta\in(0,1/L]$ and group gradients $G_{r}^{0}:=G_{r}(x)$ (with $\nabla F(x)=\sum_{r=1}^{m}G_{r}^{0}$), define
$x^{\mathrm{agg}}\ :=\ x-\eta\,\sum_{r=1}^{m}G_{r}^{0}.$ (41)
One-shot $L$-smoothness bound.
Applying $L$-smoothness with $y=x^{\mathrm{agg}}$ yields
$F(x^{\mathrm{agg}})\ \leq\ F(x)\;-\;\eta\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\tfrac{L\eta^{2}}{2}\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}.$ (42)
C.1.2 Scheduled group sequence over one refresh
Definition C.2 (Scheduled refresh).
Starting from the same point $x$, define
$x_{0}:=x,\qquad x_{r}:=x_{r-1}-\eta\,G_{r}(x_{r-1})\quad(r=1,\ldots,m),\qquad x^{\mathrm{sch}}:=x_{m}.$ (43)
Order and notation.
The within refresh order $(1,\ldots,m)$ may be fixed or randomly permuted each refresh. We write $H(\cdot)$ for the Hessian of $F$ and take $\eta\in(0,1/L]$.
Our goal is to compare upper bounds derived from $L$-smoothness for $F(x^{\text{sch}})$ and $F(x^{\text{agg}})$.
C.2 Telescoping bound for scheduled updates
Lemma 3 (Smoothness Expansion for Two Scheduled Groups).
Let $m=2$ and $G_{r}^{0}:=G_{r}(x)$. For any $\eta\in(0,1/L]$,
$F(x^{\text{sch}})$ $\leq F(x)\;-\;\eta\,\langle\nabla F(x),\,G_{1}^{0}\rangle\;+\;\tfrac{L\eta^{2}}{2}\,\|G_{1}^{0}\|^{2}$ (44)
$\quad-\;\eta\,\langle\nabla F(x),\,G_{2}(x_{1})\rangle\;+\;\tfrac{L\eta^{2}}{2}\,\|G_{2}(x_{1})\|^{2}\;+\;\eta^{2}\!\int_{0}^{1}\!\!\big\langle H(x{-}t\eta G_{1}^{0})\,G_{1}^{0},\ G_{2}(x_{1})\big\rangle dt.$
Proof sketch.
Apply the $L$-smoothness inequality at the first step to bound $F(x_{1})$. For the second step, use $L$-smoothness at $x_{1}$ and expand
$\nabla F(x_{1})=\nabla F(x)-\int_{0}^{1}H(x{-}t\eta G_{1}^{0})\,\eta G_{1}^{0}\,dt$ (45)
by the fundamental theorem of calculus along the segment $x\to x_{1}$. ∎
C.2.1 Start-of-refresh reduction under per-group lipschitzness
We adopt the following assumption whenever we compare bounds solely in terms of start-of-refresh measurements. It will be used throughout Sections C.3C.6
Assumption 6 (Per-group lipschitzness).
Each group map $G_{r}(\cdot)$ is $L_{r}$-lipschitz:
$\|G_{r}(u)-G_{r}(v)\|\ \leq\ L_{r}\,\|u-v\|\quad\text{for all }u,v.$ (46)
Under this assumption, for $m=2$ we have $G_{2}(x_{1})=G_{2}^{0}+\delta_{2}$ with $\|\delta_{2}\|\leq L_{2}\eta\|G_{1}^{0}\|$, hence
$\|G_{2}(x_{1})\|\leq\|G_{2}^{0}\|+L_{2}\eta\|G_{1}^{0}\|$ (47)
For general $m$
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\,\eta\sum_{p<r}\|G_{p}^{0}\|\qquad(r=2,\dots,m)$ (48)
When these substitutions are made in scheduled bounds, the induced drift contributions are collected into a nonnegative penalty $R_{m}(x;\eta)$
C.3 Upper bounds for scheduled and aggregated updates (general $m$)
Applying $L$-smoothness $m$ times yields the scheduled upper bound
$\mathrm{UB}_{\text{sch}}(x;\eta)$ $=F(x)\;-\;\eta\,\sum_{r=1}^{m}\big\langle\nabla F(x),\,G_{r}(x_{r-1})\big\rangle\;+\;\frac{L\eta^{2}}{2}\,\sum_{r=1}^{m}\|G_{r}(x_{r-1})\|^{2}$ (49)
$\quad+\;\eta^{2}\sum_{1\leq p<q\leq m}\ \int_{0}^{1}\!\big\langle H\!\big(x-t\eta G_{p}(x_{p-1})\big)\,G_{p}(x_{p-1}),\ G_{q}(x_{q-1})\big\rangle\,dt.$
The aggregated upper bound is the one-shot bound from Equation 42, restated as
$\mathrm{UB}_{\text{agg}}(x;\eta):=F(x)\;-\;\eta\,\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\frac{L\eta^{2}}{2}\,\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}$ (50)
The integrals in Equation 49 are over ordered pairs $p<q$ along the specific sequence $x_{0}\to x_{1}\to\cdots\to x_{m}$; the bound therefore depends on the within-refresh order. Randomizing the order yields an expected version.
In Sections C.4C.6 we express the scheduled bound in terms of $\{G_{r}^{0}\}$ under the per-group lipschitzness assumption. The associated drift terms are aggregated into $R_{m}(x;\eta)$.
C.4 Scheduled and aggregated gap at a common linearization
Define the shorthand
$I_{pq}(x;\eta)\ :=\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle\,dt$ (51)
By expanding $\mathrm{UB}_{\text{sch}}$ around $\{G_{r}^{0}\}$ and collecting the lipschitz drift penalties into $R_{m}(x;\eta)\geq 0$, we obtain:
Theorem 5 (Upper-bound gap under per-group lipschitzness).
Assuming per-group lipschitzness, for any partition $\{G_{r}\}$ and $\eta\in(0,1/L]$,
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\;\leq\;\eta^{2}\!\sum_{1\leq p<q\leq m}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;+\;I_{pq}(x;\eta)\Big)\;+\;R_{m}(x;\eta).$ (52)
Using $\|H(\cdot)\|_{\mathrm{op}}\leq L$ and Cauchy-Schwarz (Steele, 2004)
$I_{pq}(x;\eta)\ \leq\ L\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|$ (53)
which gives the envelope
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ L\eta^{2}\!\sum_{p<q}\!\big(\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|-\langle G_{p}^{0},G_{q}^{0}\rangle\big)\;+\;R_{m}(x;\eta)\ \ \geq\ 0$ (54)
Interpretation
This shows that without additional structure, the scheduled smoothness bound can be looser than the aggregated bound. The gap is governed by Hessian-weighted cross terms $I_{pq}$
Proposition 4 (Drift penalty bound under per-group lipschitzness).
Assume each group map $G_{r}$ is $L_{r}$-lipschitz. Then for $r\geq 2$,
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\eta\sum_{p<r}\|G_{p}^{0}\|\ :=\ \|G_{r}^{0}\|+L_{r}\eta\,S_{r-1},$ (55)
and the scheduled start substitution error satisfies
$R_{m}(x;\eta)\ \leq$ $\ \eta^{2}\Big(\sum_{p=1}^{m}\|G_{p}^{0}\|\Big)\,\sum_{r=2}^{m}L_{r}\,S_{r-1}$ (56)
$\ +\ \frac{L\eta^{2}}{2}\sum_{r=2}^{m}\Big(2\,\|G_{r}^{0}\|\,L_{r}\eta S_{r-1}\;+\;(L_{r}\eta S_{r-1})^{2}\Big),$
so $R_{m}(x;\eta)=O(\eta^{2})$ with constants controlled by $\{L_{r}\}$ and $\{\|G_{r}^{0}\|\}$.
C.5 Sufficient conditions for a tighter scheduled bound
The terms $I_{pq}(x;\eta)$ encode Hessian-weighted interactions between groups and determine when scheduling is advantageous at the bound level.
Assumption 7 (Hessian-weighted negative cross-terms).
There exist nonnegative margins $\{\Gamma_{pq}\}_{p<q}$ such that
$I_{pq}(x;\eta)\ =\ \int_{0}^{1}\!\big\langle H(x{-}t\eta G_{p}^{0})\,G_{p}^{0},\ G_{q}^{0}\big\rangle dt\ \leq\ -\,\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\quad\text{for all }p<q$ (57)
Theorem 6 (Strict upper-bound improvement under per-group lipschitzness and negative Hessian-weighted cross-terms).
Assuming per-group lipschitzness and 57, for any $\eta\in(0,1/L]$,
$\mathrm{UB}_{\text{sch}}(x;\eta)-\mathrm{UB}_{\text{agg}}(x;\eta)\ \leq\ \eta^{2}\!\sum_{p<q}\!\Big(\!-L\,\langle G_{p}^{0},G_{q}^{0}\rangle\;-\;\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|\Big)\ +\ R_{m}(x;\eta)$ (58)
In particular, if
$\sum_{p<q}\!\Big(\Gamma_{pq}\,\|G_{p}^{0}\|\,\|G_{q}^{0}\|+L\,\langle G_{p}^{0},G_{q}^{0}\rangle\Big)\ >\ \frac{R_{m}(x;\eta)}{\eta^{2}}$ (59)
then $\mathrm{UB}_{\text{sch}}(x;\eta)\ <\ \mathrm{UB}_{\text{agg}}(x;\eta)$
C.6 PL or strong convexity: standard rate and upper-bound gains for scheduling
Assume $F$ satisfies the Polyak–Łojasiewicz (PL) inequality with parameter $\mu>0$:
$\tfrac{1}{2}|\nabla F(x)|^{2}\ \geq\ \mu\,(F(x)-F^{\star}),\qquad\forall x$ (60)
For any $\eta\in(0,1/L]$, the single aggregated update satisfies the standard GD bound
$F(x^{\text{agg}})\ \leq\ F(x)\;-\;\eta\Big(1-\tfrac{L\eta}{2}\Big)\,|\nabla F(x)|^{2}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)$ (61)
Define the upper-bound gain (under per-group lipschitzness, so both bounds are expressed at start-of-refresh):
$\Delta{\text{UB}}(x;\eta)\ :=\ \mathrm{UB}_{\text{agg}}(x;\eta)-\mathrm{UB}_{\text{sch}}(x;\eta)\ \geq\ 0$ (62)
whenever 59 holds. Since $F(x^{\text{sch}})\leq\mathrm{UB}_{\text{sch}}(x;\eta)$ and $\mathrm{UB}_{\text{agg}}(x;\eta)$ upper-bounds the one-shot decrease term in 61, we obtain the bound-level contraction
$F(x^{\text{sch}})-F^{\star}\ \leq\ \Big(1-2\mu\eta\big(1-\tfrac{L\eta}{2}\big)\Big)\,\big(F(x)-F^{\star}\big)\;-\;\Delta_{\text{UB}}(x;\eta).$ (63)
Consequently, under per-group lipschitzness and 59, the scheduled refresh satisfies the standard gradient-descent contraction and, in addition, achieves an extra nonnegative decrement $\Delta_{\mathrm{UB}}(x;\eta)$ in the upper bound.
C.7 Why the assumptions are mild
The assumptions we use are mild. They are standard and naturally align with our training pipeline.
C.7.1 $L$-smoothness
This is the same regularity used throughout the main paper and in our baselines. Each task loss we optimize is $L_{i}$-smooth, so the overall objective is $L$-smooth. We only use this to apply the standard smoothness (descent) inequality (Nesterov, 2004; Beck, 2017).
C.7.2 Per-group lipschitzness of $G_{r}$
Each $G_{r}$ is a fixed linear combination of the task gradients assigned to group $r$. If each task gradient is $L_{i}$-lipschitz, then $G_{r}$ is lipschitz with constant $Lr\leq\sum{i\in r}L_{i}$. In other words, this property falls out of task-level smoothness. The same smoothness estimates we already use for step-size selection upper-bound the $L_{r}$.
C.7.3 Negative Hessian-weighted cross-terms
The condition we use asks that, over the short moves we actually take ($\eta\leq 1/L$), groups that are separated by the scheduler continue to exhibit negative interaction under the local Hessian (i.e., the Hessian-weighted cross-terms remain negative). This aligns with how the scheduler is built. It separates tasks that exhibit sustained negative interactions and it periodically refreshes assignments so the local geometry does not drift far. Thus the assumption matches the mechanism we deploy.
C.7.4 PL and strong convexity
We invoke PL only to convert a per-refresh decrease into a standard contraction factor. We do not require global strong convexity. A local PL inequality around the iterates is enough, which is commonly observed after warm-up and annealing we already use (Karimi et al., 2016; Zhou et al., 2021; Liu, 2025).
C.8 Concluding remarks
This appendix formalizes a bound-level comparison between scheduled and aggregated updates. Without additional structure the scheduled bound need not be tighter, but under per-group lipschitzness and negative Hessian-weighted cross-terms it becomes strictly tighter, and under PL the scheduled refresh inherits the standard GD contraction with an additional nonnegative decrement. In practice, these conditions arise naturally once the task-group assignments stabilize, so the scheduler will typically achieve tighter descent bounds without changing step sizes or gradient information.
Appendix D Computational Complexity of One Refresh (and Amortized Over Training)
We analyze the computational and memory complexity of the proposed interference-aware scheduler per refresh and its amortized cost over training. The former accounts for the cost of a single refresh operation while the latter represents the average cost distributed across all training steps. We distinguish the work required by the underlying multi-task training objective (e.g., backpropagation to obtain gradients) from the scheduler overhead (EMA maintenance, cosine computation, conflict graph construction, and color).
D.1 Notation
  • $K\in\mathbb{N}$number of tasks
  • $d\in\mathbb{N}$dimension of the gradient EMA vector per task
  • $R\in\mathbb{N}$refresh period (number of training steps between graph rebuilds)
  • $\beta\in[0,1)$exponential moving average (EMA) parameter
  • $T\in\mathbb{N}$total number of training steps
  • $\mathsf{G}>0$time to compute one backward pass to obtain a task gradient at a refresh
  • $\tau\in(0,1)$ – conflict threshold; an undirected edge $\{i,j\}$ is present iff $\widehat{C}_{ij}<-\tau$
  • $T_{\text{refresh}}>0$time cost of a single scheduler refresh
  • $S_{\text{refresh}}>0$peak additional memory used during a refresh
  • $N_{\text{refresh}}\in\mathbb{N}$number of refreshes over $T$ steps with period $R$ (satisfies $N_{\text{refresh}}\in\{\lfloor T/R\rfloor,\lceil T/R\rceil\}$ and $N_{\text{refresh}}\leq T/R+1$)
D.2 Per-refresh complexity (time and space)
At a refresh, the scheduler performs a finite sequence of deterministic operations on the current collection of task-wise exponential moving averages (EMAs) of gradients. Let
$M\in\mathbb{R}^{K\times d}$ (64)
denote the matrix whose $i$-th row $m_{i}^{\top}$ is the EMA for task $i$. A refresh first updates these rows through a scalar EMA rule
$m_{i}\leftarrow\beta m_{i}+(1-\beta)g_{i}$ (65)
using the most recent probe (or reused) gradient $g_{i}$. It then constructs the cosine-similarity matrix
$\widehat{C}=\widetilde{M}\widetilde{M}^{\top}$ (66)
where $\widetilde{M}$ is the row-normalized version of $M$. It thresholds $\widehat{C}$ at $\tau$ to obtain the conflict adjacency. Finally, it applies a graph-coloring routine to the resulting simple graph (Welsh & Powell, 1967).
EMA maintenance uses a constant number of vector operations per task: one multiply-add on each of the $d$ coordinates of $m_{i}$. Aggregating over all $K$ tasks gives a time proportional to $Kd$. The storage required to hold all EMAs is the $K\times d$ array $M$, so the working set devoted to EMAs is $\Theta(Kd)$ numbers.
The construction of $\widehat{C}$ proceeds by normalizing each row of $M$ and then multiplying $\widetilde{M}$ by its transpose. Row normalization touches each entry exactly once and therefore costs $\Theta(Kd)$ time. The Gram product $\widetilde{M}\widetilde{M}^{\top}$ consists of $K^{2}$ dot products of length $d$, which is $\Theta(K^{2}d)$ time (Kågström et al., 1998). The cosine matrix itself occupies $K^{2}$ entries. If it is retained after thresholding, it uses $\Theta(K^{2})$ space. If dropped right after graph construction, that $\Theta(K^{2})$ storage is only temporary.
Thresholding linearly scans the off-diagonal of $\widehat{C}$, adding an undirected edge when $\widehat{C}_{ij}<-\tau$; this costs $\Theta(K^{2})$ time. The result is either a dense $K\times K$ boolean array requiring $\Theta(K^{2})$ space, or a sparse adjacency whose size depends on the number of conflicts (e.g., $\Theta(kK)$ when retaining the $k$ most negative entries per row).
Putting these pieces together yields the following statement.
Proposition 5 (Per-refresh scheduler overhead).
Under the standard RAM model with dense matrix multiplication costed as $\Theta(K^{2}d)$, the time required by a single scheduler refresh is
$T_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}d)\;+\;\Theta(K^{2})\;+\;O(K^{2})\;=\;\Theta(K^{2}d),$ (67)
and the additional space required by the scheduler during the refresh is
$S_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}),$ (68)
where the $\Theta(K^{2})$ term is transient if $\widehat{C}$ is not retained after coloring.
Proof.
The EMA update costs $\Theta(Kd)$ by a direct count of coordinate-wise multiply-adds. Row normalization also costs $\Theta(Kd)$. The Gram matrix requires $K^{2}$ inner products of length $d$, which is $\Theta(K^{2}d)$. This term dominates $\Theta(Kd)$. Thresholding scans $O(K^{2})$ entries and is therefore $\Theta(K^{2})$. The greedy coloring performs a sort of $K$ keys and then assigns at most one color per edge incident on the current vertex, which is $O(K^{2})$ in the worst case. This is dominated by $\Theta(K^{2}d)$ whenever $d\geq 1$. Summing these contributions and absorbing lower-order terms yields $T_{\text{refresh}}=\Theta(K^{2}d)$. The EMA matrix occupies $\Theta(Kd)$ memory, and storing $\widehat{C}$ uses $\Theta(K^{2})$. But if $\widehat{C}$ is discarded immediately after thresholding, only $\Theta(Kd)$ remains. ∎
D.3 Amortized cost over training
Let $R\in\mathbb{N}$ denote the refresh period as the scheduler executes a refresh once every $R$ training steps. Consider a training run of length $T$ steps. The number of refreshes executed is $\lfloor T/R\rfloor$ or $\lceil T/R\rceil$ depending on whether one refresh occurs at step 0. In either case it is bounded by $T/R+1$. Multiplying the per-refresh time $T_{\text{refresh}}$ by the number of refreshes and dividing by $T$ shows that the amortized scheduler time per training step satisfies
$\frac{1}{T}\,N_{\text{refresh}}\,T_{\text{refresh}}\;\leq\;\frac{1}{T}\,\Big(\frac{T}{R}+1\Big)\,T_{\text{refresh}}\;=\;\frac{1}{R}\,T_{\text{refresh}}+\frac{1}{T}\,T_{\text{refresh}}$ (69)
Letting $T\to\infty$ (or simply taking $T$ large compared to one refresh) eliminates the $T^{-1}T_{\text{refresh}}$ boundary term, yielding the asymptotic amortized bound
$\frac{1}{R}\,T_{\text{refresh}}\;=\;\ \frac{1}{R}\,\Theta(K^{2}d)\$(70)
If probe gradients are computed only at refreshes, their contribution $K\mathsf{G}$ per refresh adds $\frac{1}{R}\Theta(K\mathsf{G})$ to the amortized time per step. If, instead, the training loop already computes task-wise gradients each step and these are reused to update the EMAs, then the probe term is absent and the amortized scheduler overhead remains $\frac{1}{R}\Theta(K^{2}d)$.
The amortized space usage is simpler. The EMA matrix $M$ must be retained throughout training and therefore contributes $\Theta(Kd)$ at all times. The cosine matrix $\widehat{C}$ and the adjacency are constructed only during the refresh. They’re released after coloring, so the $\Theta(K^{2})$ space does not persist. Consequently, the persistent memory overhead attributable to the scheduler is $\Theta(Kd)$, while the peak overhead during a refresh is $\Theta(Kd)+\Theta(K^{2})$.
D.4 Conditions for negligible overhead
Let the amortized per-step costs be
$C_{\text{sched}}=\frac{a}{R}\,K^{2}d\qquad\text{and}\qquad C_{\text{probe}}=\frac{b}{R}\,K\,\mathsf{G},$ (71)
where $a,b>0$ are platform-dependent constants and $\mathsf{G}$ denotes the per-task backpropagation cost of the optional probe at a refresh. For fixed $R$,
$\frac{C_{\text{sched}}}{C_{\text{probe}}}=\frac{a}{b}\,\frac{K^{2}d}{K\,\mathsf{G}}=\frac{a}{b}\,\frac{Kd}{\mathsf{G}}.$ (72)
Hence $C_{\text{sched}}$ is negligible relative to $C_{\text{probe}}$ whenever
$\frac{C_{\text{sched}}}{C_{\text{probe}}}\leq\varepsilon\quad\text{for some }0<\varepsilon\ll 1\quad\Longleftrightarrow\quad Kd\;\leq\;\frac{b}{a}\,\varepsilon\,\mathsf{G}.$ (73)
D.5 Reducing time complexity
In this section, we detail approaches that can be taken under certain circumstances to optimize time complexity.
D.5.1 Random projections
We replace the EMA matrix $M\in\mathbb{R}^{K\times d}$ by a lower-dimensional sketch $\widetilde{M}=MR$ with $R\in\mathbb{R}^{d\times r}$ and $r\ll d$ (Dasgupta & Gupta, 2003). The sketching multiply costs $O(Kdr)$ and the cosine Gram becomes $O(K^{2}r)$ instead of $\Theta(K^{2}d)$. Storage for the sketched EMAs is $O(Kr)$. By the Johnson-Lindenstrauss (JL) random projection guarantee, if we map the $K$ task-EMA vectors from $\mathbb{R}^{d}$ to $\mathbb{R}^{r}$ using a suitable random matrix with $r=\Theta(\epsilon^{-2}\log K)$, then after row normalization all pairwise inner products (hence cosines) are preserved within $\pm\epsilon$ with high probability. We assume a uniform row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ (which can be enforced in practice by skipping tasks with $\|m_{i}\|<\nu\ll m_{0}$) so cosine errors remain controlled. Choosing $\epsilon<\gamma$, where $\gamma$ is the cosine margin from the recovery analysis, ensures that every pair remains on the same side of the threshold $\tau$. Therefore the set $\{(i,j):\widehat{C}_{ij}<-\tau\}$ and the resulting coloring are unchanged with high probability.
In short, dimensionality drops from $d$ to $r$, the refresh cost drops from $\Theta(K^{2}d)$ to $O(Kdr+K^{2}r)$, and decisions are preserved as long as the chosen $r$ makes the embedding error smaller than the margin.
D.5.2 Deterministic covariance sketching via frequent directions
We maintain a deterministic sketch $B\in\mathbb{R}^{\ell\times d}$ of the row space of $M$ using Frequent Directions and either project rows onto $\mathrm{span}(B)$ or form an approximate Gram from the sketch (Liberty, 2013; Ghashami et al., 2016). Maintaining the sketch costs $O(Kd\ell)$, the cosine Gram in the sketch space costs $O(K^{2}\ell)$, and storage for the sketch is $O(\ell d)$. Frequent Directions gives a spectral-norm bound
$|MM^{\top}-\widehat{MM^{\top}}|_{2}\leq\epsilon|M|_{F}^{2}$ (74)
when $\ell=\Theta(\epsilon^{-2})$, which yields a uniform bound on inner-product and squared-norm errors. Assuming a row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ and applying a standard cosine perturbation bound after row normalization, one obtains
$\big|\cos(m_{i},m_{j})-\widehat{\cos}(m_{i},m_{j})\big|\;\leq\;\frac{2\,\epsilon\,\|M\|_{F}^{2}}{m_{0}^{2}}\;+\;O\!\left(\frac{\epsilon^{2}\,\|M\|_{F}^{4}}{m_{0}^{4}}\right)\$(75)
Taking $\epsilon$ small enough so that the right-hand side is $\gamma$ ensures that all threshold decisions and the resulting coloring are preserved deterministically. Thus the effective dimension drops from $d$ to $\ell$ in the worst case, and the refresh cost becomes $O(Kd\ell+K^{2}\ell)$.
D.5.3 Edge sampling for conflict graphs with adaptive refinement
We reduce the number of cosine evaluations by computing $\widehat{C}_{ij}$ for only $\tilde{O}(K\log K)$ randomly chosen task pairs to build a provisional conflict graph and then refining by evaluating additional pairs that are near the threshold or needed to certify connectivity and chromatic structure. We still compute all $K$ row norms once in $O(Kd)$ time for normalization, and the first pass costs $O(Kd\log K)$ for the sampled dot products. The total cost adds only the refinement work, which remains small when only few pairs are ambiguous. Under a planted separation model with margin $\gamma$ and reasonably dense cross-group conflicts, one can show with high probability that the sampled graph already captures the correct inter-group connectivity, so the coloring or component structure is recovered after the first pass and only boundary pairs need refinement. This reduces the pairwise work from $K^{2}$ to near $K\log K$ while preserving the final decisions under stated assumptions (Erdős & Rényi, 1960).
D.5.4 Incremental gram updates
We avoid rebuilding the full cosine matrix when only a small subset of tasks has meaningfully changed since the last refresh. If $s$ rows of $M$ cross a chosen change threshold, we first renormalize these rows and then recompute both the corresponding $s$ rows and $s$ columns of the Gram by taking dot products against all $K$ rows, which costs $O(sKd)$, with an additional $O(sd)$ to update norms, instead of $\Theta(K^{2}d)$, and we leave all unchanged entries as they are. This update is exact for the affected entries, so conflict edges and coloring decisions are preserved by construction, and the reduction is deterministic whenever $s\ll K$. To prevent slow drift in the unchanged entries, we can periodically force a full rebuild and reset the change counters.
Appendix E Descent Preservation Under $\tau$-Compatibility
E.1 Proof of Proposition 6
Proposition 6.
Let $S\subseteq\{1,\dots,K\}$ be a $\tau$-compatible task set. That is, every pair of gradients satisfies
$\bigl\langle g_{i},\,g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert,\qquad\forall\,i\neq j\in S,0\leq\tau<1$ (76)
Then
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}\;\;\geq\;\;\bigl(1-\tau(|S|-1)\bigr)\sum_{k\in S}\lVert g_{k}\rVert^{2}.$ (77)
Proof.
We begin with the polarization identity for any finite set of vectors:
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}=\sum_{k\in S}\lVert g_{k}\rVert^{2}+2\!\!\sum_{\begin{subarray}{c}i,j\in S\\ i<j\end{subarray}}\!\bigl\langle g_{i},g_{j}\bigr\rangle.$ (78)
E.1.1 Lower-bounding the cross terms
Because $S$ is $\tau$-compatible, inequality (76) gives
$\bigl\langle g_{i},g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (79)
Insert this bound into (78) to obtain
$\Bigl\lVert\textstyle\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-2\tau\!\!\sum_{i<j}\!\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (80)
E.1.2 Symmetrizing the mixed sum
Observe that
$\sum_{i<j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;=\;\frac{1}{2}\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (81)
Substituting (81) into (80) yields
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-\;\tau\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (82)
E.1.3 Bounding the mixed sum via Cauchy-Schwarz
Apply the Cauchy-Schwarz inequality in $\mathbb{R}^{|S|}$ to the vectors $a=(\lVert g_{1}\rVert,\dots,\lVert g_{|S|}\rVert)$ and $\mathbf{1}=(1,\dots,1)$:
$\sum_{k}\lVert g_{k}\rVert=\langle a,\mathbf{1}\rangle\;\leq\;\lVert a\rVert\,\lVert\mathbf{1}\rVert=\Bigl(\sum_{k}\lVert g_{k}\rVert^{2}\Bigr)^{1/2}\,\sqrt{|S|}.$ (83)
Using $\left(\sum_{k}a_{k}\right)^{2}\leq|S|\sum_{k}a_{k}^{2}$ and (84),
$\sum_{i\neq j}\|g_{i}\|\,\|g_{j}\|\;=\;\Bigl(\sum_{k}\|g_{k}\|\Bigr)^{2}\;-\;\sum_{k}\|g_{k}\|^{2},$ (84)
we obtain the standard estimate
$\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;(|S|-1)\sum_{k}\lVert g_{k}\rVert^{2}.$ (85)
Hence,
$\tau\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\tau\,\bigl(|S|-1\bigr)\sum_{k}\lVert g_{k}\rVert^{2}.$ (86)
E.1.4 Combining bounds
Insert (86) into (82):
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}-\tau\bigl(|S|-1\bigr)\sum_{k}\lVert g_{k}\rVert^{2}\;=\;\bigl(1-\tau(|S|-1)\bigr)\sum_{k}\lVert g_{k}\rVert^{2},$ (87)
which is (77). ∎
E.2 Interpretation and practical implications
Equation (77) guarantees that whenever we restrict an SGD step to a $\tau$-compatible group (i.e., a set of tasks whose gradients are not too conflicting) the resulting joint update preserves at least a $\bigl(1-\tau(|S|-1)\bigr)$ fraction of the summed squared step lengths.
Below, we provide a strictly stronger version that is assumption free.
Proposition 7 (Data-Dependent Lower Bound via the Aggregate Conflict Ratio).
Define the aggregate conflict ratio
$\tau_{\mathrm{eff}}(S)\;:=\;\frac{\displaystyle\sum_{i\neq j}\bigl(-\langle g_{i},g_{j}\rangle\bigr)_{+}}{\displaystyle\sum_{k}\|g_{k}\|^{2}},\qquad(x)_{+}:=\max\{x,0\}.$ (88)
Then, without additional assumptions,
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}\;\geq\;\Bigl(1-\tau_{\mathrm{eff}}(S)\Bigr)\sum_{k\in S}\lVert g_{k}\rVert^{2},$ (89)
and under $\tau$-compatibility we always have $\tau_{\mathrm{eff}}(S)\leq\tau(|S|-1)$, so (89) is never weaker than (77).
Our takeaways from this are as follows:
  1. (i)
    Descent direction safety. The aggregated step is guaranteed to be a descent direction whenever $\tau_{\mathrm{eff}}(S)<1$ (data-dependent) and, in particular, whenever $\tau(|S|-1)<1$ (worst-case).
  2. (ii)
    Convergence-rate constant. In analyses for smooth SGD, one may replace $\lVert g_{t}\rVert^{2}$ by the right-hand side of either (89) (which is tighter) or (77) (worst-case), leading respectively to constants involving $\tau_{\mathrm{eff}}(S_{t})$ or $\tau(|S_{t}|-1)$.
Appendix F Convergence rate with $\tau$ -dependent constant
Theorem 7 (Baseline $O(1/\sqrt{T})$ convergence of the full gradient).
Let $F(\theta)=\sum_{k=1}^{K}\mathcal{L}_{k}(\theta,\phi_{k})$ be LL–smooth in the shared parameters $\theta$. Assume the stochastic gradient $g_{t}$ obtained at step $t$ satisfies $\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t})$ and $\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}\mid\theta_{t}]\leq\sigma^{2}$. Let the step size be $\eta=\tfrac{c}{\sqrt{T}}$ with $0<c\leq\frac{1}{L}$, and suppose the scheduler selects a $\tau$-compatible task set $S_{t}$ at each step (this will be used below for a refinement). Then
$\min_{1\leq t\leq T}\;\mathbb{E}\!\left[\lVert\nabla F(\theta_{t})\rVert^{2}\right]\;\leq\;\frac{2\bigl(F_{0}-F^{\star}\bigr)}{c\sqrt{T}}\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}.$ (90)
Proof.
Because $F$ is LL–smooth, for any $\eta\leq\tfrac{1}{L}$ the standard non-convex SGD inequality (Ghadimi & Lan 2013, Lemma 3.2) gives
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\;\leq\;\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\frac{\eta}{2}\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (91)
Summing equation 91 over $t=0,\dots,T-1$ and using $\mathbb{E}[F(\theta_{T})]\geq F^{\star}$ yields
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (92)
Dividing by $T$, using $\min_{t}x_{t}\leq\tfrac{1}{T}\sum_{t}x_{t}$, and substituting $\eta=\tfrac{c}{\sqrt{T}}$ gives equation 90. ∎
Data-dependent $\tau$-refinement for the scheduled gradient energy.
For a finite set $S$, define the aggregate conflict ratio
$\tau_{\mathrm{eff}}(S):=\frac{\sum_{i\neq j\in S}\bigl(-\langle g_{i},g_{j}\rangle\bigr)_{+}}{\sum_{k\in S}\|g_{k}\|^{2}}\in[0,\infty),\qquad(x)_{+}=\max\{x,0\}.$ (93)
Then for every step $t$,
$\Bigl\|\textstyle\sum_{k\in S_{t}}g_{k,t}\Bigr\|^{2}\;\geq\;\Bigl(1-\tau_{\mathrm{eff}}(S_{t})\Bigr)\sum_{k\in S_{t}}\|g_{k,t}\|^{2}.$ (94)
Consequently,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\sum_{k\in S_{t}}\|g_{k,t}\|^{2}\Bigr]\;\leq\;\underbrace{\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\frac{1}{1-\tau_{\mathrm{eff}}(S_{t})}\Bigr]}_{=:~\Gamma_{T}}\cdot\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\|g_{t}\|^{2}\bigr].$ (95)
Using $\mathbb{E}\|g_{t}\|^{2}=\mathbb{E}\|\nabla F(\theta_{t})\|^{2}+\mathbb{E}\|g_{t}-\nabla F(\theta_{t})\|^{2}\leq\mathbb{E}\|\nabla F(\theta_{t})\|^{2}+\sigma^{2}$ and the average version of equation 91,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\|\nabla F(\theta_{t})\|^{2}\bigr]\;\leq\;\frac{2(F_{0}-F^{\star})}{\eta T}\;+\;L\eta\sigma^{2},$ (96)
we obtain the $\tau$-dependent, data-driven control
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\sum_{k\in S_{t}}\|g_{k,t}\|^{2}\Bigr]\;\leq\;\Gamma_{T}\Biggl(\frac{2(F_{0}-F^{\star})}{\eta T}+L\eta\sigma^{2}+\sigma^{2}\Biggr).$ (97)
If, in addition, each $S_{t}$ is pairwise $\tau$-compatible with $|S_{t}|=s_{t}$ and $\tau\,(s_{t}-1)\leq\rho<1$ uniformly in $t$, then $\tau_{\mathrm{eff}}(S_{t})\leq\tau(s_{t}-1)\leq\rho$ and hence $\Gamma_{T}\leq\tfrac{1}{1-\rho}.$ With $\eta=\tfrac{c}{\sqrt{T}}$, equation 97 becomes
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\Bigl[\sum_{k\in S_{t}}\|g_{k,t}\|^{2}\Bigr]\;\leq\;\frac{1}{1-\rho}\left(\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}\;+\;\sigma^{2}\right).$ (98)
F.1 Discussion and intuition
Equation equation 90 is the classical $O(1/\sqrt{T})$ rate for non-convex SGD with unbiased and bounded variance gradients and constant-over-time step size $\eta=c/\sqrt{T}$. Under these conditions, the convergence rate in terms of the full gradient norm $\|\nabla F(\theta_{t})\|^{2}$ does not depend on $\tau$. However, the scheduler’s $\tau$ structure does control the per step energy of the scheduled gradient through equation 95–equation 98. Less cross-task conflict (smaller $\Gamma_{T}$) results in a tighter bound on $\tfrac{1}{T}\sum_{t}\sum_{k\in S_{t}}\|g_{k,t}\|^{2}$, which is the quantity governed by the descent preservation inequalities used throughout the analysis.
Appendix G Bounded Staleness via Greedy Graph Coloring
Proposition 8 (Staleness Bound).
Let $G=(\mathcal{T},E)$ be the task–conflict graph whose vertices are tasks and whose edges connect pairs with interference coefficient exceeding the threshold $\tau$. Denote by $\Delta$ its maximum degree. Greedy graph coloring produces a proper coloring $C_{1},\dots,C_{m}$ with
$m\;\leq\;\Delta+1.$ (99)
If the scheduler activates the color classes in the cyclic order $C_{1}\!\to\!C_{2}\!\to\!\dots\!\to\!C_{m}\!\to\!C_{1}\!\to\!\dots$, then every task is updated at least once every
$s_{\max}\;=\;m-1\;\leq\;\Delta$ (100)
iterations. In particular, the schedule enforces a bounded inter-update delay of at most $\Delta$ iterations per task, consistent with the bounded-delay assumption of Recht et al. (Niu et al., 2011).
Proof.
We proceed in two parts.
Part A: Color count bound.
A greedy algorithm scans vertices in some order and assigns to each vertex the smallest available color not used by its already colored neighbors. When the $i$-th vertex $v$ is reached, at most $\deg(v)\leq\Delta$ of its neighbors are already colored, so at most $\Delta$ colors are unavailable. Therefore one of the first $\Delta+1$ colors is always free, implying $m\leq\Delta+1$ (Lovász, 2006).
Part B: Staleness of cyclic execution.
Fix any task $T\in\mathcal{T}$ and let it belong to color $C_{j}$ for some $1\leq j\leq m$. Under cyclic scheduling, $C_{j}$ is executed at steps $t=j,\;j+m,\;j+2m,\dots$. The number of intervening steps between two consecutive executions of $C_{j}$ is exactly m1m-1. Hence task $T$ never waits more than $s_{\max}=m-1$ iterations for an update. Combining with Equation 8 yields $s_{\max}\leq\Delta$. ∎
G.1 Interpretation
The bound (Equation 100) guarantees that the shared parameters used by any task are refreshed at least once every $\Delta$ iterations in the worst case (e.g., when the conflict graph is a clique of size $\Delta+1$). This aligns with the bounded-delay assumption common in analyses of asynchronous SGD and lock-free training, so convergence proofs built under that assumption apply to our cyclic schedule with delay parameter at most $\Delta$ when iterations are used as the unit of delay (Niu et al., 2011; Lian et al., 2015). In practice $\Delta$ is often much smaller than the total number of tasks, so the scheduler achieves low interference and low parameter staleness simultaneously.
Appendix H Greedy Graph-Coloring Uses at Most $\Delta\!+\!1$ Colors
H.1 Proof of Proposition 9
Proposition 9 (Coloring Period Bound).
Let $G=(V,E)$ be a finite, simple, undirected graph with maximum degree $\Delta\,\coloneqq\!\max_{v\in V}\deg(v)$. The greedy (first-fit) coloring algorithm (e.g., Welsh–Powell order) [^1] Order the vertices in non–increasing degree and assign to each the smallest positive integer (color) not used by its previously colored neighbors. produces a proper vertex coloring with no more than
$\chi_{\text{greedy}}(G)\;\leq\;\Delta+1$ (101)
distinct colors. Consequently, when the scheduler activates the color classes in a cyclic order, the cycle length is bounded by $\Delta+1$. This is a quantity depending only on the structure of the conflict graph.
Proof.
Let the vertices be processed in the chosen order $v_{1},v_{2},\dots,v_{|V|}\,$ (e.g., Welsh-Powell). Assume inductively that after coloring the first k1k-1 vertices the algorithm has used at most $\Delta+1$ colors. Consider vertex $v_{k}$. Since $\deg(v_{k})\leq\Delta$, at most $\Delta$ neighbors of $v_{k}$ can appear before $v_{k}$ in the ordering. Hence, at the moment of coloring $v_{k}$, at most $\Delta$ colors are forbidden (one for each previously colored neighbor). Among the palette $\{1,2,\dots,\Delta+1\}$ there is therefore at least one color still available. Assigning the smallest such color to $v_{k}$ maintains a proper coloring and never introduces a new color beyond $\Delta+1$.
Proceeding vertex-by-vertex, no step ever requires more than $\Delta+1$ colors, establishing equation 101. ∎
H.2 Implications for the scheduler
A coloring with at most $\Delta+1$ classes means the scheduler’s cycle period (the number of batches needed before every task reappears) is bounded by a graph invariant independent of the number of tasks. Even if thousands of tasks exist, as long as each one conflicts with at most $\Delta$ others, the memory footprint (one shared backbone plus $\Delta+1$ sets of head activations) and the maximum waiting time between successive updates for any task (bounded by $\Delta$, see Proposition 8) remain predictable and small. This guarantee is essential for scaling the scheduler to large, heterogeneous tasks.
Appendix I Baseline Non-Convex SGD Convergence Rate
I.1 Proof of Theorem 8
Theorem 8 (Classical $O(1/\sqrt{T})$ bound).
Let $F:\mathbb{R}^{d}\!\to\!\mathbb{R}$ be an $L$-smooth, possibly non-convex objective and suppose the stochastic gradient $g_{t}$ computed at iteration $t$ satisfies
$\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t}),\qquad\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}\mid\theta_{t}]\;\leq\;\sigma^{2}.$ (102)
Run SGD with the constant step size $\eta=\tfrac{c}{\sqrt{T}},\;0<c\leq\frac{1}{L},$ for $T$ iterations starting from $\theta_{0}$. Then
$\min_{0\leq t<T}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;\frac{2\bigl(F_{0}-F^{\ast}\bigr)}{c\sqrt{T}}+\frac{cL\sigma^{2}}{\sqrt{T}},$ (103)
where $F^{\ast}=\inf_{\theta}F(\theta)$.
Proof.
The proof is a streamlined restatement of ((Ghadimi & Lan, 2013; Nemirovski et al., 2009)). By $L$-smoothness,
$F(\theta_{t+1})\;\leq\;F(\theta_{t})+\langle\nabla F(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\lVert\theta_{t+1}-\theta_{t}\rVert^{2}.$ (104)
With $\theta_{t+1}=\theta_{t}-\eta\,g_{t}$ and taking conditional expectation,
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\leq\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\eta\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L}{2}\,\mathbb{E}\!\bigl[\lVert g_{t}\rVert^{2}\bigr].$ (105)
Decompose the squared stochastic gradient:
$\mathbb{E}[\lVert g_{t}\rVert^{2}]=\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}]\leq\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\sigma^{2}$ (106)
Thus, and using $\eta\leq 1/L$ so that $\eta-\tfrac{L\eta^{2}}{2}\geq\tfrac{\eta}{2}$,
$\mathbb{E}[F(\theta_{t+1})]\leq\mathbb{E}[F(\theta_{t})]-\frac{\eta}{2}\,\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (107)
Summing from $t=0$ to T1T-1 and telescoping gives
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}[\lVert\nabla F(\theta_{t})\rVert^{2}]\leq F_{0}-F^{\ast}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (108)
Dividing by $\eta T$ and inserting $\eta=c/\sqrt{T}$ yields equation 103. ∎
I.2 Connection to the scheduler
At $\tau=0$, pairs with negative inner product are incompatible, so the conflict graph on tasks can be colored into $m$ classes $\{C_{1},\dots,C_{m}\}$, and a simple policy activates one color class per step. Under a deterministic (cyclic) activation order, the update $g_{t}=\sum_{k\in S_{t}}g_{k,t}$ generally satisfies
$\mathbb{E}\bigl[g_{t}\mid\theta_{t}\bigr]=\sum_{k\in S_{t}}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t})\;\neq\;\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t}),$ (109)
so it is biased for the full gradient.
I.2.1 Consistency with the unbiased SGD assumption
The analysis in Theorem 8 assumes an unbiased stochastic gradient, $\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t})$. This assumption is met under either of the following implementations.
(i) Randomized class sampling with scaling. Draw $J_{t}\sim\mathrm{Unif}\{1,\ldots,m\}$ independently each step and set
$\tilde{g}_{t}\;=\;m\!\!\sum_{k\in C_{J_{t}}}\!g_{k,t}.$ (110)
Then $\mathbb{E}[\tilde{g}_{t}\mid\theta_{t}]=\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\theta_{t},\phi_{k,t})=\nabla F(\theta_{t}),$ so Theorem 8 applies (with the variance bound adjusted for the scaled estimator). Equivalently, one may keep $g_{t}=\sum_{k\in C_{J_{t}}}g_{k,t}$ and use an effective step size $m\eta$.
(ii) Deterministic cyclic schedule. If the classes are visited in a fixed periodic order, then generally $\mathbb{E}[g_{t}\mid\theta_{t}]\neq\nabla F(\theta_{t})$ at the per-step level. Nonetheless, standard analyses of nonconvex smooth cyclic block updates yield an $O(1/\sqrt{T})$ decay of the average gradient norm under usual step-size conditions, with constants depending on the number of blocks.
Either implementation delivers an $O(1/\sqrt{T})$ convergence guarantee.
Appendix J Experimental Setup for Datasets
Table 3: Information on the datasets utilized in experimentation. (*Some samples were removed during preprocessing)
Dataset Main Tasks (+) Aux. Tasks (-) Aux Tasks Modalities Samples
NYUv2 Semantic Segmentation [2pt] Depth Estimation [2pt] Surface Normal Prediction Color Temp. Estimation Image 250*
CIFAR-10 Image Classification Quadrant Localization [2pt] Texture Classification Corruption‐Type Prediction [2pt] Rotation Angle Prediction Image 2,500*
AV-MNIST Digit Classification Digit Parity Audio, Image 56.0k
$M$-IMDb Genre Classification Release Decade Title-Iniial Classification Image, Text 25.9k
STOCKS-F&B $4\times$ Stock Return Prediction Five-Day Rolling Volatility [2pt] Sector-Average Next-Day Return Day of the Week Prediction [2pt] Lag-0 Reconstruction of Today’s Open-Price Timeseries $\times 18$ 75.5k
STOCKS-HEALTH $7\times$ Stock Return Prediction Five-Day Rolling Volatility [2pt] Sector-Average Next-Day Return Day of the Week Prediction [2pt] Lag-0 Reconstruction of Today’s Open-Price Timeseries $\times 63$ 75.5k
We evaluate the proposed scheduler alongside numerous baselines and state-of-the-art models across multiple datasets to reliably assess its general performance relative to other approaches. In total, it is evaluated across 6 datasets.
Across all datasets, we incorporate positive and/or negative auxiliary tasks into training. Positive auxiliary tasks share structure or predictive signals with the main tasks (e.g., common features or correlated outputs) and so can improve the learned representations by providing relevant supervision. In contrast, negative auxiliary tasks are uncorrelated or directly conflicting with the main objectives, inducing gradient interference that can slow or degrade primary performance. Including both creates controlled variation in task alignment, letting us test whether SON-GOKU (1) groups compatible tasks, (2) separates conflicting tasks, and (3) maintains main-task performance under interference created by auxiliary tasks.
J.1 NYUv2
The NYU Depth Dataset v2 (NYUv2) (Silberman et al., 2012) consists of RGB-D indoor scenes with 1,449 densely labeled pairs of RGB and depth images. To demonstrate auxiliary task value in data-scarce conditions, we employ a subset of 250 training samples randomly selected from the original training set.
We formulate a multi-main-task setup with three primary objectives: (1) semantic segmentation (14 classes), (2) depth estimation where the model predicts per-pixel depth values from RGB images, and (3) surface normal prediction where 3-channel surface normals are estimated from RGB input. The negative auxiliary task is color temperature estimation, a synthetically generated task that predicts global color temperature properties designed to interfere with the main tasks by emphasizing global color distribution rather than local semantic and geometric features.
All tasks utilize RGB images as the sole input modality, with depth maps and surface normals serving as prediction targets rather than input features. A ResNet-18 (He et al., 2015) backbone trained from scratch processes the RGB input, with task-specific decoder heads for segmentation (with 32 $\times$ upsampling), depth regression, surface normal regression, and color temperature estimation.
J.2 CIFAR-10
The CIFAR-10 (Krizhevsky et al., 2009) dataset contains 60,000 $32\times 32$ color images across 10 generic classes. To evaluate our interference-aware scheduler in a data-scarce environment where auxiliary tasks provide maximum benefit, we employ a subset of 2,500 training samples (250 per class) from the original 50,000 training images.
For the multi-task learning setup, we set image classification as the main task and construct three auxiliary tasks synthetically from the RGB images. The positive auxiliary tasks include: (1) quadrant localization, where the model predicts which quadrant contains the primary object, and (2) texture classification using Gabor filter responses clustered into 8 texture categories via k-means clustering. The negative auxiliary tasks consist of: (3) corruption-type prediction, where images are artificially corrupted using 15 different corruption types from the ImageNet-C corruption suite (Hendrycks & Dietterich, 2019), and (4) rotation angle prediction, where images are rotated by 0°, 90°, 180°, or 270° and the model predicts the rotation angle.
All tasks share a ResNet-18 (He et al., 2015) backbone trained from scratch without pretraining, with task-specific heads for each auxiliary task.
J.3 AV-MNIST
The AV‑MNIST benchmark (Vielzeuf et al., 2018) pairs MNIST images (Lecun et al., 1998) with a log‑mel spectrogram of the corresponding spoken digit from TIDIGITS (Leonard & Doddington, 1993). It is a synthetic benchmark that has significant noise applied to audio and feature reduction applied to images, making it far more difficult than the original MNIST.
We use all paired samples in our experiments. Our primary task is 10‑way digit classification. Following (Vielzeuf et al., 2018), we encode images with a small 4‑layer convolutional network and spectrograms with a 2‑layer CNN, both built and trained from scratch. These embeddings are projected and fused for processing by a simple MLP in intermediate fusion (Boulahia et al., 2021; Guarrasi et al., 2025), as are the models trained on MM-IMDb and STOCKS. We include only one positive auxiliary class, Digital Parity. This task aims to identify the digits as either even or odd, which has been shown to be a positive auxiliary task for improving representations on MNIST-like datasets (Tacchetti et al., 2018; Mohammadi et al., 2020).
J.4 MM-IMDb
The MM‑IMDb dataset (Arevalo et al., 2017) contains 25,959 movies with genre annotations over 23 categories. We extract poster images and plot summaries for every movie in the dataset.
The images and summaries are encoded by a frozen VGG16 (Simonyan & Zisserman, 2014) and Google word2vec (Mikolov et al., 2013) model, respectively. Our main task is movie genre prediction. We add one positive auxiliary task, Release Decade, and one negative auxiliary task, the classification of the title’s first word as either a vowel or consonant.
J.5 STOCKS
The STOCKS datasets we use, introduced in (Liang et al., 2021), contain stock market timeseries data across two categories. Specifically: (1) STOCKS-F&B, which has 14 input and 4 output stocks in the GICS Restaurants or Packaged Food & Meats category (MSCI Inc. & S&P Dow Jones Indices, 2024), and (2) STOCKS-HEALTH, which contains 56 input and 7 output stocks in the Health Care category.
Every input stock consists of 500 trading days, with the goal of predicting returns over the next day. We discretize the continuous return variable $R$ into three non-overlapping categories: (1) $Low$, where $0\leq R<0.1$, (2) $Medium$, where $0.1\leq R<0.5$, and (3) $High$, where $R\geq 0.5$. Mean Absolute Error (MAE) is calculated by mapping the three classes to numbers ($Low\rightarrow 0$, $Medium\rightarrow 1$, $High\rightarrow 2$) and then deriving MAE as usual. Each input series is encoded by the same CNN-BiLSTM network. This consists of 3 CNNs and 1 BiLSTM (Cui et al., 2018).
We augment the main prediction task with two positive auxiliaries and two negative auxiliaries. The first positive task, Five-Day Rolling Volatility, is calculated as the standard deviation of daily logarithmic returns over a sliding five‑trading‑day window. This feature captures short-term fluctuations in a stock’s price. In Sector-Average Next-Day Return, for each date we compute the mean of the actual next‑day returns of all stocks within the same GICS sector, providing a simple measure of sector‑level momentum and drift
The negative tasks focus on useless information that is meant to distract the model. Namely, day of the week prediction (in the range of Monday to Friday) and Lag-0 Open-Price Reconstruction, which requires the model to reproduce the same day’s opening price verbatim. The first is information that contains little to no signals that would contribute to overall performance, and the second is a trivial identity mapping that contributes no real predictive challenge.
Appendix K Models Used for Comparison
K.1 Baseline models
  1. 1.
    Uniform. This baseline assigns equal weights to all tasks throughout training, representing the simplest approach where all task losses are weighted equally.
  2. 2.
    Gradnorm (Chen et al., 2018). Balances task learning rates by normalizing gradient magnitudes relative to target loss ratios. This maintains consistent training dynamics across tasks.
  3. 3.
    MGDA (Sener & Koltun, 2018). Formulates multi-task learning as a multi-objective optimization problem, finding Pareto-optimal solutions (Lockwood, 2008; Pareto, 2014) through gradient descent in the convex hull of gradients (Fliege & Svaiter, 2000; Miettinen, 1999).
K.2 State-of-the-art models
  1. 1.
    PCGrad (Yu et al., 2020). Projects conflicting gradients onto orthogonal subspaces when negative cosine similarity is detected, eliminating destructive interference between task gradients.
  2. 2.
    CAGrad (Liu et al., 2021). Extends PCGrad by adaptively adjusting gradient magnitudes based on conflict severity. This proves more nuanced modifications to gradients than binary projection.
  3. 3.
    Adatask (Yang et al., 2023). Dynamically reweighs task losses using relative loss changes, adapting to varying task learning rates during training.
  4. 4.
    FAMO (Liu et al., 2023). Fast Adaptive Multitask Optimization dynamically adjusts task weights to equalize each task’s rate of loss improvement. It uses an online, per-step rule (no pairwise gradient ops), adding negligible overhead while remaining robust to loss-scale differences.
  5. 5.
    Fair Resource Allocation in MTL (FairGrad) (Ban & Ji, 2024). Views the shared update as a limited resource and chooses it to maximize an $\alpha$-fair utility of per-task improvements. The parameter $\alpha$ controls the trade-off between average performance and fairness.
  6. 6.
    Nash-MTL (Navon et al., 2022). Frames multitask training as a bargaining game and computes a scale-invariant weighted combination of task gradients given by the Nash bargaining solution. Weights are obtained by solving a small inner problem (e.g., via CCP) using the gradient Gram matrix. Updates are balanced across tasks.
Appendix L Expanded Wall-Clock Time Study
Table 4: We present wall-clock time (seconds $\pm$ standard deviation) across all $K$ and scheduler refresh rates $R\in\{4,32,256\}$. We split results into sub-tables by $R$ for readability. Non-scheduler methods do not depend on $R$, so they are shown in the $R=4$ sub-table and omitted in the R=32,256R{=}32,256 subtables to avoid redundancy.
(a) R=4R{=}4 (all methods)
Method K=3 K=6 K=16 K=40
Uniform 0.2656 $\pm$ 0.1201 0.3240 $\pm$ 0.0629 0.3798 $\pm$ 0.1050 0.4054 $\pm$ 0.1190
GradNorm 5.4714 $\pm$ 0.7137 5.1201 $\pm$ 0.6112 4.9042 $\pm$ 0.5869 4.7372 $\pm$ 0.9286
MGDA 97.1081 $\pm$ 5.4645 121.4371 $\pm$ 9.0923 132.4913 $\pm$ 3.1752 134.0878 $\pm$ 2.2621
PCGrad 3.6212 $\pm$ 0.3517 23.1266 $\pm$ 0.8773 176.7566 $\pm$ 2.8171 1127.1337 $\pm$ 34.2603
CAGrad 102.8651 $\pm$ 18.3422 136.1034 $\pm$ 2.4218 134.3585 $\pm$ 4.0791 132.7034 $\pm$ 1.2412
AdaTask 2.1816 $\pm$ 0.0934 2.1032 $\pm$ 0.1012 2.2853 $\pm$ 0.0718 2.2278 $\pm$ 0.1370
FAMO 2.0725 $\pm$ 0.2073 1.9980 $\pm$ 0.1998 2.1710 $\pm$ 0.2171 2.1164 $\pm$ 0.2116
FairGrad 3.8020 $\pm$ 0.5703 15.2079 $\pm$ 2.2812 108.1450 $\pm$ 16.2218 675.9065 $\pm$ 101.3860
Nash-MTL 5.7030 $\pm$ 1.1406 22.8118 $\pm$ 4.5624 162.2176 $\pm$ 32.4435 1013.8598 $\pm$ 202.7720
SON-GOKU 2.0904 $\pm$ 0.3506 3.6770 $\pm$ 0.4974 6.3225 $\pm$ 0.7895 14.3280 $\pm$ 1.4073
SON-GOKU ++ AdaTask 4.1011 $\pm$ 0.4174 5.2126 $\pm$ 0.6066 7.6798 $\pm$ 0.7107 14.7528 $\pm$ 1.8671
SON-GOKU ++ GradNorm 7.3223 $\pm$ 0.4994 8.5898 $\pm$ 0.8203 12.1065 $\pm$ 2.5850 16.8329 $\pm$ 1.9803
SON-GOKU ++ PCGrad 2.3489 $\pm$ 0.3258 3.5925 $\pm$ 0.4100 6.1549 $\pm$ 0.8461 12.5729 $\pm$ 1.2657
(b) R=32R{=}32 (scheduler-based approaches)
Method K=3 K=6 K=16 K=40
SON-GOKU 1.9896 $\pm$ 0.3651 3.3202 $\pm$ 0.5745 6.0897 $\pm$ 0.9425 12.1432 $\pm$ 1.2044
SON-GOKU ++ AdaTask 3.7718 $\pm$ 0.9654 5.0511 $\pm$ 0.6531 7.5903 $\pm$ 1.1920 14.5182 $\pm$ 2.0660
SON-GOKU ++ GradNorm 7.0202 $\pm$ 1.0711 8.1661 $\pm$ 0.9355 10.7227 $\pm$ 2.2088 16.5760 $\pm$ 1.8418
SON-GOKU ++ PCGrad 1.9834 $\pm$ 0.3586 3.4971 $\pm$ 0.3840 6.1395 $\pm$ 0.9425 10.9097 $\pm$ 1.5263
(c) R=256R{=}256 (scheduler-based approaches)
Method K=3 K=6 K=16 K=40
SON-GOKU 1.7593 $\pm$ 0.2280 3.0024 $\pm$ 0.3942 4.8411 $\pm$ 0.7302 11.4162 $\pm$ 1.6076
SON-GOKU ++ AdaTask 3.7224 $\pm$ 0.2696 4.4548 $\pm$ 0.5837 7.5276 $\pm$ 0.6230 13.0608 $\pm$ 3.2925
SON-GOKU ++ GradNorm 6.0221 $\pm$ 1.0418 7.8659 $\pm$ 0.7917 9.5029 $\pm$ 1.2168 15.6860 $\pm$ 2.3680
SON-GOKU ++ PCGrad 1.6776 $\pm$ 0.4104 3.0189 $\pm$ 0.7854 5.9893 $\pm$ 1.3797 7.1915 $\pm$ 0.2021
We provide more results from our wall-clock time study. The expanded table includes results from testing refresh rates $R\in\{4,32,256\}$ for scheduler-based methods.
L.1 Experimental Setup for Wall-Clock Time Study
We benchmark wall-clock time with a controlled synthetic workload to remove the effects of data loading and I/O. For each configuration (number of tasks $K$ and scheduler refresh rates $R$), we pre-generate a fixed sequence of per-task gradient vectors and loss values directly on the target device, and then feed the same exact tensors, in the same exact order, to every method. We set the gradient dimensionality to 1024. Timing uses a high-resolution clock with a device synchronize before starting and after finishing to capture only on-device compute. We also accumulate the norm of the combined gradient into a scalar accumulator (also known as a scalar sink) so the backend must realize the computation, avoiding lazy evaluation. Each MTL approach is run for 900 steps and repeated 10 times.
Appendix M Extended Related Work
Multi-task learning (MTL) methods have evolved from simple loss-weighting approaches to larger and more sophisticated optimization techniques that manage task conflict and cooperation (Yang et al., 2023). Early adaptive-weighting approaches sought to balance losses automatically (Vandenhende et al., 2022; Fan et al., 2023), while more recent work modifies gradients directly (Yu et al., 2020). Task scheduling and grouping methods, though far less popular than adaptive weighting techniques (Torbarina et al., 2023), have contributed to the field by controlling the timing of updates.
M.1 Tuned Loss Weighting
From early MTL work it became clear that simply summing task losses often favors one objective at the expense of others (Kurin et al., 2022; Zhao et al., 2024; Mueller et al., 2022), especially when losses have different scales or noise levels. To address this, practitioners manually tuned per‑task weight coefficients ($\lambda$-values) to rebalance learning (Argyriou et al., 2007; Ando & Zhang, 2005; Evgeniou et al., 2005; Kang et al., 2011; Liang & Zhang, 2020; Lin et al., 2022; Yu et al., 2021), but this process was laborious and dataset-specific. Thus, researchers began to develop automated methods.
M.2 Adaptive Loss Weighting
(Kendall et al., 2018) introduced uncertainty weighting, learning each task’s homoscedastic (constant‑variance) (Bishop, 2006) noise to scale losses automatically and improve depth and semantics on NYUv2 (Silberman et al., 2012).
GradNorm automatically balances multiple loss functions by tuning each task’s gradient magnitude so that all tasks train at comparable speeds (Chen et al., 2018). It does this by introducing a single asymmetry hyperparameter $\alpha$ that governs how much each task’s loss is scaled. This eliminates the need for expensive grid searches over manual weights. GradNorm was also a major leap empirically as it surpassed exhaustive search baselines on both regression and classification tasks. Dynamic Weight Averaging (DWA) extended this idea by adjusting weights based on loss rate of change, reducing oscillations between tasks (Liu et al., 2019).
More recently AdaTask applies task-specific learning rates that adapt to each head’s gradient norm, yielding significant gains on multi-label classification benchmarks (Yang et al., 2023).
M.3 Gradient-Level Conflict Mitigation
Rather than rescaling losses, gradient surgery methods alter update directions. PCGrad projects gradients that conflict (negative cosine) onto each other’s normal plane, significantly boosting efficiency on supervised vision and RL problems (Yu et al., 2020). CAGrad frames task balance as a min-max optimization, finding updates that maximize the worst-case task improvement (Liu et al., 2021). The Multiple Gradient Descent Algorithm (MGDA) computes a Pareto-optimal convex combination of task gradients, ensuring no task is harmed (Sener & Koltun, 2018). More recent variants such as SAM-GS incorporate momentum into conflict detection, smoothing gradient estimates while preserving the benefits of surgery (Borsani et al., 2025).
M.4 Task Grouping
Task grouping aims to decide which tasks should train together so that helpful transfer is amplified and harmful interference is limited. It typically groups tasks into subsets that update jointly, rather than updating all tasks at once. This is different from approaches that keep all tasks active or reweight the joint gradient (adaptive loss weighting, gradient surgery).
Early approaches under this category used round-robin and random sampling-based approaches that ignored any task relationships (McCann et al., 2018; Zamir et al., 2020). Standley et al. (2020) exhaustively searches over small subsets to identify beneficial groupings, demonstrating the potential of selective updates but failing to scale beyond eight tasks due to computational complexity.
Task Affinity Groupings (TAG) (Fifty et al., 2021) performs one joint training run to measure inter-task ’affinity’. It quantifies how an update for task $i$ (its gradient) would change task $j$’s loss, and it uses these cross-effects to select partitions of tasks that should share updates. The key idea is to treat grouping as an outcome of measured gradient interactions.
Ayman et al. (Ayman et al., 2023) train a predictor that maps single-task statistics and dataset features to an estimate of whether two or more tasks should be grouped. They then use that predictor to guide a randomized search over groups, which dramatically reduces the number of multi-task trainings (or ’MTL trials’) needed to find a good partition.
Using a completely different approach, Towards Principled Task Grouping (PTG) (Wang et al., 2024) formulates grouping as a mathematical program with a theoretically motivated objective capturing beneficial transfer while respecting resource constraints (e.g., compute budgets). It builds a principled optimization over candidate groups that is meant to generalize across application domains.
Scalable Task Grouping via Training Dynamics (STG-MLT) (Sherif et al., 2024) avoids expensive affinity estimation by extracting Data Maps (Swayamdipta et al., 2020) (simple summaries of training dynamics per task) and then clustering tasks using those features. The clusters are intended to push for positive transfer at larger scale. This approach essentially replaces gradient cross-effects with more compact trajectory features that are cheap to compute and easy to cluster.